The NetCDF Fortran 77 Interface Guide

NetCDF Version 4.0.1 January 2009

Russ Rew, Glenn Davis, Steve Emmerson, and Harvey Davies Unidata Program Center

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2005-2006 University Corporation for Atmospheric Research

Permission is granted to make and distribute verbatim copies of this manual provided that the copyright notice and these paragraphs are preserved on all copies. The software and any accompanying written materials are provided “as is” without warranty of any kind. UCAR expressly disclaims all warranties of any kind, either expressed or implied, including but not limited to the implied warranties of merchantability and fitness for a particular purpose. The Unidata Program Center is managed by the University Corporation for Atmospheric Research and sponsored by the National Science Foundation. Any opinions, findings, conclusions, or recommendations expressed in this publication are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Mention of any commercial company or product in this document does not constitute an endorsement by the Unidata Program Center. Unidata does not authorize any use of information from this publication for advertising or publicity purposes.

Table of Contents

1 Use of the NetCDF Library. . . . . . . . . . . . . . . . . . . . 1

1.1 Creating a NetCDF Dataset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Reading a NetCDF Dataset with Known Names . . . . . . . . . . . . . . . . 2 1.3 Reading a netCDF Dataset with Unknown Names . . . . . . . . . . . . . . 3 1.4 Adding New Dimensions, Variables, Attributes. . . . . . . . . . . . . . . . . . 4 1.5 Error Handling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.6 Compiling and Linking with the NetCDF Library . . . . . . . . . . . . . . . 5

2 Datasets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.1 Datasets Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2 NetCDF Library Interface Descriptions . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.3 NF STRERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.4 Get netCDF library version: NF INQ LIBVERS . . . . . . . . . . . . . . . 8 2.5 NF CREATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.6 NF CREATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.7 NF CREATE PAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.8 NF OPEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.9 NF OPEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.10 NF OPEN PAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.11 NF REDEF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.12 NF ENDDEF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.13 NF ENDDEF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.14 NF CLOSE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.15 NF INQ Family. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 2.16 NF SYNC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.17 NF ABORT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 2.18 NF SET FILL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 2.19 NF SET DEFAULT FORMAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

3 Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

3.1 Find a Group ID: NF INQ NCID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.2 Get a List of Groups in a Group: NF INQ GRPS . . . . . . . . . . . . . 30 3.3 Find all the Variables in a Group: NF INQ VARIDS . . . . . . . . . . 30 3.4 Find all Dimensions Visible in a Group: NF INQ DIMIDS …. 31 3.5 Find the Length of a Group’s Name: NF INQ GRPNAME LEN

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 3.6 Find a Group’s Name: NF INQ GRPNAME . . . . . . . . . . . . . . . . . . 33 3.7 Find a Group’s Full Name: NF INQ GRPNAME FULL. . . . . . . 34 3.8 Find a Group’s Parent: NF INQ GRP PARENT . . . . . . . . . . . . . . 35 3.9 Create a New Group: NF DEF GRP. . . . . . . . . . . . . . . . . . . . . . . . . . 35

4 Dimensions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

4.1 Dimensions Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.2 NF DEF DIM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.3 NF INQ DIMID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.4 NF INQ DIM Family . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.5 NF RENAME DIM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

5 User Defined Data Types . . . . . . . . . . . . . . . . . . . . . 43

5.1 User Defined Types Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.2 Learn the IDs of All Types in Group: NF INQ TYPEIDS . . . . . 43 5.3 Learn About a User Defined Type: NF INQ TYPE . . . . . . . . . . . 44 5.4 Learn About a User Defined Type: NF INQ USER TYPE …. 45 5.5 Compound Types Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

5.5.1 Creating a Compound Type: NF DEF COMPOUND. . . . . 46 5.5.2 Inserting a Field into a Compound Type:

NF INSERT COMPOUND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.5.3 Inserting an Array Field into a Compound Type:

NF INSERT ARRAY COMPOUND. . . . . . . . . . . . . . . . . . . . . . . . 49 5.5.4 Learn About a Compound Type: NF INQ COMPOUND

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 5.5.5 Learn About a Field of a Compound Type:

NF INQ COMPOUND FIELD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 5.6 Variable Length Array Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

5.6.1 Define a Variable Length Array (VLEN): NF DEF VLEN

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 5.6.2 Learning about a Variable Length Array (VLEN) Type:

NF INQ VLEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 5.6.3 Releasing Memory for a Variable Length Array (VLEN)

Type: NF FREE VLEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 5.6.4 Set a Variable Length Array with

NF PUT VLEN ELEMENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 5.6.5 Set a Variable Length Array with

NF GET VLEN ELEMENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.7 Opaque Type Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

5.7.1 Creating Opaque Types: NF DEF OPAQUE . . . . . . . . . . . . . 58 5.7.2 Learn About an Opaque Type: NF INQ OPAQUE . . . . . . . 59

5.8 Enum Type Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 5.8.1 Creating a Enum Type: NF DEF ENUM . . . . . . . . . . . . . . . . 60 5.8.2 Inserting a Field into a Enum Type: NF INSERT ENUM

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 5.8.3 Learn About a Enum Type: NF INQ ENUM . . . . . . . . . . . . . 62 5.8.4 Learn the Name of a Enum Type: nf inq enum member . . 63 5.8.5 Learn the Name of a Enum Type: NF INQ ENUM IDENT

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

6 Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

6.1 Variables Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.2 Language Types Corresponding to netCDF external data types

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.3 Create a Variable: NF_DEF_VAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 6.4 Define Chunking Parameters for a Variable:

NF_DEF_VAR_CHUNKING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 6.5 Learn About Chunking Parameters for a Variable:

NF_INQ_VAR_CHUNKING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 6.6 Define Fill Parameters for a Variable: nf_def_var_fill . . . . . . . 70 6.7 Learn About Fill Parameters for a Variable: NF_INQ_VAR_FILL

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 6.8 Define Compression Parameters for a Variable:

NF_DEF_VAR_DEFLATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 6.9 Learn About Deflate Parameters for a Variable:

NF_INQ_VAR_DEFLATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 6.10 Define Checksum Parameters for a Variable:

NF_DEF_VAR_FLETCHER32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 6.11 Learn About Checksum Parameters for a Variable:

NF_INQ_VAR_FLETCHER32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 6.12 Define Endianness of a Variable: NF_DEF_VAR_ENDIAN . . . . . . . . 77 6.13 Learn About Endian Parameters for a Variable:

NF_INQ_VAR_ENDIAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 6.14 Get a Variable ID from Its Name: NF INQ VARID . . . . . . . . . . 80 6.15 Get Information about a Variable from Its ID: NF INQ VAR

family . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 6.16 Write a Single Data Value: NF PUT VAR1 type . . . . . . . . . . . . 82 6.17 Write an Entire Variable: NF PUT VAR type . . . . . . . . . . . . . . 84 6.18 Write an Array of Values: NF PUT VARA type . . . . . . . . . . . . 85 6.19 NF PUT VARS type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 6.20 NF PUT VARM type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 6.21 NF GET VAR1 type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 6.22 NF GET VAR type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 6.23 NF GET VARA type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 6.24 NF GET VARS type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 6.25 NF GET VARM type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 6.26 Reading and Writing Character String Values . . . . . . . . . . . . . . . 105 6.27 Fill Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 6.28 NF RENAME VAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 6.29 Change between Collective and Independent Parallel Access:

NF VAR PAR ACCESS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108

7 Attributes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

7.1 Attributes Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 7.2 Attribute Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 7.3 NF PUT ATT type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 7.4 NF INQ ATT Family . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 7.5 NF GET ATT type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 7.6 NF COPY ATT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 7.7 NF RENAME ATT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 7.8 NF DEL ATT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

Appendix A NetCDF 2 to NetCDF 3 Fortran 77

Transition Guide . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

A.1 Overview of FORTRAN interface changes. . . . . . . . . . . . . . . . . . . . 125 A.2 The New FORTRAN Interface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 A.3 Function Naming Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 A.4 Type Conversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

Appendix B Summary of FORTRAN 77 Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

1 Use of the NetCDF Library

You can use the netCDF library without knowing about all of the netCDF interface. If you are creating a netCDF dataset, only a handful of routines are required to define the necessary dimensions, variables, and attributes, and to write the data to the netCDF dataset. (Even less are needed if you use the ncgen utility to create the dataset before running a program using netCDF library calls to write data. See Section “ncgen” in The NetCDF Users Guide.) Similarly, if you are writing software to access data stored in a particular netCDF object, only a small subset of the netCDF library is required to open the netCDF dataset and access the data. Authors of generic applications that access arbitrary netCDF datasets need to be familiar with more of the netCDF library.

In this chapter we provide templates of common sequences of netCDF calls needed for common uses. For clarity we present only the names of routines; omit declarations and error checking; omit the type-specific suffixes of routine names for variables and attributes; indent statements that are typically invoked multiple times; and use ... to represent arbitrary sequences of other statements. Full parameter lists are described in later chapters.

1.1 Creating a NetCDF Dataset

Here is a typical sequence of netCDF calls used to create a new netCDF dataset:

         NF_CREATE           ! create netCDF dataset: enter define mode
              ...
            NF_DEF_DIM       ! define dimensions: from name and length
              ...
            NF_DEF_VAR       ! define variables: from name, type, dims
              ...
            NF_PUT_ATT       ! assign attribute values
              ...
         NF_ENDDEF           ! end definitions: leave define mode
              ...
            NF_PUT_VAR       ! provide values for variable
              ...
         NF_CLOSE            ! close: save new netCDF dataset

Only one call is needed to create a netCDF dataset, at which point you will be in the first of two netCDF modes. When accessing an open netCDF dataset, it is either in define mode or data mode. In define mode, you can create dimensions, variables, and new attributes, but you cannot read or write variable data. In data mode, you can access data and change existing attributes, but you are not permitted to create new dimensions, variables, or attributes.

One call to NF DEF DIM is needed for each dimension created. Similarly, one call to NF DEF VAR is needed for each variable creation, and one call to a member of the NF PUT ATT family is needed for each attribute defined and assigned a value. To leave define mode and enter data mode, call NF ENDDEF.

Once in data mode, you can add new data to variables, change old values, and change values of existing attributes (so long as the attribute changes do not require more storage space). Single values may be written to a netCDF variable with one of the members of the NF PUT VAR1 family, depending on what type of data you have to write. All the values of a variable may be written at once with one of the members of the NF PUT VAR family. Arrays or array cross-sections of a variable may be written using members of the NF PUT VARA family. Subsampled array sections may be written using members of the NF PUT VARS family. Mapped array sections may be written using members of the NF PUT VARM family. (Subsampled and mapped access are general forms of data access that are explained later.)

Finally, you should explicitly close all netCDF datasets that have been opened for writing by calling NF CLOSE. By default, access to the file system is buffered by the netCDF library. If a program terminates abnormally with netCDF datasets open for writing, your most recent modifications may be lost. This default buffering of data is disabled by setting the NF SHARE flag when opening the dataset. But even if this flag is set, changes to attribute values or changes made in define mode are not written out until NF SYNC or NF CLOSE is called.

1.2 Reading a NetCDF Dataset with Known Names

Here we consider the case where you know the names of not only the netCDF datasets, but also the names of their dimensions, variables, and attributes. (Otherwise you would have to do "inquire" calls.) The order of typical C calls to read data from those variables in a netCDF dataset is:

         NF_OPEN               ! open existing netCDF dataset
              ...
            NF_INQ_DIMID       ! get dimension IDs
              ...
            NF_INQ_VARID       ! get variable IDs
              ...
            NF_GET_ATT         ! get attribute values
              ...
            NF_GET_VAR         ! get values of variables
              ...
         NF_CLOSE              ! close netCDF dataset

First, a single call opens the netCDF dataset, given the dataset name, and returns a netCDF ID that is used to refer to the open netCDF dataset in all subsequent calls.

Next, a call to NF INQ DIMID for each dimension of interest gets the dimension ID from the dimension name. Similarly, each required variable ID is determined from its name by a call to NF INQ VARID.Once variable IDs are known, variable attribute values can be retrieved using the netCDF ID, the variable ID, and the desired attribute name as input to a member of the NF GET ATT family (typically NF GET ATT TEXT or NF GET ATT DOUBLE) for each desired attribute. Variable data values can be directly accessed from the netCDF dataset with calls to members of the NF GET VAR1 family for single values, the NF GET VAR family for entire variables, or various other members of the NF GET VARA, NF GET VARS, or NF GET VARM families for array, subsampled or mapped access.

Finally, the netCDF dataset is closed with NF CLOSE. There is no need to close a dataset open only for reading.

1.3 Reading a netCDF Dataset with Unknown Names

It is possible to write programs (e.g., generic software) which do such things as processing every variable, without needing to know in advance the names of these variables. Similarly, the names of dimensions and attributes may be unknown.

Names and other information about netCDF objects may be obtained from netCDF datasets by calling inquire functions. These return information about a whole netCDF dataset, a dimension, a variable, or an attribute. The following template illustrates how they are used:

         NF_OPEN                   ! open existing netCDF dataset
           ...
         NF_INQ                    ! find out what is in it
              ...
            NF_INQ_DIM             ! get dimension names, lengths
              ...
            NF_INQ_VAR             ! get variable names, types, shapes
                 ...
               NF_INQ_ATTNAME      ! get attribute names
                 ...
               NF_INQ_ATT          ! get attribute values
                 ...
               NF_GET_ATT          ! get attribute values
                 ...
            NF_GET_VAR             ! get values of variables
              ...
         NF_CLOSE                  ! close netCDF dataset

As in the previous example, a single call opens the existing netCDF dataset, returning a netCDF ID. This netCDF ID is given to the NF INQ routine, which returns the number of dimensions, the number of variables, the number of global attributes, and the ID of the unlimited dimension, if there is one.

All the inquire functions are inexpensive to use and require no I/O, since the information they provide is stored in memory when a netCDF dataset is first opened.

Dimension IDs use consecutive integers, beginning at 1. Also dimensions, once created, cannot be deleted. Therefore, knowing the number of dimension IDs in a netCDF dataset means knowing all the dimension IDs: they are the integers 1, 2, 3, ... up to the number of dimensions. For each dimension ID, a call to the inquire function NF INQ DIM returns the dimension name and length.

Variable IDs are also assigned from consecutive integers 1, 2, 3, ... up to the number of variables. These can be used in NF INQ VAR calls to find out the names, types, shapes, and the number of attributes assigned to each variable.

Once the number of attributes for a variable is known, successive calls to NF INQ ATTNAME return the name for each attribute given the netCDF ID, variable ID, and attribute number. Armed with the attribute name, a call to NF INQ ATT returns its type and length. Given the type and length, you can allocate enough space to hold the attribute values. Then a call to a member of the NF GET ATT family returns the attribute values.

Once the IDs and shapes of netCDF variables are known, data values can be accessed by calling a member of the NF GET VAR1 family for single values, or members of the NF GET VAR, NF GET VARA, NF GET VARS, or NF GET VARM for various kinds of array access.

1.4 Adding New Dimensions, Variables, Attributes

An existing netCDF dataset can be extensively altered. New dimensions, variables, and attributes can be added or existing ones renamed, and existing attributes can be deleted. Existing dimensions, variables, and attributes can be renamed. The following code template lists a typical sequence of calls to add new netCDF components to an existing dataset:

         NF_OPEN             ! open existing netCDF dataset
           ...
         NF_REDEF            ! put it into define mode
             ...
           NF_DEF_DIM        ! define additional dimensions (if any)
             ...
           NF_DEF_VAR        ! define additional variables (if any)
             ...
           NF_PUT_ATT        ! define other attributes (if any)
             ...
         NF_ENDDEF           ! check definitions, leave define mode
             ...
           NF_PUT_VAR        ! provide new variable values
             ...
         NF_CLOSE            ! close netCDF dataset

A netCDF dataset is first opened by the NF OPEN call. This call puts the open dataset in data mode, which means existing data values can be accessed and changed, existing attributes can be changed (so long as they do not grow), but nothing can be added. To add new netCDF dimensions, variables, or attributes you must enter define mode, by calling NF REDEF.In define mode, call NF DEF DIM to define new dimensions, NF DEF VAR to define new variables, and a member of the NF PUT ATT family to assign new attributes to variables or enlarge old attributes.

You can leave define mode and reenter data mode, checking all the new definitions for consistency and committing the changes to disk, by calling NF ENDDEF. If you do not wish to reenter data mode, just call NF CLOSE, which will have the effect of first calling NF ENDDEF.

Until the NF ENDDEF call, you may back out of all the redefinitions made in define mode and restore the previous state of the netCDF dataset by calling NF ABORT. You may also use the NF ABORT call to restore the netCDF dataset to a consistent state if the call to NF ENDDEF fails. If you have called NF CLOSE from definition mode and the implied call to NF ENDDEF fails, NF ABORT will automatically be called to close the netCDF dataset and leave it in its previous consistent state (before you entered define mode).

At most one process should have a netCDF dataset open for writing at one time. The library is designed to provide limited support for multiple concurrent readers with one writer, via disciplined use of the NF SYNC function and the NF SHARE flag. If a writer makes changes in define mode, such as the addition of new variables, dimensions, or attributes, some means external to the library is necessary to prevent readers from making concurrent accesses and to inform readers to call NF SYNC before the next access.

1.5 Error Handling

The netCDF library provides the facilities needed to handle errors in a flexible way. Each netCDF function returns an integer status value. If the returned status value indicates an error, you may handle it in any way desired, from printing an associated error message and exiting to ignoring the error indication and proceeding (not recommended!). For simplicity, the examples in this guide check the error status and call a separate function to handle any errors.

The NF STRERROR function is available to convert a returned integer error status into an error message string.

Occasionally, low-level I/O errors may occur in a layer below the netCDF library. For example, if a write operation causes you to exceed disk quotas or to attempt to write to a device that is no longer available, you may get an error from a layer below the netCDF library, but the resulting write error will still be reflected in the returned status value.

1.6 Compiling and Linking with the NetCDF Library

Details of how to compile and link a program that uses the netCDF C or FORTRAN interfaces differ, depending on the operating system, the available compilers, and where the netCDF library and include files are installed. Nevertheless, we provide here examples of how to compile and link a program that uses the netCDF library on a Unix platform, so that you can adjust these examples to fit your installation.

Every FORTRAN file that references netCDF functions or constants must contain an appropriate INCLUDE statement before the first such reference:

     INCLUDE 'netcdf.inc'

Unless the netcdf.inc file is installed in a standard directory where the FORTRAN compiler always looks, you must use the -I option when invoking the compiler, to specify a directory where netcdf.inc is installed, for example:

     f77 -c -I/usr/local/netcdf/include myprogram.f

Alternatively, you could specify an absolute path name in the INCLUDE statement, but then your program would not compile on another platform where netCDF is installed in a different location.

Unless the netCDF library is installed in a standard directory where the linker always looks, you must use the -L and -l options to link an object file that uses the netCDF library. For example:

     f77 -o myprogram myprogram.o -L/usr/local/netcdf/lib -lnetcdf

Alternatively, you could specify an absolute path name for the library:

     f77 -o myprogram myprogram.o -l/usr/local/netcdf/lib/libnetcdf.

2 Datasets

2.1 Datasets Introduction

This chapter presents the interfaces of the netCDF functions that deal with a netCDF dataset or the whole netCDF library.

A netCDF dataset that has not yet been opened can only be referred to by its dataset name. Once a netCDF dataset is opened, it is referred to by a netCDF ID, which is a small nonnegative integer returned when you create or open the dataset. A netCDF ID is much like a file descriptor in C or a logical unit number in FORTRAN. In any single program, the netCDF IDs of distinct open netCDF datasets are distinct. A single netCDF dataset may be opened multiple times and will then have multiple distinct netCDF IDs; however at most one of the open instances of a single netCDF dataset should permit writing. When an open netCDF dataset is closed, the ID is no longer associated with a netCDF dataset.

Functions that deal with the netCDF library include:

• Get version of library.

• Get error message corresponding to a returned error code.

The operations supported on a netCDF dataset as a single object are:

• Create, given dataset name and whether to overwrite or not.

• Open for access, given dataset name and read or write intent.

• Put into define mode, to add dimensions, variables, or attributes.

• Take out of define mode, checking consistency of additions.

• Close, writing to disk if required.

• Inquire about the number of dimensions, number of variables, number of global attributes, and ID of the unlimited dimension, if any.

• Synchronize to disk to make sure it is current.

• Set and unset nofill mode for optimized sequential writes.

• After a summary of conventions used in describing the netCDF interfaces, the rest of this chapter presents a detailed description of the interfaces for these operations.

2.2 NetCDF Library Interface Descriptions

Each interface description for a particular netCDF function in this and later chapters contains:

• a description of the purpose of the function;

• a FORTRAN function prototype that presents the type and order of the formal parameters to the function;

• a description of each formal parameter in the C interface;

• a list of possible error conditions; and

• an example of a FORTRAN program fragment calling the netCDF function (and perhaps other netCDF functions).

The examples follow a simple convention for error handling, always checking the error status returned from each netCDF function call and calling a handle error function in case an error was detected. For an example of such a function, see Section 5.2 "Get error message corresponding to error status: nf strerror".

2.3 NF STRERROR

The function NF STRERROR returns a static reference to an error message string corresponding to an integer netCDF error status or to a system error number, presumably returned by a previous call to some other netCDF function. The list of netCDF error status codes is available in the appropriate include file for each language binding.

Usage

     CHARACTER*80 FUNCTION NF_STRERROR(INTEGER NCERR)

NCERR An error status that might have been returned from a previous call to some netCDF function.

Errors

If you provide an invalid integer error status that does not correspond to any netCDF error message or or to any system error message (as understood by the system strerror function), NF STRERROR returns a string indicating that there is no such error status.

Example

Here is an example of a simple error handling function that uses NF STRERROR to print the error message corresponding to the netCDF error status returned from any netCDF function call and then exit:

     INCLUDE 'netcdf.inc'
        ...
     SUBROUTINE HANDLE_ERR(STATUS)
     INTEGER STATUS
     IF (STATUS .NE. NF_NOERR) THEN
       PRINT *, NF_STRERROR(STATUS)
       STOP 'Stopped'
     ENDIF
     END

2.4 Get netCDF library version: NF INQ LIBVERS

The function NF INQ LIBVERS returns a string identifying the version of the netCDF library, and when it was built.

Usage

     CHARACTER*80 FUNCTION NF_INQ_LIBVERS()

Errors

This function takes no arguments, and thus no errors are possible in its invocation.

Example

Here is an example using nf inq libvers to print the version of the netCDF library with which the program is linked:

INCLUDE 'netcdf.inc'
...
PRINT *, NF_INQ_LIBVERS()

2.5 NF CREATE

This function creates a new netCDF dataset, returning a netCDF ID that can subsequently be used to refer to the netCDF dataset in other netCDF function calls. The new netCDF dataset opened for write access and placed in define mode, ready for you to add dimensions, variables, and attributes.

A creation mode flag specifies whether to overwrite any existing dataset with the same name and whether access to the dataset is shared.

Usage

INTEGER FUNCTION NF_CREATE (CHARACTER*(*) PATH, INTEGER CMODE,
                       INTEGER ncid)

PATH The file name of the new netCDF dataset.

CMODE  The creation mode flag.  The following flags are available:

NF NOCLOBBER, NF SHARE, NF 64BIT OFFSET, NF NETCDF4 and NF CLASSIC MODEL. You can combine the affect of multiple flags in a single argument by using the bitwise OR operator. For example, to specify both NF NOCLOBBER and NF SHARE, you could provide the argument OR(NF NOCLOBBER, NF SHARE). A zero value (defined for convenience as NF CLOBBER) specifies the default behavior: overwrite any existing dataset with the same file name and buffer and cache accesses for efficiency. The dataset will be in netCDF classic format. See Section “NetCDF Classic Format Limitations” in The NetCDF Users Guide. Setting NF NOCLOBBER means you do not want to clobber (overwrite) an existing dataset; an error (NF EEXIST) is returned if the specified dataset already exists. The NF SHARE flag is appropriate when one process may be writing the dataset and one or more other processes reading the dataset concurrently; it means that dataset accesses are not buffered and caching is limited. Since the buffering scheme is optimized for sequential access, programs that do not access data sequentially may see some performance improvement by setting the NF SHARE flag. This only applied to classic and 64-bit offset format files. Setting NF 64BIT OFFSET causes netCDF to create a 64-bit offset format file, instead of a netCDF classic format file. The 64-bit offset format imposes far fewer restrictions on very large (i.e. over 2 GB) data files. See Section “Large File Support” in The NetCDF Users Guide. Setting NF NETCDF4 causes netCDF to create a netCDF-4/HDF5 format file. Oring NF CLASSIC MODEL with NF NETCDF4 causes the netCDF library to create a netCDF-4/HDF5 data file, with the netCDF classic model enforced - none of the new features of the netCDF-4 data model may be usedin such a file, for example groups and user-defined types.

ncid Returned netCDF ID.

Errors

NF CREATE returns the value NF NOERR if no errors occurred. Possible causes of errors include:

• Passing a dataset name that includes a directory that does not exist.

• Specifying a dataset name of a file that exists and also specifying NF NOCLOBBER.

• Specifying a meaningless value for the creation mode.

• Attempting to create a netCDF dataset in a directory where you don’t have permission to create files.

Example

In this example we create a netCDF dataset named foo.nc; we want the dataset to be created in the current directory only if a dataset with that name does not already exist:

     INCLUDE 'netcdf.inc'
       ...
     INTEGER NCID, STATUS
     ...
     STATUS = NF_CREATE('foo.nc', NF_NOCLOBBER, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

2.6 NF CREATE

This function is a variant of NF CREATE, NF CREATE (note the double underscore) allows users to specify two tuning parameters for the file that it is creating. These tuning parameters are not written to the data file, they are only used for so long as the file remains open after an NF CREATE.

This function creates a new netCDF dataset, returning a netCDF ID that can subsequently be used to refer to the netCDF dataset in other netCDF function calls. The new netCDF dataset opened for write access and placed in define mode, ready for you to add dimensions, variables, and attributes.

A creation mode flag specifies whether to overwrite any existing dataset with the same name and whether access to the dataset is shared.

Usage

     INTEGER FUNCTION NF_CREATE (CHARACTER*(*) PATH, INTEGER CMODE, INTEGER INITIALSZ,
                                 INTEGER CHUNKSIZEHINT, INTEGER ncid)

PATH The file name of the new netCDF dataset.

CMODE  The creation mode flag.  The following flags are available:

NF NOCLOBBER, NF SHARE, NF 64BIT OFFSET, NF NETCDF4, and NF CLASSIC MODEL. Setting NF NOCLOBBER means you do not want to clobber (overwrite) an existing dataset; an error (NF EEXIST) is returned if the specified dataset already exists. The NF SHARE flag is appropriate when one process may be writing the dataset and one or more other processes reading the dataset concurrently; it means that dataset accesses are not buffered and caching is limited. Since the buffering scheme is optimized for sequential access, programs that do not access data sequentially may see some performance improvement by setting the NF SHARE flag. This flag has no effect with netCDF-4/HDF5 files. Setting NF 64BIT OFFSET causes netCDF to create a 64-bit offset format file, instead of a netCDF classic format file. The 64-bit offset format imposes far fewer restrictions on very large (i.e. over 2 GB) data files. See Section “Large File Support” in The NetCDF Users Guide. Setting NF CLASSIC MODEL causes netCDF to enforce the classic data model in this file. (This only has effect for netCDF-4/HDF5 files, as classic and 64-bit offset files always use the classic model.) When used with NF NETCDF4, this flag ensures that the resulting netCDF-4/HDF5 file may never contain any new constructs from the enhanced data model. That is, it cannot contain groups, user defined types, multiple unlimited dimensions, or new atomic types. The advantage of this restriction is that such files are guarenteed to work with existing netCDF software. A zero value (defined for convenience as NF CLOBBER) specifies the default behavior: overwrite any existing dataset with the same file name and buffer and cache accesses for efficiency. The dataset will be in netCDF classic format. See Section “NetCDF Classic Format Limitations” in The NetCDF Users Guide.

initialsz

This parameter sets the initial size of the file at creation time.

chunksizehintp

The argument referenced by chunksizehintp controls a space versus time tradeoff, memory allocated in the netcdf library versus number of system calls. Because of internal requirements, the value may not be set to exactly the value requested. The actual value chosen is returned by reference. Using the value NF SIZEHINT DEFAULT causes the library to choose a default. How the system chooses the default depends on the system. On many systems, the "preferred I/O block size" is available from the stat() system call, struct stat member st blksize. If this is available it is used. Lacking that, twice the system pagesize is used. Lacking a call to discover the system pagesize, we just set default chunksize to The chunksize is a property of a given open netcdf descriptor ncid, it is not a persistent property of the netcdf dataset.

ncid Returned netCDF ID.

Errors

NF CREATE returns the value NF NOERR if no errors occurred. Possible causes of errors include:

• Passing a dataset name that includes a directory that does not exist.

• Specifying a dataset name of a file that exists and also specifying NF NOCLOBBER.

• Specifying a meaningless value for the creation mode.

• Attempting to create a netCDF dataset in a directory where you don’t have permission to create files.

Example

In this example we create a netCDF dataset named foo.nc; we want the dataset to be created in the current directory only if a dataset with that name does not already exist:

     INCLUDE 'netcdf.inc'
       ...
     INTEGER NCID, STATUS, INITIALSZ, CHUNKSIZEHINT
     ...
     INITIALSZ = 2048
     CHUNKSIZEHINT = 1024
     STATUS = NF_CREATE('foo.nc', NF_NOCLOBBER, INITIALSZ, CHUNKSIZEHINT, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

2.7 NF CREATE PAR

This function is a variant of nf create, nf create par allows users to open a file on a MPI/IO or MPI/Posix parallel file system.

The parallel parameters are not written to the data file, they are only used for so long as the file remains open after an nf create par.

This function is only available if the netCDF library was built with a HDF5 library for which –enable-parallel was used, and which was linked (like HDF5) to MPI libraries.

This function creates a new netCDF dataset, returning a netCDF ID that can subsequently be used to refer to the netCDF dataset in other netCDF function calls. The new netCDF dataset opened for write access and placed in define mode, ready for you to add dimensions, variables, and attributes.

This function is only available for netCDF-4 files. The creation mode flag must include NF NETCDF4.

When a netCDF-4 file is created for parallel access, collective operations are the default. To use independent access on a variable, See Section 6.29 [NF VAR PAR ACCESS], page 108.

Usage

     INTEGER FUNCTION NF_CREATE_PAR(CHARACTER*(*) PATH, INTEGER CMODE,
                                    INTEGER MPI_COMM, INTEGER MPI_INFO,
                                    INTEGER ncid)

PATH The file name of the new netCDF dataset.

CMODE The creation mode flag. The following flags are available: NF NOCLOBBER, NF NETCDF4 and NF CLASSIC MODEL. You can combine the affect of multiple flags in a single argument by using the bitwise OR operator. For example, to specify both NF NOCLOBBER and NF NETCDF4, you could provide the argument OR(NF NOCLOBBER, NF NETCDF4). Setting NF NETCDF4 causes netCDF to create a netCDF-4/HDF5 format file. Oring NF CLASSIC MODEL with NF NETCDF4 causes the netCDF library to create a netCDF-4/HDF5 data file, with the netCDF classic model enforced - none of the new features of the netCDF-4 data model may be usedin such a file, for example groups and user-defined types. Only netCDF-4/HDF5 files may be used with parallel I/O.

MPI_COMM The MPI communicator.

MPI_INFO The MPI info.

ncid Returned netCDF ID.

Errors

NF CREATE returns the value NF NOERR if no errors occurred. Possible causes of errors include:

• Passing a dataset name that includes a directory that does not exist.

• Specifying a dataset name of a file that exists and also specifying NF NOCLOBBER.

• Specifying a meaningless value for the creation mode.

• Attempting to create a netCDF dataset in a directory where you don’t have permission to create files.

Example

This example is from test program nf test/ftst parallel.F.

     !     Create the netCDF file.
           mode_flag = IOR(nf_netcdf4, nf_classic_model)
           retval = nf_create_par(FILE_NAME, mode_flag, MPI_COMM_WORLD,
          $     MPI_INFO_NULL, ncid)
           if (retval .ne. nf_noerr) stop 2

2.8 NF OPEN

The function NF OPEN opens an existing netCDF dataset for access.

Usage

INTEGER FUNCTION NF_OPEN(CHARACTER*(*) PATH, INTEGER OMODE, INTEGER ncid)

PATH File name for netCDF dataset to be opened. This may be an OPeNDAP URL if DAP support is enabled.

OMODE A zero value (or NF NOWRITE) specifies the default behavior: open the dataset with read-only access, buffering and caching accesses for efficiency. Otherwise, the creation mode is NF WRITE, NF SHARE, or OR(NF WRITE, NF SHARE). Setting the NF WRITE flag opens the dataset with read-write access. ("Writing" means any kind of change to the dataset, including appending or changing data, adding or renaming dimensions, variables, and attributes, or deleting attributes.) The NF SHARE flag is appropriate when one process may be writing the dataset and one or more other processes reading the dataset concurrently; it means that dataset accesses are not buffered and caching is limited. Since the buffering scheme is optimized for sequential access, programs that do not access data sequentially may see some performance improvement by setting the NF SHARE flag.

ncid Returned netCDF ID.

Errors

NF OPEN returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The specified netCDF dataset does not exist.

• A meaningless mode was specified.

Example

Here is an example using NF OPEN to open an existing netCDF dataset named foo.nc for read-only, non-shared access:

INCLUDE 'netcdf.inc'
...
INTEGER NCID, STATUS
...
STATUS = NF_OPEN('foo.nc', 0, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

2.9 NF OPEN

The function NF) OPEN opens an existing netCDF dataset for access, with a performance tuning parameter.

Usage

INTEGER FUNCTION NF__OPEN(CHARACTER*(*) PATH, INTEGER OMODE, INTEGER
CHUNKSIZEHINT, INTEGER ncid)

PATH File name for netCDF dataset to be opened.

OMODE A zero value (or NF NOWRITE) specifies the default behavior: open the dataset with read-only access, buffering and caching accesses for efficiency Otherwise, the creation mode is NF WRITE, NF SHARE, or OR(NF WRITE,NF SHARE). Setting the NF WRITE flag opens the dataset with read-write access. ("Writing" means any kind of change to the dataset, including appending or changing data, adding or renaming dimensions, variables, and attributes, or deleting attributes.) The NF SHARE flag is appropriate when one process may be writing the dataset and one or more other processes reading the dataset concurrently; it means that dataset accesses are not buffered and caching is limited. Since the buffering scheme is optimized for sequential access, programs that do not access data sequentially may see some performance improvement by setting the NF SHARE flag.

CHUNKSIZEHINT

This argument controls a space versus time tradeoff, memory allocated in the netcdf library versus number of system calls. Because of internal requirements, the value may not be set to exactly the value requested. The actual value chosen is returned by reference. Using the value NF SIZEHINT DEFAULT causes the library to choose a default. How the system chooses the default depends on the system. On many systems, the "preferred I/O block size" is available from the stat() system call, struct stat member st blksize. If this is available it is used. Lacking that, twice the system pagesize is used. Lacking a call to discover the system pagesize, we just set default chunksize to 8192. The chunksize is a property of a given open netcdf descriptor ncid, it is not a persistent property of the netcdf dataset.

ncid Returned netCDF ID.

Errors

NF OPEN returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The specified netCDF dataset does not exist.

• A meaningless mode was specified.

Example

Here is an example using NF OPEN to open an existing netCDF dataset named foo.nc for read-only, non-shared access:

INCLUDE 'netcdf.inc'
...
INTEGER NCID, STATUS, CHUNKSIZEHINT
...
CHUNKSIZEHINT = 1024
STATUS = NF_OPEN('foo.nc', 0, CHUNKSIZEHINT, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

2.10 NF OPEN PAR

This function opens a netCDF-4 dataset for parallel access.

This function is only available if the netCDF library was built with a HDF5 library for which –enable-parallel was used, and which was linked (like HDF5) to MPI libraries.

This opens the file using either MPI-IO or MPI-POSIX. The file must be a netCDF-4 file. (That is, it must have been created using NF NETCDF4 in the creation mode).

This function is only available if netCDF-4 was build with a version of the HDF5 library which was built with –enable-parallel.

Before either HDF5 or netCDF-4 can be installed with support for parallel programming, and MPI layer must also be installed on the machine, and usually a parallel file system.

NetCDF-4 exposes the parallel access functionality of HDF5. For more information about what is required to install and use the parallel access functions, see the HDF5 web site.

When a netCDF-4 file is opened for parallel access, collective operations are the default. To use independent access on a variable, See Section 6.29 [NF VAR PAR ACCESS], page 108.

Usage

     INTEGER FUNCTION NF_OPEN_PAR(CHARACTER*(*) PATH, INTEGER OMODE,
                                  INTEGER MPI_COMM, INTEGER MPI_INFO,
                                  INTEGER ncid)

PATH File name for netCDF dataset to be opened.

OMODE A zero value (or NF NOWRITE) specifies the default behavior: open the dataset with read-only access. Otherwise, the mode may be NF WRITE. Setting the NF WRITE flag opens the dataset with read-write access. ("Writing" means any kind of change to the dataset, including appending or changing data, adding or renaming dimensions, variables, and attributes, or deleting attributes.) Setting NF NETCDF4 is not necessary (or allowed). The file type is detected automatically.

MPI_COMM The MPI communicator.

MPI_INFO The MPI info.

ncid Returned netCDF ID.

Errors

NF OPEN returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The specified netCDF dataset does not exist.

• A meaningless mode was specified.

• Not a netCDF-4 file.

Example

This example is from the test program nf test/ftst parallel.F.

     !     Reopen the file.
           retval = nf_open_par(FILE_NAME, nf_nowrite, MPI_COMM_WORLD,
          $     MPI_INFO_NULL, ncid)
           if (retval .ne. nf_noerr) stop 2

2.11 NF REDEF

The function NF REDEF puts an open netCDF dataset into define mode, so dimensions, variables, and attributes can be added or renamed and attributes can be deleted.

Usage

     INTEGER FUNCTION NF_REDEF(INTEGER NCID)

NCID netCDF ID, from a previous call to NF OPEN or NF CREATE.

Errors

NF REDEF returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The specified netCDF dataset is already in define mode.

• The specified netCDF dataset was opened for read-only.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF REDEF to open an existing netCDF dataset named foo.nc and put it into define mode:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER NCID, STATUS
        ...
     STATUS = NF_OPEN('foo.nc', NF_WRITE, NCID)   ! open dataset
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     STATUS = NF_REDEF(NCID)                      ! put in define mode
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

2.12 NF ENDDEF

The function NF ENDDEF takes an open netCDF dataset out of define mode. The changes made to the netCDF dataset while it was in define mode are checked and committed to disk if no problems occurred. Non-record variables may be initialized to a "fill value" as well (see Section 2.18 [NF SET FILL], page 24). The netCDF dataset is then placed in data mode, so variable data can be read or written.

This call may involve copying data under some circumstances. See Section “File Structure and Performance” in NetCDF Users’ Guide.

Usage

     INTEGER FUNCTION NF_ENDDEF(INTEGER NCID)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

Errors

NF ENDDEF returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The specified netCDF dataset is not in define mode.

• The specified netCDF ID does not refer to an open netCDF dataset. The size of one or more variables exceed the size constraints for whichever variant of the file format is in use). See Section “Large File Support” in The NetCDF Users Guide.

•

Example

Here is an example using NF ENDDEF to finish the definitions of a new netCDF dataset named foo.nc and put it into data mode:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER NCID, STATUS
        ...
     STATUS = NF_CREATE('foo.nc', NF_NOCLOBBER, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...   ! create dimensions, variables, attributes
     STATUS = NF_ENDDEF(NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

2.13 NF ENDDEF

The function NF ENDDEF takes an open netCDF dataset out of define mode. The changes made to the netCDF dataset while it was in define mode are checked and committed to disk if no problems occurred. Non-record variables may be initialized to a "fill value" as well (see Section 2.18 [NF SET FILL], page 24). The netCDF dataset is then placed in data mode, so variable data can be read or written.

This call may involve copying data under some circumstances. See Section “File Structure and Performance” in NetCDF Users’ Guide.

This function assumes specific characteristics of the netcdf version 1 and version 2 file formats. Users should use nf enddef in most circumstances. Although this function will be available in future netCDF implementations, it may not continue to have any effect on performance.

The current netcdf file format has three sections, the "header" section, the data section for fixed size variables, and the data section for variables which have an unlimited dimension (record variables).

The header begins at the beginning of the file. The index (offset) of the beginning of the other two sections is contained in the header. Typically, there is no space between the sections. This causes copying overhead to accrue if one wishes to change the size of the sections, as may happen when changing names of things, text attribute values, adding attributes or adding variables. Also, for buffered i/o, there may be advantages to aligning sections in certain ways.

The minfree parameters allow one to control costs of future calls to nf redef, nf enddef by requesting that minfree bytes be available at the end of the section.

The align parameters allow one to set the alignment of the beginning of the corresponding sections. The beginning of the section is rounded up to an index which is a multiple of the align parameter. The flag value ALIGN CHUNK tells the library to use the chunksize (see above) as the align parameter.

The file format requires mod 4 alignment, so the align parameters are silently rounded up to multiples of 4. The usual call,

     nf_enddef(ncid);

is equivalent to

     nf_enddef(ncid, 0, 4, 0, 4);

The file format does not contain a "record size" value, this is calculated from the sizes of the record variables. This unfortunate fact prevents us from providing minfree and alignment control of the "records" in a netcdf file. If you add a variable which has an unlimited dimension, the third section will always be copied with the new variable added.

Usage

     INTEGER FUNCTION NF_ENDDEF(INTEGER NCID, INTEGER H_MINFREE, INTEGER V_ALIGN,
                    INTEGER V_MINFREE, INTEGER R_ALIGN)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

H_MINFREE

Sets the pad at the end of the "header" section.

V_ALIGN Controls the alignment of the beginning of the data section for fixed size variables.

V_MINFREE

Sets the pad at the end of the data section for fixed size variables.

R_ALIGN Controls the alignment of the beginning of the data section for variables which have an unlimited dimension (record variables).

Errors

NF ENDDEF returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The specified netCDF dataset is not in define mode.

• The specified netCDF ID does not refer to an open netCDF dataset.

• The size of one or more variables exceed the size constraints for whichever variant of the file format is in use). See Section “Large File Support” in The NetCDF Users Guide.

Example

Here is an example using NF ENDDEF to finish the definitions of a new netCDF dataset named foo.nc and put it into data mode:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER NCID, STATUS, H_MINFREE, V_ALIGN, V_MINFREE, R_ALIGN
        ...
     STATUS = NF_CREATE('foo.nc', NF_NOCLOBBER, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...   ! create dimensions, variables, attributes
     H_MINFREE = 512
     V_ALIGN = 512
     V_MINFREE = 512
     R_ALIGN = 512
     STATUS = NF_ENDDEF(NCID, H_MINFREE, V_ALIGN, V_MINFREE, R_ALIGN)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

2.14 NF CLOSE

The function NF CLOSE closes an open netCDF dataset. If the dataset is in define mode, NF ENDDEF will be called before closing. (In this case, if NF ENDDEF returns an error, NF ABORT will automatically be called to restore the dataset to the consistent state before define mode was last entered.) After an open netCDF dataset is closed, its netCDF ID may be reassigned to the next netCDF dataset that is opened or created.

Usage

     INTEGER FUNCTION NF_CLOSE(INTEGER NCID)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

Errors

NF CLOSE returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• Define mode was entered and the automatic call made to NF ENDDEF failed.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF CLOSE to finish the definitions of a new netCDF dataset named foo.nc and release its netCDF ID:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER NCID, STATUS
        ...

     STATUS = NF_CREATE('foo.nc', NF_NOCLOBBER, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...   ! create dimensions, variables, attributes
     STATUS = NF_CLOSE(NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

2.15 NF INQ Family

Members of the NF INQ family of functions return information about an open netCDF dataset, given its netCDF ID. Dataset inquire functions may be called from either define mode or data mode. The first function, NF INQ, returns values for the number of dimensions, the number of variables, the number of global attributes, and the dimension ID of the dimension defined with unlimited length, if any. The other functions in the family each return just one of these items of information.

For FORTRAN, these functions include NF INQ, NF INQ NDIMS, NF INQ NVARS, NF INQ NATTS, and NF INQ UNLIMDIM. An additional function, NF INQ FORMAT, returns the (rarely needed) format version.

No I/O is performed when these functions are called, since the required information is available in memory for each open netCDF dataset.

Usage

     INTEGER FUNCTION NF_INQ          (INTEGER NCID, INTEGER ndims,
                                       INTEGER nvars,INTEGER ngatts,
                                       INTEGER unlimdimid)
     INTEGER FUNCTION NF_INQ_NDIMS    (INTEGER NCID, INTEGER ndims)
     INTEGER FUNCTION NF_INQ_NVARS    (INTEGER NCID, INTEGER nvars)
     INTEGER FUNCTION NF_INQ_NATTS    (INTEGER NCID, INTEGER ngatts)
     INTEGER FUNCTION NF_INQ_UNLIMDIM (INTEGER NCID, INTEGER unlimdimid)
     INTEGER FUNCTION NF_INQ_FORMAT   (INTEGER NCID, INTEGER format)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

ndims Returned number of dimensions defined for this netCDF dataset.

nvars Returned number of variables defined for this netCDF dataset.

ngatts Returned number of global attributes defined for this netCDF dataset.

unlimdimid

Returned ID of the unlimited dimension, if there is one for this netCDF dataset. If no unlimited length dimension has been defined, -1 is returned.

format Returned format version, one of NF FORMAT CLASSIC, NF FORMAT 64BIT, NF FORMAT NETCDF4, NF FORMAT NETCDF4 CLASSIC.

Errors

All members of the NF INQ family return the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF INQ to find out about a netCDF dataset named foo.nc:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER STATUS, NCID, NDIMS, NVARS, NGATTS, UNLIMDIMID
        ...
     STATUS = NF_OPEN('foo.nc', NF_NOWRITE, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     STATUS = NF_INQ(NCID, NDIMS, NVARS, NGATTS, UNLIMDIMID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

2.16 NF SYNC

The function NF SYNC offers a way to synchronize the disk copy of a netCDF dataset with in-memory buffers. There are two reasons you might want to synchronize after writes:

• To minimize data loss in case of abnormal termination, or

• To make data available to other processes for reading immediately after it is written. But note that a process that already had the dataset open for reading would not see the number of records increase when the writing process calls NF SYNC; to accomplish this, the reading process must call NF SYNC.

This function is backward-compatible with previous versions of the netCDF library. The intent was to allow sharing of a netCDF dataset among multiple readers and one writer, by having the writer call NF SYNC after writing and the readers call NF SYNC before each read. For a writer, this flushes buffers to disk. For a reader, it makes sure that the next read will be from disk rather than from previously cached buffers, so that the reader will see changes made by the writing process (e.g., the number of records written) without having to close and reopen the dataset. If you are only accessing a small amount of data, it can be expensive in computer resources to always synchronize to disk after every write, since you are giving up the benefits of buffering.

An easier way to accomplish sharing (and what is now recommended) is to have the writer and readers open the dataset with the NF SHARE flag, and then it will not be necessary to call NF SYNC at all. However, the NF SYNC function still provides finer granularity than the NF SHARE flag, if only a few netCDF accesses need to be synchronized among processes.

It is important to note that changes to the ancillary data, such as attribute values, are not propagated automatically by use of the NF SHARE flag. Use of the NF SYNC function is still required for this purpose.

Sharing datasets when the writer enters define mode to change the data schema requires extra care. In previous releases, after the writer left define mode, the readers were left looking at an old copy of the dataset, since the changes were made to a new copy. The only way readers could see the changes was by closing and reopening the dataset. Now the changes are made in place, but readers have no knowledge that their internal tables are now inconsistent with the new dataset schema. If netCDF datasets are shared across redefinition, some mechanism external to the netCDF library must be provided that prevents access by readers during redefinition and causes the readers to call NF SYNC before any subsequent access.

When calling NF SYNC, the netCDF dataset must be in data mode. A netCDF dataset in define mode is synchronized to disk only when NF ENDDEF is called. A process that is reading a netCDF dataset that another process is writing may call NF SYNC to get updated with the changes made to the data by the writing process (e.g., the number of records written), without having to close and reopen the dataset.

Data is automatically synchronized to disk when a netCDF dataset is closed, or whenever you leave define mode.

Usage

     INTEGER FUNCTION NF_SYNC(INTEGER NCID)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

Errors

NF SYNC returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The netCDF dataset is in define mode.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF SYNC to synchronize the disk writes of a netCDF dataset named foo.nc:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER STATUS, NCID
        ...
     STATUS = NF_OPEN('foo.nc', NF_WRITE, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     ! write data or change attributes
        ...
     STATUS = NF_SYNC(NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

2.17 NF ABORT

You no longer need to call this function, since it is called automatically by NF CLOSE in case the dataset is in define mode and something goes wrong with committing the changes.

The function NF ABORT just closes the netCDF dataset, if not in define mode. If the dataset is being created and is still in define mode, the dataset is deleted. If define mode was entered by a call to NF REDEF, the netCDF dataset is restored to its state before definition mode was entered and the dataset is closed.

Usage

     INTEGER FUNCTION NF_ABORT(INTEGER NCID)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

Errors

NF ABORT returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• When called from define mode while creating a netCDF dataset, deletion of the dataset failed.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF ABORT to back out of redefinitions of a dataset named foo.nc:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER STATUS, NCID, LATID
        ...
     STATUS = NF_OPEN('foo.nc', NF_WRITE, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     STATUS = NF_REDEF(NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     STATUS = NF_DEF_DIM(NCID, 'LAT', 18, LATID)
     IF (STATUS .NE. NF_NOERR) THEN  ! dimension definition failed
        CALL HANDLE_ERR(STATUS)
        STATUS = NF_ABORT(NCID)  ! abort redefinitions
        IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
     ENDIF
     ...

2.18 NF SET FILL

This function is intended for advanced usage, to optimize writes under some circumstances described below. The function NF SET FILL sets the fill mode for a netCDF dataset open for writing and returns the current fill mode in a return parameter. The fill mode can be specified as either NF FILL or NF NOFILL. The default behavior corresponding to NF FILL is that data is pre-filled with fill values, that is fill values are written when you create non-record variables or when you write a value beyond data that has not yet been written. This makes it possible to detect attempts to read data before it was written. See Section 6.27 [Fill Values], page 106, for more information on the use of fill values. See Section 7.2 [Attribute Conventions], page 111, for information about how to define your own fill values.

The behavior corresponding to NF NOFILL overrides the default behavior of prefilling data with fill values. This can be used to enhance performance, because it avoids the duplicate writes that occur when the netCDF library writes fill values that are later overwritten with data.

A value indicating which mode the netCDF dataset was already in is returned. You can use this value to temporarily change the fill mode of an open netCDF dataset and then restore it to the previous mode.

After you turn on NF NOFILL mode for an open netCDF dataset, you must be certain to write valid data in all the positions that will later be read. Note that nofill mode is only a transient property of a netCDF dataset open for writing: if you close and reopen the dataset, it will revert to the default behavior. You can also revert to the default behavior by calling NF SET FILL again to explicitly set the fill mode to NF FILL.

There are three situations where it is advantageous to set nofill mode:

1. Creating and initializing a netCDF dataset. In this case, you should set nofill mode before calling NF ENDDEF and then write completely all non-record variables and the initial records of all the record variables you want to initialize.

2. Extending an existing record-oriented netCDF dataset. Set nofill mode after opening the dataset for writing, then append the additional records to the dataset completely, leaving no intervening unwritten records.

3. Adding new variables that you are going to initialize to an existing netCDF dataset. Set nofill mode before calling NF ENDDEF then write all the new variables completely.

If the netCDF dataset has an unlimited dimension and the last record was written while in nofill mode, then the dataset may be shorter than if nofill mode was not set, but this will be completely transparent if you access the data only through the netCDF interfaces.

The use of this feature may not be available (or even needed) in future releases. Programmers are cautioned against heavy reliance upon this feature.

Usage

     INTEGER FUNCTION NF_SET_FILL(INTEGER NCID, INTEGER FILLMODE,
                                  INTEGER old_mode)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

FILLMODE Desired fill mode for the dataset, either NF NOFILL or NF FILL.

old_mode Returned current fill mode of the dataset before this call, either NF NOFILL or NF FILL.

Errors

NF SET FILL returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The specified netCDF ID does not refer to an open netCDF dataset.

• The specified netCDF ID refers to a dataset open for read-only access.

• The fill mode argument is neither NF NOFILL nor NF FILL..

Example

Here is an example using NF SET FILL to set nofill mode for subsequent writes of a netCDF dataset named foo.nc:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER NCID, STATUS, OMODE
        ...
     STATUS = NF_OPEN('foo.nc', NF_WRITE, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     ! write data with default prefilling behavior
        ...
     STATUS = NF_SET_FILL(NCID, NF_NOFILL, OMODE)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     ! write data with no prefilling
        ...

2.19 NF SET DEFAULT FORMAT

This function is intended for advanced users.

In version 3.6, netCDF introduced a new data format, the first change in the underlying binary data format since the netCDF interface was released. The new format, 64-bit offset format, was introduced to greatly relax the limitations on creating very large files.

In version 4.0, another new binary format was introduced: netCDF-4/HDF5.

Users are warned that creating files in the 64-bit offset format makes them unreadable by the netCDF library prior to version 3.6.0, and creating files in netcdf-4/HDF5 format makes them unreadable by the netCDF library prior to version 4.0. For reasons of compatibility, users should continue to create files in netCDF classic format.

Users who do want to use 64-bit offset or netCDF-4/HDF5 format files can create them directory from NF CREATE, using the proper cmode flag. (see Section 2.5 [NF CREATE], page 9).

The function NF SET DEFAULT FORMAT allows the user to change the format of the netCDF file to be created by future calls to NF CREATE without changing the cmode flag.

This allows the user to convert a program to use the new formats without changing all calls the NF CREATE.

Once the default format is set, all future created files will be in the desired format. Constants are provided in the netcdf.inc file to be used with this function:

nf format classic, nf format 64bit, nf format netcdf4 and nf format netcdf4 classic.

Usage

INTEGER FUNCTION NF_SET_DEFAULT_FORMAT(INTEGER FORMAT, INTEGER OLD_FORMT)

FORMAT  Either nf format classic,  nf format 64bit,  nf format netcdf4 or

nf format netcdf4 classic.

OLD_FORMAT

The default format at the time the function is called is returned here.

Errors

The following error codes may be returned by this function:

• An NF EINVAL error is returned if an invalid default format is specified.

Example

INCLUDE 'netcdf.inc'
   ...
INTEGER STATUS, OLD_FORMAT
   ...
STATUS = NF_SET_DEFAULT_FORMAT(nf_format_64bit, OLD_FORMAT)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...

3 Groups

NetCDF-4 added support for hierarchical groups within netCDF datasets.

Groups are identified with a ncid, which identifies both the open file, and the group within that file. When a file is opened with NF OPEN or NF CREATE, the ncid for the root group of that file is provided. Using that as a starting point, users can add new groups, or list and navigate existing groups.

All netCDF calls take a ncid which determines where the call will take its action. For example, the NF DEF VAR function takes a ncid as its first parameter. It will create a variable in whichever group its ncid refers to. Use the root ncid provided by NF CREATE or NF OPEN to create a variable in the root group. Or use NF DEF GRP to create a group and use its ncid to define a variable in the new group.

Variable are only visible in the group in which they are defined. The same applies to attributes. “Global” attributes are defined in whichever group is refered to by the ncid.

Dimensions are visible in their groups, and all child groups.

Group operations are only permitted on netCDF-4 files - that is, files created with the HDF5 flag in nf create. (see Section 2.5 [NF CREATE], page 9). Groups are not compatible with the netCDF classic data model, so files created with the NF CLASSIC MODEL file cannot contain groups (except the root group).

3.1 Find a Group ID: NF INQ NCID

Given an ncid and group name (NULL or "" gets root group), return ncid of the named group.

Usage

     INTEGER FUNCTION NF_INQ_NCID(INTEGER NCID, CHARACTER*(*) NAME, INTEGER GRPID)

NCID The group id for this operation.

NAME A character array that holds the name of the desired group. Must be less then NF MAX NAME.

GRPID The ID of the group will go here.

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag HDF5. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/ftst groups.F.

     C     Check getting the group by name
           retval = nf_inq_ncid(ncid, group_name, grpid_in)
           if (retval .ne. nf_noerr) call handle_err(retval)

3.2 Get a List of Groups in a Group: NF INQ GRPS

Given a location id, return the number of groups it contains, and an array of their ncids.

Usage

     INTEGER FUNCTION NF_INQ_GRPS(INTEGER NCID, INTEGER NUMGRPS, INTEGER NCIDS)

NCID The group id for this operation.

NUMGRPS An integer which will get number of groups in this group.

NCIDS An array of ints which will receive the IDs of all the groups in this group.

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag HDF5. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/ftst groups.F.

     C     What groups are there from the root group?
           retval = nf_inq_grps(ncid, ngroups_in, grpids)
           if (retval .ne. nf_noerr) call handle_err(retval)

3.3 Find all the Variables in a Group: NF INQ VARIDS

Find all varids for a location.

Usage

     INTEGER FUNCTION NF_INQ_VARIDS(INTEGER NCID, INTEGERS VARIDS)

NCID The group id for this operation.

VARIDS An already allocated array to store the list of varids. Use nf inq nvars to find out how many variables there are. (see Section 2.15 [NF INQ Family], page 21).

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag HDF5. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/ftst groups.F.

     C     Check varids in subgroup.
           retval = nf_inq_varids(subgrp_in, nvars, varids_in)
           if (retval .ne. nf_noerr) call handle_err(retval)

3.4 Find all Dimensions Visible in a Group:

NF INQ DIMIDS

Find all dimids for a location. This finds all dimensions in a group, or any of its parents.

Usage

     INTEGER FUNCTION NF_INQ_DIMIDS(INTEGER NCID, INTEGER DIMIDS, INTEGER INCLUDE_PARENTS)

NCID The group id for this operation.

DIMIDS An array of ints when the dimids of the visible dimensions will be stashed. Use nf inq ndims to find out how many dims are visible from this group. (see Section 2.15 [NF INQ Family], page 21).

INCLUDE_PARENTS

If zero, only the group specified by NCID will be searched for dimensions. Otherwise parent groups will be searched too.

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag HDF5. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/ftst groups.F.

     C     Check dimids in subgroup.
           retval = nf_inq_dimids(subgrp_in, ndims, dimids_in, 0)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (ndims .ne. 2 .or. dimids_in(1) .ne. dimids(1) .or.
          &     dimids_in(2) .ne. dimids(2)) stop 2

3.5 Find the Length of a Group’s Name:

NF INQ GRPNAME LEN

Given ncid, find length of the full name. (Root group is named "/", with length 1.)

Usage

     INTEGER FUNCTION NF_INQ_GRPNAME_LEN(INTEGER NCID, INTEGER LEN)

NCID The group id for this operation.

LEN An integer where the length will be placed.

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag HDF5. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/ftst groups.F.

     C     Check the length of the full name.
           retval = nf_inq_grpname_len(grpids(1), full_name_len)
           if (retval .ne. nf_noerr) call handle_err(retval)

3.6 Find a Group’s Name: NF INQ GRPNAME

Given ncid, find relative name of group. (Root group is named "/").

The name provided by this function is relative to the parent group. For a full path name for the group is, with all parent groups included, separated with a forward slash (as in Unix directory names) See Section 3.7 [NF INQ GRPNAME FULL], page 34.

Usage

     INTEGER FUNCTION NF_INQ_GRPNAME(INTEGER NCID, CHARACTER*(*) NAME)

NCID The group id for this operation.

NAME The name of the group will be copied to this character array. The name will be less than NF MAX NAME in length.

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag HDF5. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/ftst groups.F.

     C     Check the name of the root group.
           retval = nf_inq_grpname(ncid, name_in)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (name_in(1:1) .ne. '/') stop 2

3.7 Find a Group’s Full Name: NF INQ GRPNAME FULL

Given ncid, find complete name of group. (Root group is named "/").

The name provided by this function is a full path name for the group is, with all parent groups included, separated with a forward slash (as in Unix directory names). For a name relative to the parent group See Section 3.6 [NF INQ GRPNAME], page 33.

To find the length of the full name See Section 3.5 [NF INQ GRPNAME LEN], page 32.

Usage

     INTEGER FUNCTION NF_INQ_GRPNAME_FULL(INTEGER NCID, INTEGER LEN, CHARACTER*(*) NAME)

NCID The group id for this operation.

LEN The length of the full group name will go here.

NAME The name of the group will be copied to this character array.

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag HDF5. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/ftst groups.F.

     C     Check the full name.
           retval = nf_inq_grpname_full(grpids(1), full_name_len, name_in2)
           if (retval .ne. nf_noerr) call handle_err(retval)

3.8 Find a Group’s Parent: NF INQ GRP PARENT

Given ncid, find the ncid of the parent group.

When used with the root group, this function returns the NF ENOGRP error (since the root group has no parent.)

Usage

     INTEGER FUNCTION NF_INQ_GRP_PARENT(INTEGER NCID, INTEGER PARENT_NCID)

NCID The group id.

PARENT_NCID

The ncid of the parent group will be copied here.

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENOGRP

No parent group found (i.e. this is the root group).

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag HDF5. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/ftst groups.F.

     C     Check the parent ncid.
           retval = nf_inq_grp_parent(grpids(1), grpid_in)
           if (retval .ne. nf_noerr) call handle_err(retval)

3.9 Create a New Group: NF DEF GRP

Create a group. Its location id is returned in new ncid.

Usage

     INTEGER FUNCTION NF_DEF_GRP(INTEGER PARENT_NCID, CHARACTER*(*) NAME,
             INTEGER NEW_NCID)

PARENT_NCID

The group id of the parent group.

NAME The name of the new group.

NEW_NCID The ncid of the new group will be placed there.

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENAMEINUSE

That name is in use. Group names must be unique within a group.

NF_EMAXNAME

Name exceed max length NF MAX NAME.

NF_EBADNAME

Name contains illegal characters.

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag HDF5. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

NF_EPERM Attempt to write to a read-only file.

NF_ENOTINDEFINE

Not in define mode.

Example

In this exampe rom nf test/ftst groups.F, a groups is reated, and then a sub-group is created in that group.

     C     Create the netCDF file.
           retval = nf_create(file_name, NF_NETCDF4, ncid)
           if (retval .ne. nf_noerr) call handle_err(retval)
     C     Create a group and a subgroup.
           retval = nf_def_grp(ncid, group_name, grpid)
           if (retval .ne. nf_noerr) call handle_err(retval)
           retval = nf_def_grp(grpid, sub_group_name, sub_grpid)
           if (retval .ne. nf_noerr) call handle_err(retval)

4 Dimensions

4.1 Dimensions Introduction

Dimensions for a netCDF dataset are defined when it is created, while the netCDF dataset is in define mode. Additional dimensions may be added later by reentering define mode. A netCDF dimension has a name and a length. At most one dimension in a netCDF dataset can have the unlimited length, which means variables using this dimension can grow along this dimension.

There is a suggested limit (100) to the number of dimensions that can be defined in a single netCDF dataset. The limit is the value of the predefined macro NF MAX DIMS. The purpose of the limit is to make writing generic applications simpler. They need only provide an array of NF MAX DIMS dimensions to handle any netCDF dataset. The implementation of the netCDF library does not enforce this advisory maximum, so it is possible to use more dimensions, if necessary, but netCDF utilities that assume the advisory maximums may not be able to handle the resulting netCDF datasets.

Ordinarily, the name and length of a dimension are fixed when the dimension is first defined. The name may be changed later, but the length of a dimension (other than the unlimited dimension) cannot be changed without copying all the data to a new netCDF dataset with a redefined dimension length.

A netCDF dimension in an open netCDF dataset is referred to by a small integer called a dimension ID. In the FORTRAN interface, dimension IDs are 1, 2, 3, ..., in the order in which the dimensions were defined.

Operations supported on dimensions are:

• Create a dimension, given its name and length.

• Get a dimension ID from its name.

• Get a dimension’s name and length from its ID.

• Rename a dimension.

4.2 NF DEF DIM

The function NF DEF DIM adds a new dimension to an open netCDF dataset in define mode. It returns (as an argument) a dimension ID, given the netCDF ID, the dimension name, and the dimension length. At most one unlimited length dimension, called the record dimension, may be defined for each netCDF dataset.

Usage

     INTEGER FUNCTION NF_DEF_DIM (INTEGER NCID, CHARACTER*(*) NAME,
                               INTEGER LEN, INTEGER dimid)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

NAME Dimension name.

LEN Length of dimension; that is, number of values for this dimension as an index to variables that use it. This should be either a positive integer or the predefined constant NF UNLIMITED. dimid Returned dimension ID.

Errors

NF DEF DIM returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The netCDF dataset is not in definition mode.

• The specified dimension name is the name of another existing dimension.

• The specified length is not greater than zero.

• The specified length is unlimited, but there is already an unlimited length dimension defined for this netCDF dataset.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF DEF DIM to create a dimension named lat of length 18 and a unlimited dimension named rec in a new netCDF dataset named foo.nc:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER STATUS, NCID, LATID, RECID
        ...
     STATUS = NF_CREATE('foo.nc', NF_NOCLOBBER, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     STATUS = NF_DEF_DIM(NCID, 'lat', 18, LATID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
     STATUS = NF_DEF_DIM(NCID, 'rec', NF_UNLIMITED, RECID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

4.3 NF INQ DIMID

The function NF INQ DIMID returns (as an argument) the ID of a netCDF dimension, given the name of the dimension. If ndims is the number of dimensions defined for a netCDF dataset, each dimension has an ID between 1 and ndims.

Usage

     INTEGER FUNCTION NF_INQ_DIMID (INTEGER NCID, CHARACTER*(*) NAME,
                                    INTEGER dimid)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

NAME Dimension name.

dimid Returned dimension ID.

Errors

NF INQ DIMID returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The name that was specified is not the name of a dimension in the netCDF dataset.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF INQ DIMID to determine the dimension ID of a dimension named lat, assumed to have been defined previously in an existing netCDF dataset named foo.nc:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER STATUS, NCID, LATID
        ...
     STATUS = NF_OPEN('foo.nc', NF_NOWRITE, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     STATUS = NF_INQ_DIMID(NCID, 'lat', LATID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

4.4 NF INQ DIM Family

This family of functions returns information about a netCDF dimension. Information about a dimension includes its name and its length. The length for the unlimited dimension, if any, is the number of records written so far.

The functions in this family include NF INQ DIM, NF INQ DIMNAME, and NF INQ DIMLEN. The function NF INQ DIM returns all the information about a dimension; the other functions each return just one item of information.

Usage

     INTEGER FUNCTION NF_INQ_DIM     (INTEGER NCID, INTEGER DIMID,
                                      CHARACTER*(*) name, INTEGER len)
     INTEGER FUNCTION NF_INQ_DIMNAME (INTEGER NCID, INTEGER DIMID,
                                      CHARACTER*(*) name)
     INTEGER FUNCTION NF_INQ_DIMLEN  (INTEGER NCID, INTEGER DIMID,
                                      INTEGER len)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

DIMID Dimension ID, from a previous call to NF INQ DIMID or NF DEF DIM.

NAME Returned dimension name. The caller must allocate space for the returned name. The maximum possible length, in characters, of a dimension name is given by the predefined constant NF MAX NAME.

len Returned length of dimension. For the unlimited dimension, this is the current maximum value used for writing any variables with this dimension, that is the maximum record number.

Errors

These functions return the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The dimension ID is invalid for the specified netCDF dataset.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF INQ DIM to determine the length of a dimension named lat, and the name and current maximum length of the unlimited dimension for an existing netCDF dataset named foo.nc:

INCLUDE 'netcdf.inc'
   ...
INTEGER STATUS, NCID, LATID, LATLEN, RECID, NRECS
CHARACTER*(NF_MAX_NAME) LATNAM, RECNAM
   ...
STATUS = NF_OPEN('foo.nc', NF_NOWRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
! get ID of unlimited dimension
STATUS = NF_INQ_UNLIMDIM(NCID, RECID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_DIMID(NCID, 'lat', LATID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
! get lat length
STATUS = NF_INQ_DIMLEN(NCID, LATID, LATLEN)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
! get unlimited dimension name and current length
STATUS = NF_INQ_DIM(NCID, RECID, RECNAME, NRECS)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

4.5 NF RENAME DIM

The function NF RENAME DIM renames an existing dimension in a netCDF dataset open for writing. If the new name is longer than the old name, the netCDF dataset must be in define mode. You cannot rename a dimension to have the same name as another dimension.

Usage

INTEGER FUNCTION NF_RENAME_DIM (INTEGER NCID, INTEGER DIMID,
                                CHARACTER*(*) NAME)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

DIMID Dimension ID, from a previous call to NF INQ DIMID or NF DEF DIM.

NAME New dimension name.

Errors

NF RENAME DIM returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The new name is the name of another dimension.

• The dimension ID is invalid for the specified netCDF dataset.

• The specified netCDF ID does not refer to an open netCDF dataset.

• The new name is longer than the old name and the netCDF dataset is not in define mode.

Example

Here is an example using NF RENAME DIM to rename the dimension lat to latitude in an existing netCDF dataset named foo.nc:

INCLUDE 'netcdf.inc'
   ...
INTEGER STATUS, NCID, LATID
   ...
STATUS = NF_OPEN('foo.nc', NF_WRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
! put in define mode to rename dimension
STATUS = NF_REDEF(NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_INQ_DIMID(NCID, 'lat', LATID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_RENAME_DIM(NCID, LATID, 'latitude')
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
! leave define mode
STATUS = NF_ENDDEF(NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

5 User Defined Data Types

5.1 User Defined Types Introduction

NetCDF-4 has added support for four different user defined data types.

compound type

Like a C struct, a compound type is a collection of types, including other user defined types, in one package.

variable length array type

The variable length array may be used to store ragged arrays.

opaque type

This type has only a size per element, and no other type information.

enum type Like an enumeration in C, this type lets you assign text values to integer values, and store the integer values.

Users may construct user defined type with the various NF DEF * functions described in this section. They may learn about user defined types by using the NF INQ functions defined in this section.

Once types are constructed, define variables of the new type with NF DEF VAR (see Section 6.3 [NF DEF VAR], page 66). Write to them with NF PUT VAR1, NF PUT VAR, NF PUT VARA, or NF PUT VARS (see Chapter 6 [Variables], page 65). Read data of user-defined type with NF GET VAR1, NF GET VAR, NF GET VARA, or NF GET VARS (see Chapter 6 [Variables], page 65).

Create attributes of the new type with NF PUT ATT (see Section 7.3 [NF PUT ATT type], page 114). Read attributes of the new type with NF GET ATT (see Section 7.5 [NF GET ATT type], page 118).

5.2 Learn the IDs of All Types in Group:

NF INQ TYPEIDS

Learn the number of types defined in a group, and their IDs.

Usage

     INTEGER FUNCTION NF_INQ_TYPEIDS(INTEGER NCID, INTEGER NTYPES,
             INTEGER TYPEIDS)

NCID The group id.

NTYPES A pointer to int which will get the number of types defined in the group. If NULL, ignored.

TYPEIDS A pointer to an int array which will get the typeids. If NULL, ignored.

Errors

NF_NOERR No error.
NF_BADID Bad ncid.

Example

The following example is from the test program nf test/ftst vars3.F.

           retval = nf_inq_typeids(ncid, num_types, typeids)
           if (retval .ne. nf_noerr) call handle_err(retval)

5.3 Learn About a User Defined Type: NF INQ TYPE

Given an ncid and a typeid, get the information about a type. This function will work on any type, including atomic and any user defined type, whether compound, opaque, enumeration, or variable length array.

For even more information about a user defined type Section 5.4 [NF INQ USER TYPE], page 45.

Usage

     INTEGER FUNCTION NF_INQ_TYPE(INTEGER NCID, INTEGER XTYPE, CHARACTER*(*) NAME, INTEGER SIZE)

NCID The ncid for the group containing the type (ignored for atomic types).

XTYPE The typeid for this type, as returned by NF DEF COMPOUND, NF DEF OPAQUE, NF DEF ENUM, NF DEF VLEN, or NF INQ VAR, or as found in netcdf.inc in the list of atomic types (NF CHAR, NF INT, etc.).

NAME The name of the user defined type will be copied here. It will be NF MAX NAME bytes or less. For atomic types, the type name from CDL will be given.

SIZEP The size of the type (in bytes) will be copied here. VLEN type size is the size of one element of the VLEN. String size is returned as zero, since it varies from string to string.

Return Codes

NF_NOERR No error.
NF_EBADTYPEID

Bad typeid.

NF_ENOTNC4

Seeking a user-defined type in a netCDF-3 file.

NF_ESTRICTNC3

Seeking a user-defined type in a netCDF-4 file for which classic model has been turned on.

NF_EBADGRPID

Bad group ID in ncid.

NF_EBADID

Type ID not found.

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from the test program nf test/ftst vars3.F, and it uses all the possible inquiry functions on an enum type.

     C     Check the enum type.
           retval = NF_INQ_TYPEIDS(ncid, num_types, typeids)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (num_types .ne. MAX_TYPES) stop 2
           retval = nf_inq_enum(ncid, typeids(1), type_name, base_type,
          &     base_size, num_members)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (base_type .ne. NF_INT .or. num_members .ne. 2) stop 2
           retval = nf_inq_enum_member(ncid, typeids(1), 1, member_name,
          &     member_value)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (member_name(1:len(one_name)) .ne. one_name) stop 2

5.4 Learn About a User Defined Type:

NF INQ USER TYPE

Given an ncid and a typeid, get the information about a user defined type. This function will work on any user defined type, whether compound, opaque, enumeration, or variable length array.

Usage

     INTEGER FUNCTION NF_INQ_USER_TYPE(INTEGER NCID, INTEGER XTYPE,
             CHARACTER*(*) NAME, INTEGER SIZE, INTEGER BASE_NF_TYPE,
             INTEGER NFIELDS, INTEGER CLASS)

NCID The ncid for the group containing the user defined type.

XTYPE The typeid for this type, as returned by NF DEF COMPOUND, NF DEF OPAQUE, NF DEF ENUM, NF DEF VLEN, or NF INQ VAR.

NAME The name of the user defined type will be copied here. It will be NF MAX NAME bytes or less.

SIZE The size of the user defined type will be copied here.

BASE_NF_TYPE

The base typeid will be copied here for vlen and enum types.

NFIELDS The number of fields will be copied here for enum and compound types.

CLASS The class of the user defined type, NF VLEN, NF OPAQUE, NF ENUM, or NF COMPOUND, will be copied here.

Errors

NF_NOERR No error.

NF_EBADTYPEID

Bad typeid.

NF_EBADFIELDID

Bad fieldid.

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/ftst types2.F.

     C     Check the type.
           retval = nf_inq_user_type(ncid, typeids(1), name_in, size_in,
          &     base_type_in, nfields_in, class_in)
           if (retval .ne. nf_noerr) call handle_err(retval)

5.5 Compound Types Introduction

NetCDF-4 added support for compound types, which allow users to construct a new type - a combination of other types, like a C struct.

Compound types are not supported in classic or 64-bit offset format files.

To write data in a compound type, first use nf def compound to create the type, multiple calls to nf insert compound to add to the compound type, and then write data with the appropriate nf put var1, nf put vara, nf put vars, or nf put varm call.

To read data written in a compound type, you must know its structure. Use the NF INQ COMPOUND functions to learn about the compound type.

In Fortran a character buffer must be used for the compound data. The user must read the data from within that buffer in the same way that the C compiler which compiled netCDF would store the structure.

The use of compound types introduces challenges and portability issues for Fortran users.

5.5.1 Creating a Compound Type: NF DEF COMPOUND

Create a compound type. Provide an ncid, a name, and a total size (in bytes) of one element of the completed compound type.

After calling this function, fill out the type with repeated calls to NF INSERT COMPOUND (see Section 5.5.2 [NF INSERT COMPOUND], page 47). Call NF INSERT COMPOUND once for each field you wish to insert into the compound type.

Note that there does not seem to be a way to read such types into structures in Fortran 90 (and there are no structures in Fortran 77).

Fortran users may use character buffers to read and write compound types.

Usage

     INTEGER FUNCTION NF_DEF_COMPOUND(INTEGER NCID, INTEGER SIZE,
             CHARACTER*(*) NAME, INTEGER TYPEIDP)

NCID The groupid where this compound type will be created.

SIZE The size, in bytes, of the compound type.

NAME The name of the new compound type.

TYPEIDP The typeid of the new type will be placed here.

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENAMEINUSE

That name is in use. Compound type names must be unique in the data file.

NF_EMAXNAME

Name exceeds max length NF MAX NAME.

NF_EBADNAME

Name contains illegal characters.

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag NF NETCDF4. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

NF_EPERM Attempt to write to a read-only file.

NF_ENOTINDEFINE

Not in define mode.

Example

This example is from nf test/ftst types2.F.

     C     Define a compound type.
           retval = nf_def_compound(ncid, cmp_size, type_name,
          &     cmp_typeid)
           if (retval .ne. nf_noerr) call handle_err(retval)

5.5.2 Inserting a Field into a Compound Type:

NF INSERT COMPOUND

Insert a named field into a compound type.

Usage

     INTEGER FUNTION NF_INSERT_COMPOUND(INTEGER TYPEID, CHARACTER*(*) NAME, INTEGER OFFSET,
             INTEGER FIELD_TYPEID)

TYPEID The typeid for this compound type, as returned by NF DEF COMPOUND, or NF INQ VAR.

NAME The name of the new field.

OFFSET Offset in byte from the beginning of the compound type for this field.

FIELD_TYPEID

The type of the field to be inserted.

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENAMEINUSE

That name is in use. Field names must be unique within a compound type.

NF_EMAXNAME

Name exceed max length NF MAX NAME.

NF_EBADNAME

Name contains illegal characters.

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag NF NETCDF4. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

NF_ENOTINDEFINE

Not in define mode.

Example

This example is from nf test/ftst types.F.

     C     Define a compound type.
           retval = nf_def_compound(ncid, WIND_T_SIZE, type_name,
          &     wind_typeid)
           if (retval .ne. nf_noerr) call handle_err(retval)
           retval = nf_insert_compound(ncid, wind_typeid, u_name, 0, NF_INT)
           if (retval .ne. nf_noerr) call handle_err(retval)

           retval = nf_insert_compound(ncid, wind_typeid, v_name, 4, NF_INT)
           if (retval .ne. nf_noerr) call handle_err(retval)

5.5.3 Inserting an Array Field into a Compound Type:

NF INSERT ARRAY COMPOUND

Insert a named array field into a compound type.

Usage

     INTEGER FUNCTION NF_INSERT_ARRAY_COMPOUND(INTEGER NCID, INTEGER XTYPE,
             CHARACTER*(*) NAME, INTEGER OFFSET, INTEGER FIELD_TYPEID,
             INTEGER NDIMS, INTEGER DIM_SIZES)

NCID The ID of the file that contains the array type and the compound type.

XTYPE The typeid for this compound type, as returned by nf def compound, or nf inq var.

NAME The name of the new field.

OFFSET Offset in byte from the beginning of the compound type for this field.

FIELD_TYPEID

The base type of the array to be inserted.

NDIMS The number of dimensions for the array to be inserted.

DIM_SIZES

An array containing the sizes of each dimension.

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENAMEINUSE

That name is in use. Field names must be unique within a compound type.

NF_EMAXNAME

Name exceed max length NF MAX NAME.

NF_EBADNAME

Name contains illegal characters.

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag NF NETCDF4. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

NF_ENOTINDEFINE

Not in define mode.

NF_ETYPEDEFINED

Attempt to change type that has already been committed. The first time the file leaves define mode, all defined types are committed, and can’t be changed. If you wish to add an array to a compound type, you must do so before the compound type is committed.

Example

This example is from nf test/ftst types2.F.

     C     Define a compound type.
           retval = nf_def_compound(ncid, cmp_size, type_name,
          &     cmp_typeid)
           if (retval .ne. nf_noerr) call handle_err(retval)
     C     Include an array.
           dim_sizes(1) = NX
           dim_sizes(2) = NY
           retval = nf_insert_array_compound(ncid, cmp_typeid, ary_name, 0,
          &     NF_INT, NDIMS, dim_sizes)
           if (retval .ne. nf_noerr) call handle_err(retval)

5.5.4 Learn About a Compound Type: NF INQ COMPOUND

Get the number of fields, length in bytes, and name of a compound type.

In addtion to the NF INQ COMPOUND function, three additional functions are provided which get only the name, size, and number of fields.

Usage

     INTEGER FUNCTION NF_INQ_COMPOUND(INTEGER NCID, INTEGER XTYPE,
             CHARACTER*(*) NAME, INTEGER SIZEP, INTEGER NFIELDSP)
     INTEGER FUNCTION NF_INQ_COMPOUND_NAME(INTEGER NCID, INTEGER XTYPE,
             CHARACTER*(*) NAME)
     INTEGER FUNCTION NF_INQ_COMPOUND_SIZE(INTEGER NCID, INTEGER XTYPE,
             INTEGER SIZEP)
     INTEGER FUNCTION NF_INQ_COMPOUND_NFIELDS(INTEGER NCID, INTEGER XTYPE,
             INTEGER NFIELDSP)

NCID The ID of any group in the file that contains the compound type.

XTYPE The typeid for this compound type, as returned by NF DEF COMPOUND, or NF INQ VAR.

NAME Character array which will get the name of the compound type. It will have a maximum length of NF MAX NAME.

SIZEP The size of the compound type in bytes will be put here.

NFIELDSP The number of fields in the compound type will be placed here.

Return Codes

NF_NOERR No error.
NF_EBADID

Couldn’t find this ncid.

NF_ENOTNC4

Not a netCDF-4/HDF5 file.

NF_ESTRICTNC3

A netCDF-4/HDF5 file, but with CLASSIC MODEL. No user defined types are allowed in the classic model.

NF_EBADTYPE

This type not a compound type.

NF_EBADTYPEID

Bad type id.

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/ftst types.F.

     C     Check it differently.
           retval = nf_inq_compound(ncid, typeids(1), name_in, size_in,
          &     nfields_in)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (name_in(1:len(type_name)) .ne. type_name .or.
          &     size_in .ne. WIND_T_SIZE .or. nfields_in .ne. 2) stop 2
     C     Check it one piece at a time.
           retval = nf_inq_compound_nfields(ncid, typeids(1), nfields_in)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (nfields_in .ne. 2) stop 2
           retval = nf_inq_compound_size(ncid, typeids(1), size_in)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (size_in .ne. WIND_T_SIZE) stop 2
           retval = nf_inq_compound_name(ncid, typeids(1), name_in)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (name_in(1:len(type_name)) .ne. type_name) stop 2

5.5.5 Learn About a Field of a Compound Type:

NF INQ COMPOUND FIELD

Get information about one of the fields of a compound type.

Usage

     INTEGER FUNCTION NF_INQ_COMPOUND_FIELD(INTEGER NCID, INTEGER XTYPE,
             INTEGER FIELDID, CHARACTER*(*) NAME, INTEGER OFFSETP,
             INTEGER FIELD_TYPEIDP, INTEGER NDIMSP, INTEGER DIM_SIZESP)
     INTEGER FUNCTION NF_INQ_COMPOUND_FIELDNAME(INTEGER TYPEID,
             INTEGER FIELDID, CHARACTER*(*) NAME)
     INTEGER FUNCTION NF_INQ_COMPOUND_FIELDINDEX(INTEGER TYPEID,
             CHARACTER*(*) NAME, INTEGER FIELDIDP)
     INTEGER FUNCTION NF_INQ_COMPOUND_FIELDOFFSET(INTEGER TYPEID,
             INTEGER FIELDID, INTEGER OFFSETP)
     INTEGER FUNCTION NF_INQ_COMPOUND_FIELDTYPE(INTEGER TYPEID,
             INTEGER FIELDID, INTEGER FIELD_TYPEIDP)
     INTEGER FUNCTION NF_INQ_COMPOUND_FIELDNDIMS(INTEGER NCID,
             INTEGER XTYPE, INTEGER FIELDID, INTEGER NDIMSP)
     INTEGER FUNCTION NF_INQ_COMPOUND_FIELDDIM_SIZES(INTEGER NCID,
             INTEGER XTYPE, INTEGER FIELDID, INTEGER DIM_SIZES)

NCID The groupid where this compound type exists.

XTYPE The typeid for this compound type, as returned by NF DEF COMPOUND, or NF INQ VAR.

FIELDID A one-based index number specifying a field in the compound type.

NAME A character array which will get the name of the field. The name will be NF MAX NAME characters, at most.

OFFSETP An integer which will get the offset of the field.

FIELD_TYPEID

An integer which will get the typeid of the field.

NDIMSP An integer which will get the number of dimensions of the field.

DIM_SIZESP

An integer array which will get the dimension sizes of the field.

Errors

NF_NOERR No error.
NF_EBADTYPEID

Bad type id.

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/fst types.F.

C     Check the first field of the compound type.
retval = nf_inq_compound_field(ncid, typeids(1), 1, name_in,
&     offset_in, field_typeid_in, ndims_in, dim_sizes_in)
if (retval .ne. nf_noerr) call handle_err(retval)
if (name_in(1:len(u_name)) .ne. u_name .or. offset_in .ne. 0 .or.
&     field_typeid_in .ne. NF_INT .or. ndims_in .ne. 0) stop 2
retval = nf_inq_compound_fieldname(ncid, typeids(1), 1, name_in)
if (retval .ne. nf_noerr) call handle_err(retval)
if (name_in(1:len(u_name)) .ne. u_name) stop 2
retval = nf_inq_compound_fieldoffset(ncid, typeids(1), 1,
&     offset_in)
if (retval .ne. nf_noerr) call handle_err(retval)
if (offset_in .ne. 0) stop 2
retval = nf_inq_compound_fieldtype(ncid, typeids(1), 1,
&     field_typeid_in)
if (retval .ne. nf_noerr) call handle_err(retval)
if (field_typeid_in .ne. NF_INT) stop 2
retval = nf_inq_compound_fieldndims(ncid, typeids(1), 1,
&     ndims_in)
if (retval .ne. nf_noerr) call handle_err(retval)
if (ndims_in .ne. 0) stop 2

5.6 Variable Length Array Introduction

NetCDF-4 added support for a variable length array type. This is not supported in classic or 64-bit offset files, or in netCDF-4 files which were created with the NF CLASSIC MODEL flag.

A variable length array is represented in C as a structure from HDF5, the nf vlen t structure. It contains a len member, which contains the length of that array, and a pointer to the array.

So an array of VLEN in C is an array of nc vlen t structures. The only way to handle this in Fortran is with a character buffer sized correctly for the platform.

The extra access functions NF GET VLEN ELEMENT and NF PUT VLEN ELEMENT to get and put one VLEN element. (That is, one array of variable length.) When calling the put, the data are not copied from the source. When calling the get the data are copied from VLEN allocated memory, which must still be freed (see below).

VLEN arrays are handled differently with respect to allocation of memory. Generally, when reading data, it is up to the user to malloc (and subsequently free) the memory needed to hold the data. It is up to the user to ensure that enough memory is allocated.

With VLENs, this is impossible. The user cannot know the size of an array of VLEN until after reading the array. Therefore when reading VLEN arrays, the netCDF library will allocate the memory for the data within each VLEN.

It is up to the user, however, to eventually free this memory. This is not just a matter of one call to free, with the pointer to the array of VLENs; each VLEN contains a pointer which must be freed.

5.6.1 Define a Variable Length Array (VLEN): NF DEF VLEN

Use this function to define a variable length array type.

Usage

     INTEGER FUNCTION NF_DEF_VLEN(INTEGER NCID, CHARACTER*(*) NAME,
             INTEGER BASE_TYPEID, INTEGER XTYPEP)

NCID The ncid of the file to create the VLEN type in.

NAME A name for the VLEN type.

BASE_TYPEID

The typeid of the base type of the VLEN. For example, for a VLEN of shorts, the base type is NF SHORT. This can be a user defined type.

XTYPEP The typeid of the new VLEN type will be set here.

Errors

NF_NOERR No error.
NF_EMAXNAME

NF MAX NAME exceeded.

NF_ENAMEINUSE

Name is already in use.

NF_EBADNAME

Attribute or variable name contains illegal characters.

NF_EBADID

ncid invalid.

NF_EBADGRPID

Group ID part of ncid was invalid.

NF_EINVAL

Size is invalid.

NF_ENOMEM

Out of memory.

Example

This example is from nf test/ftst vars4.F.

     C     Create the vlen type.
           retval = nf_def_vlen(ncid, vlen_type_name, nf_int, vlen_typeid)
           if (retval .ne. nf_noerr) call handle_err(retval)

5.6.2 Learning about a Variable Length Array (VLEN) Type: NF INQ VLEN

Use this type to learn about a vlen.

Usage

     INTEGER FUNCTION NF_INQ_VLEN(INTEGER NCID, INTEGER XTYPE,
             CHARACTER*(*) NAME, INTEGER DATUM_SIZEP, INTEGER
             BASE_NF_TYPEP)

NCID The ncid of the file that contains the VLEN type.

XTYPE The type of the VLEN to inquire about.

NAME The name of the VLEN type. The name will be NF MAX NAME characters or less.

DATUM_SIZEP

A pointer to a size t, this will get the size of one element of this vlen.

BASE_NF_TYPEP

An integer that will get the type of the VLEN base type. (In other words, what type is this a VLEN of?)

Errors

NF_NOERR No error.
NF_EBADTYPE

Can’t find the typeid.

NF_EBADID

ncid invalid.

NF_EBADGRPID

Group ID part of ncid was invalid.

Example

This example is from nf test/ftst vars4.F.

     C     Use nf_inq_vlen and make sure we get the same answers as we did
     C     with nf_inq_user_type.
           retval = nf_inq_vlen(ncid, typeids(1), type_name, base_size,
          &     base_type)
           if (retval .ne. nf_noerr) call handle_err(retval)

5.6.3 Releasing Memory for a Variable Length Array (VLEN) Type: NF FREE VLEN

When a VLEN is read into user memory from the file, the HDF5 library performs memory allocations for each of the variable length arrays contained within the VLEN structure. This memory must be freed by the user to avoid memory leaks.

This violates the normal netCDF expectation that the user is responsible for all memory allocation. But, with VLEN arrays, the underlying HDF5 library allocates the memory for the user, and the user is responsible for deallocating that memory.

Usage

     INTEGER FUNCTION NF_FREE_VLEN(CHARACTER VL);

VL The variable length array structure which is to be freed.

Errors

NF_NOERR No error.
NF_EBADTYPE

Can’t find the typeid.

Example

5.6.4 Set a Variable Length Array with NF PUT VLEN ELEMENT

Use this to set the element of the (potentially) n-dimensional array of VLEN. That is, this sets the data in one variable length array.

Usage

     INTEGER FUNCTION NF_PUT_VLEN_ELEMENT(INTEGER NCID, INTEGER XTYPE,
             CHARACTER*(*) VLEN_ELEMENT, INTEGER LEN, DATA)

NCID The ncid of the file that contains the VLEN type.

XTYPE The type of the VLEN.

VLEN_ELEMENT

The VLEN element to be set.

LEN The number of entries in this array.

DATA The data to be stored. Must match the base type of this VLEN.

Errors

NF_NOERR No error.
NF_EBADTYPE

Can’t find the typeid.

NF_EBADID

ncid invalid.

NF_EBADGRPID

Group ID part of ncid was invalid.

Example

This example is from nf test/ftst vars4.F.

     C     Set up the vlen with this helper function, since F77 can't deal
     C     with pointers.
           retval = nf_put_vlen_element(ncid, vlen_typeid, vlen,
          &     vlen_len, data1)
           if (retval .ne. nf_noerr) call handle_err(retval)

5.6.5 Set a Variable Length Array with NF GET VLEN ELEMENT

Use this to set the element of the (potentially) n-dimensional array of VLEN. That is, this sets the data in one variable length array.

Usage

     INTEGER FUNCTION NF_GET_VLEN_ELEMENT(INTEGER NCID, INTEGER XTYPE,
             CHARACTER*(*) VLEN_ELEMENT, INTEGER LEN, DATA)

NCID The ncid of the file that contains the VLEN type.

XTYPE The type of the VLEN.

VLEN_ELEMENT

The VLEN element to be set.

LEN This will be set to the number of entries in this array.

DATA The data will be copied here. Sufficient storage must be available or bad things will happen to you.

Errors

NF_NOERR No error.
NF_EBADTYPE

Can’t find the typeid.

NF_EBADID

ncid invalid.

NF_EBADGRPID

Group ID part of ncid was invalid.

Example

This example is from nf test/ftst vars4.F.

     C     Read the vlen attribute.
           retval = nf_get_att(ncid, NF_GLOBAL, 'att1', vlen_in)

           if (retval .ne. nf_noerr) call handle_err(retval)
     C     Get the data from the vlen we just read.
           retval = nf_get_vlen_element(ncid, vlen_typeid, vlen_in,
          &     vlen_len_in, data1_in)
           if (retval .ne. nf_noerr) call handle_err(retval)

5.7 Opaque Type Introduction

NetCDF-4 added support for the opaque type. This is not supported in classic or 64-bit offset files.

The opaque type is a type which is a collection of objects of a known size. (And each object is the same size). Nothing is known to netCDF about the contents of these blobs of data, except their size in bytes, and the name of the type.

To use an opaque type, first define it with Section 5.7.1 [NF DEF OPAQUE], page 58. If encountering an enum type in a new data file, use Section 5.7.2 [NF INQ OPAQUE], page 59 to learn its name and size.

5.7.1 Creating Opaque Types: NF DEF OPAQUE

Create an opaque type. Provide a size and a name.

Usage

     INTEGER FUNCTION NF_DEF_OPAQUE(INTEGER NCID, CHARACTER*(*) NAME,
             INTEGER SIZE, INTEGER TYPEIDP)

NCID The groupid where the type will be created. The type may be used anywhere in the file, no matter what group it is in.

NAME The name for this type. Must be shorter than NF MAX NAME.

SIZE The size of each opaque object.

TYPEIDP Pointer where the new typeid for this type is returned. Use this typeid when defining variables of this type with Section 6.3 [NF DEF VAR], page 66.

Errors

NF_NOERR No error.
NF_EBADTYPEID

Bad typeid.

NF_EBADFIELDID

Bad fieldid.

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/ftst vars3.F.

     C     Create the opaque type.
           retval = nf_def_opaque(ncid, opaque_size, opaque_type_name,
          &     opaque_typeid)
           if (retval .ne. nf_noerr) call handle_err(retval)

5.7.2 Learn About an Opaque Type: NF INQ OPAQUE

Given a typeid, get the information about an opaque type.

Usage

     INTEGER FUNCTION NF_INQ_OPAQUE(INTEGER NCID, INTEGER XTYPE,
             CHARACTER*(*) NAME, INTEGER SIZEP)

NCID The ncid for the group containing the opaque type.

XTYPE The typeid for this opaque type, as returned by NF DEF COMPOUND, or NF INQ VAR.

NAME The name of the opaque type will be copied here. It will be NF MAX NAME bytes or less.

SIZEP The size of the opaque type will be copied here.

Errors

NF_NOERR No error.
NF_EBADTYPEID

Bad typeid.

NF_EBADFIELDID

Bad fieldid.

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/ftst vars3.F.

     C     Use nf_inq_opaque and make sure we get the same answers as we did
     C     with nf_inq_user_type.
           retval = nf_inq_opaque(ncid, typeids(2), type_name, base_size)
           if (retval .ne. nf_noerr) call handle_err(retval)

5.8 Enum Type Introduction

NetCDF-4 added support for the enum type. This is not supported in classic or 64-bit offset files.

5.8.1 Creating a Enum Type: NF DEF ENUM

Create an enum type. Provide an ncid, a name, and a base integer type.

After calling this function, fill out the type with repeated calls to NF INSERT ENUM (see Section 5.8.2 [NF INSERT ENUM], page 61). Call NF INSERT ENUM once for each value you wish to make part of the enumeration.

Usage

     INTEGER FUNCTION NF_DEF_ENUM(INTEGER NCID, INTEGER BASE_TYPEID,
             CHARACTER*(*) NAME, INTEGER TYPEIDP)

NCID The groupid where this compound type will be created.

BASE_TYPEID

The base integer type for this enum. Must be one of: NF BYTE, NF UBYTE, NF SHORT, NF USHORT, NF INT, NF UINT, NF INT64, NF UINT64.

NAME The name of the new enum type.

TYPEIDP The typeid of the new type will be placed here.

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENAMEINUSE

That name is in use. Compound type names must be unique in the data file.

NF_EMAXNAME

Name exceeds max length NF MAX NAME.

NF_EBADNAME

Name contains illegal characters.

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag NF NETCDF4. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

NF_EPERM Attempt to write to a read-only file.

NF_ENOTINDEFINE

Not in define mode.

This example is from nf test/ftst vars3.F.

     C     Create the enum type.
           retval = nf_def_enum(ncid, NF_INT, enum_type_name, enum_typeid)
           if (retval .ne. nf_noerr) call handle_err(retval)

5.8.2 Inserting a Field into a Enum Type: NF INSERT ENUM

Insert a named member into a enum type.

Usage

     INTEGER FUNCTION NF_INSERT_ENUM(INTEGER NCID, INTEGER XTYPE,
             CHARACTER IDENTIFIER, INTEGER VALUE)

NCID The ncid of the group which contains the type.

TYPEID The typeid for this enum type, as returned by nf def enum, or nf inq var.

IDENTIFIER

The identifier of the new member.

VALUE The value that is to be associated with this member.

Errors

NF_NOERR No error.
NF_EBADID

Bad group id.

NF_ENAMEINUSE

That name is in use. Field names must be unique within a enum type.

NF_EMAXNAME

Name exceed max length NF MAX NAME.

NF_EBADNAME

Name contains illegal characters.

NF_ENOTNC4

Attempting a netCDF-4 operation on a netCDF-3 file. NetCDF-4 operations can only be performed on files defined with a create mode which includes flag NF NETCDF4. (see Section 2.8 [NF OPEN], page 13).

NF_ESTRICTNC3

This file was created with the strict netcdf-3 flag, therefore netcdf-4 operations are not allowed. (see Section 2.8 [NF OPEN], page 13).

NF_EHDFERR

An error was reported by the HDF5 layer.

NF_ENOTINDEFINE

Not in define mode.

Example

This example is from nf test/ftst vars3.F.

           one = 1
           zero = 0
           retval = nf_insert_enum(ncid, enum_typeid, zero_name, zero)
           if (retval .ne. nf_noerr) call handle_err(retval)
           retval = nf_insert_enum(ncid, enum_typeid, one_name, one)
           if (retval .ne. nf_noerr) call handle_err(retval)

5.8.3 Learn About a Enum Type: NF INQ ENUM

Get information about a user-defined enumeration type.

Usage

     INTEGER FUNCTION NF_INQ_ENUM(INTEGER NCID, INTEGER XTYPE,
             CHARACTER*(*) NAME, INTEGER BASE_NF_TYPE, INTEGER BASE_SIZE,
             INTEGER NUM_MEMBERS)

NCID The group ID of the group which holds the enum type.

XTYPE The typeid for this enum type, as returned by NF DEF ENUM, or NF INQ VAR.

NAME Character array which will get the name. It will have a maximum length of NF MAX NAME.

BASE_NF_TYPE

An integer which will get the base integer type of this enum.

BASE_SIZE

An integer which will get the size (in bytes) of the base integer type of this enum.

NUM_MEMBERS

An integer which will get the number of members defined for this enumeration type.

Errors

NF_NOERR No error.
NF_EBADTYPEID

Bad type id.

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

In this example from nf test/ftst vars3.F, an enum type is created and then examined:

           retval = nf_inq_enum(ncid, typeids(1), type_name, base_type,
          &     base_size, num_members)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (base_type .ne. NF_INT .or. num_members .ne. 2) stop 2

5.8.4 Learn the Name of a Enum Type: nf inq enum member

Get information about a member of an enum type.

Usage

     INTEGER FUNCTION NF_INQ_ENUM_MEMBER(INTEGER NCID, INTEGER XTYPE,
             INTEGER IDX, CHARACTER*(*) NAME, INTEGER VALUE)

NCID The groupid where this enum type exists.

XTYPE The typeid for this enum type.

IDX The one-based index number for the member of interest.

NAME A character array which will get the name of the member. It will have a maximum length of NF MAX NAME.

VALUE An integer that will get the value associated with this member.

Errors

NF_NOERR No error.
NF_EBADTYPEID

Bad type id.

NF_EHDFERR

An error was reported by the HDF5 layer.

Example

This example is from nf test/ftst vars3.F:

     C     Check the members of the enum type.
           retval = nf_inq_enum_member(ncid, typeids(1), 1, member_name,
          &     member_value)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (member_name(1:len(zero_name)) .ne. zero_name .or.
          &     member_value .ne. 0) stop 2
           retval = nf_inq_enum_member(ncid, typeids(1), 2, member_name,
          &     member_value)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (member_name(1:len(one_name)) .ne. one_name .or.
          &     member_value .ne. 1) stop 2

5.8.5 Learn the Name of a Enum Type: NF INQ ENUM IDENT

Get the name which is associated with an enum member value.

This is similar to NF INQ ENUM MEMBER, but instead of using the index of the member, you use the value of the member.

Usage

     INTEGER FUNCTION NF_INQ_ENUM_IDENT(INTEGER NCID, INTEGER XTYPE,
             INTEGER VALUE, CHARACTER*(*) IDENTIFIER)

NCID The groupid where this enum type exists.

XTYPE The typeid for this enum type.

VALUE The value for which an identifier is sought.

IDENTIFIER

A character array that will get the identifier. It will have a maximum length of NF MAX NAME.

Return Code

NF_NOERR No error.
NF_EBADTYPEID

Bad type id, or not an enum type.

NF_EHDFERR

An error was reported by the HDF5 layer.

NF_EINVAL

The value was not found in the enum.

Example

In this example from nf test/ftst vars3.F, the values for 0 and 1 are checked in an enum.

           retval = nf_inq_enum_ident(ncid, typeids(1), 0, member_name)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (member_name(1:len(zero_name)) .ne. zero_name) stop 2
           retval = nf_inq_enum_ident(ncid, typeids(1), 1, member_name)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (member_name(1:len(one_name)) .ne. one_name) stop 2

6 Variables

6.1 Variables Introduction

Variables for a netCDF dataset are defined when the dataset is created, while the netCDF dataset is in define mode. Other variables may be added later by reentering define mode. A netCDF variable has a name, a type, and a shape, which are specified when it is defined. A variable may also have values, which are established later in data mode.

Ordinarily, the name, type, and shape are fixed when the variable is first defined. The name may be changed, but the type and shape of a variable cannot be changed. However, a variable defined in terms of the unlimited dimension can grow without bound in that dimension.

A netCDF variable in an open netCDF dataset is referred to by a small integer called a variable ID.

Variable IDs reflect the order in which variables were defined within a netCDF dataset. Variable IDs are 1, 2, 3,..., in the order in which the variables were defined. A function is available for getting the variable ID from the variable name and vice-versa.

Attributes (see Chapter 7 [Attributes], page 111) may be associated with a variable to specify such properties as units.

Operations supported on variables are:

• Create a variable, given its name, data type, and shape.

• Get a variable ID from its name.

• Get a variable’s name, data type, shape, and number of attributes from its ID.

• Put a data value into a variable, given variable ID, indices, and value.

• Put an array of values into a variable, given variable ID, corner indices, edge lengths, and a block of values.

• Put a subsampled or mapped array-section of values into a variable, given variable ID, corner indices, edge lengths, stride vector, index mapping vector, and a block of values.

• Get a data value from a variable, given variable ID and indices.

• Get an array of values from a variable, given variable ID, corner indices, and edge lengths.

• Get a subsampled or mapped array-section of values from a variable, given variable ID, corner indices, edge lengths, stride vector, and index mapping vector.

• Rename a variable.

6.2 Language Types Corresponding to netCDF external

data types

The following table gives the netCDF external data types and the corresponding type constants for defining variables in the FORTRAN interface:

Type  FORTRAN API Mnemonic  Bits
byte  NF BYTE  8

char  NF CHAR  8
short  NF SHORT  16
int  NF INT  32
float  NF FLOAT  32
double  NF DOUBLE  64

The first column gives the netCDF external data type, which is the same as the CDL data type. The next column gives the corresponding FORTRAN parameter for use in netCDF functions (the parameters are defined in the netCDF FORTRAN include-file netcdf.inc). The last column gives the number of bits used in the external representation of values of the corresponding type.

Note that there are no netCDF types corresponding to 64-bit integers or to characters wider than 8 bits in the current version of the netCDF library.

6.3 Create a Variable: NF_DEF_VAR

The function NF DEF VAR adds a new variable to an open netCDF dataset in define mode. It returns (as an argument) a variable ID, given the netCDF ID, the variable name, the variable type, the number of dimensions, and a list of the dimension IDs.

Usage

     INTEGER FUNCTION NF_DEF_VAR(INTEGER NCID, CHARACTER*(*) NAME,
                                 INTEGER XTYPE, INTEGER NVDIMS,
                                 INTEGER VDIMS(*), INTEGER varid)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

NAME Variable name.

XTYPE One of the set of predefined netCDF external data types. The type of this parameter, NF TYPE, is defined in the netCDF header file. The valid netCDF external data types are NF BYTE, NF CHAR, NF SHORT, NF INT, NF FLOAT, and NF DOUBLE. If the file is a NetCDF-4/HDF5 file, the additional types NF UBYTE, NF USHORT, NF UINT, NF INT64, NF UINT64, and NF STRING may be used, as well as a user defined type ID.

NVDIMS Number of dimensions for the variable. For example, 2 specifies a matrix, 1 specifies a vector, and 0 means the variable is a scalar with no dimensions. Must not be negative or greater than the predefined constant NF MAX VAR DIMS.

VDIMS Vector of ndims dimension IDs corresponding to the variable dimensions. If the ID of the unlimited dimension is included, it must be first. This argument is ignored if ndims is 0. For expanded model netCDF4/HDF5 files, there may be any number of unlimited dimensions, and they may be used in any element of the dimids array.

varid Returned variable ID.

Errors

NF DEF VAR returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The netCDF dataset is not in define mode.

• The specified variable name is the name of another existing variable.

• The specified type is not a valid netCDF type.

• The specified number of dimensions is negative or more than the constant NF MAX VAR DIMS, the maximum number of dimensions permitted for a netCDF variable.

• One or more of the dimension IDs in the list of dimensions is not a valid dimension ID for the netCDF dataset.

• The number of variables would exceed the constant NF MAX VARS, the maximum number of variables permitted in a netCDF dataset.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF DEF VAR to create a variable named rh of type double with three dimensions, time, lat, and lon in a new netCDF dataset named foo.nc:

INCLUDE 'netcdf.inc'
   ...
INTEGER  STATUS, NCID
INTEGER  LATDIM, LONDIM, TIMDIM  ! dimension IDs
INTEGER  RHID                    ! variable ID
INTEGER  RHDIMS(3)               ! variable shape
   ...
STATUS = NF_CREATE ('foo.nc', NF_NOCLOBBER, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
                                 ! define dimensions
STATUS = NF_DEF_DIM(NCID, 'lat', 5, LATDIM)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_DEF_DIM(NCID, 'lon', 10, LONDIM)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_DEF_DIM(NCID, 'time', NF_UNLIMITED, TIMDIM)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
         ...
                                 ! define variable
RHDIMS(1) = LONDIM
RHDIMS(2) = LATDIM
RHDIMS(3) = TIMDIM
STATUS = NF_DEF_VAR (NCID, 'rh', NF_DOUBLE, 3, RHDIMS, RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.4 Define Chunking Parameters for a Variable: NF_DEF_VAR_

    CHUNKING

The function NF DEF VAR CHUNKING sets the storage parameters for a variable in a netCDF-4 file. It can set the chunk sizes to get chunked storage, or it can set the contiguous flag to get contiguous storage.

The total size of a chunk must be less than 4 GiB. That is, the product of all chunksizes and the size of the data (or the size of nc vlen t for VLEN types) must be less than 4 GiB.

This function may only be called after the variable is defined, but before nf enddef is called. Once the chunking parameters are set for a variable, they cannot be changed.

Usage

     NF_DEF_VAR_CHUNKING(INTEGER NCID, INTEGER VARID, INTEGER STORAGE, INTEGER CHUNKSIZES)

ncid NetCDF ID, from a previous call to NF OPEN or NF CREATE.

varid Variable ID.

storage If NF CONTIGUOUS, then contiguous storage is used for this variable. Variables with one or more unlimited dimensions cannot use contiguous storage. If contiguous storage is turned on, the chunksizes parameter is ignored. If NF CHUNKED, then chunked storage is used for this variable. Chunk sizes may be specified with the chunksizes parameter. Default sizes will be used if chunking is required and this function is not called.

chunksizes

An array of chunk sizes. The array must have the one chunksize for each dimension in the variable. If contiguous storage is used, then the chunksizes parameter is ignored.

Errors

NF DEF VAR CHUNKING returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error.

Possible return codes include:

NF_NOERR No error.
NF_BADID Bad ncid.
NF_ENOTNC4

Not a netCDF-4 file.

NF_ENOTVAR

Can’t find this variable.

NF_ELATEDEF

This variable has already been the subject of a NF ENDDEF call. In netCDF-4 files NF ENDDEF will be called automatically for any data read or write. Once enddef has been called, it is impossible to set the chunking for a variable.

NF_ENOTINDEFINE

Not in define mode. This is returned for netCDF classic or 64-bit offset files, or for netCDF-4 files, when they were been created with NF STRICT NC3 flag. (see Section 2.5 [NF CREATE], page 9).

NF_ESTRICTNC3

Trying to create a var some place other than the root group in a netCDF file with NF STRICT NC3 turned on.

Example

In this example from nf test/ftst vars.F, a file is created, two dimensions and a variable are defined, and the chunksizes of the data are set to the size of the data (that is, data will be written in one chunk).

C     Create the netCDF file.
retval = nf_create(FILE_NAME, NF_NETCDF4, ncid)
if (retval .ne. nf_noerr) call handle_err(retval)
C     Define the dimensions.
retval = nf_def_dim(ncid, "x", NX, x_dimid)
if (retval .ne. nf_noerr) call handle_err(retval)
retval = nf_def_dim(ncid, "y", NY, y_dimid)
if (retval .ne. nf_noerr) call handle_err(retval)
C     Define the variable.
dimids(1) = y_dimid
dimids(2) = x_dimid
retval = NF_DEF_VAR(ncid, "data", NF_INT, NDIMS, dimids, varid)
if (retval .ne. nf_noerr) call handle_err(retval)
C     Turn on chunking.
chunks(1) = NY
chunks(2) = NX
retval = NF_DEF_VAR_chunking(ncid, varid, NF_CHUNKED, chunks)
if (retval .ne. nf_noerr) call handle_err(retval)

6.5 Learn About Chunking Parameters for a Variable: NF_

INQ_VAR_CHUNKING

The function NF INQ VAR CHUNKING returns the chunking settings for a variable in a netCDF-4 file.

Usage

NF_INQ_VAR_CHUNKING(INTEGER NCID, INTEGER VARID, INTEGER STORAGE, INTEGER CHUNKSIZES);

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

STORAGE On return, set to NF CONTIGUOUS if this variable uses contiguous storage, NF CHUNKED if it uses chunked storage.

CHUNKSIZES

An array of chunk sizes. The length of CHUNKSIZES must be the same as the number of dimensions of the variable.

Errors

NF INQ VAR CHUNKING returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error.

Possible return codes include:

NF_NOERR No error.
NF_BADID Bad ncid.
NF_ENOTNC4

Not a netCDF-4 file.

NF_ENOTVAR

Can’t find this variable.

Example

In this example from nf test/ftst vars.F, a variable with chunked storage is check to ensure that the chunksizes are set to expected values.

     C     Is everything set that is supposed to be?
           retval = nf_inq_var_chunking(ncid, varid, storage, chunks_in)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (storage .ne. NF_CHUNKED) stop 2
           if (chunks(1) .ne. chunks_in(1)) stop 2
           if (chunks(2) .ne. chunks_in(2)) stop 2

6.6 Define Fill Parameters for a Variable: nf_def_var_fill

The function NF DEF VAR FILL sets the fill parameters for a variable in a netCDF-4 file.

This function must be called after the variable is defined, but before NF ENDDEF is called.

Usage

     NF_DEF_VAR_FILL(INTEGER NCID, INTEGER VARID, INTEGER NO_FILL, FILL_VALUE);

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

NO_FILL Set to non-zero value to set no fill mode on a variable. When this mode is on, fill values will not be written for the variable. This is helpful in high performance applications. For netCDF-4/HDF5 files (whether classic model or not), this may only be changed after the variable is defined, but before it is committed to disk (i.e. before the first NF ENDDEF after the NF DEF VAR.) For classic and 64-bit offset file, the no fill mode may be turned on and off at any time.

FILL_VALUE

A value which will be used as the fill value for the variable. Must be the same type as the variable. This will be written to a FillValue attribute, created for this purpose. If NULL, this argument will be ignored.

Return Codes

NF_NOERR No error.
NF_BADID Bad ncid.
NF_ENOTNC4

Not a netCDF-4 file.

NF_ENOTVAR

Can’t find this variable.

NF_ELATEDEF

This variable has already been the subject of a NF ENDDEF call. In netCDF-4 files NF ENDDEF will be called automatically for any data read or write. Once enddef has been called, it is impossible to set the fill for a variable.

NF_ENOTINDEFINE

Not in define mode. This is returned for netCDF classic or 64-bit offset files, or for netCDF-4 files, when they were been created with NF STRICT NC3 flag. (see Section 2.5 [NF CREATE], page 9).

NF_EPERM Attempt to create object in read-only file.

Example

6.7 Learn About Fill Parameters for a Variable: NF_INQ_VAR_

    FILL

The function NF INQ VAR FILL returns the fill settings for a variable in a netCDF-4 file.

Usage

     NF_INQ_VAR_FILL(INTEGER NCID, INTEGER VARID, INTEGER NO_FILL, FILL_VALUE)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

NO_FILL An integer which will get a 1 if no fill mode is set for this variable, and a zero if it is not set

FILL_VALUE

This will get the fill value for this variable. This parameter will be ignored if it is NULL.

Return Codes

NF_NOERR No error.
NF_BADID Bad ncid.
NF_ENOTNC4

Not a netCDF-4 file.

NF_ENOTVAR

Can’t find this variable.

Example

6.8 Define Compression Parameters for a Variable: NF_DEF_

    VAR_DEFLATE

The function NF DEF VAR DEFLATE sets the deflate parameters for a variable in a netCDF-4 file.

When using parallel I/O for writing data, deflate cannot be used. This is because the compression makes it impossible for the HDF5 library to exactly map the data to disk location.

(Deflated data can be read with parallel I/O).

NF DEF VAR DEFLATE must be called after the variable is defined, but before NF ENDDEF is called.

Usage

     NF_DEF_VAR_DEFLATE(INTEGER NCID, INTEGER VARID, INTEGER SHUFFLE, INTEGER DEFLATE,
                        INTEGER DEFLATE_LEVEL);

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

SHUFFLE If non-zero, turn on the shuffle filter.

DEFLATE If non-zero, turn on the deflate filter at the level specified by the deflate level parameter.

DEFLATE_LEVEL

Must be between 0 (no deflate, the default) and 9 (slowest, but “best” deflate).

If set to zero, no deflation takes place and the def var deflate call is ignored. This is slightly different from HDF5 handing of 0 deflate, which turns on the filter but makes only trivial changes to the data.

Informal testing at NetCDF World Headquarters suggests that there is little to be gained (with the limited set of test data used here), in setting the deflate level above 2 or 3.

Errors

NF DEF VAR DEFLATE returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error.

Possible return codes include:

NF_NOERR No error.
NF_BADID Bad ncid.
NF_ENOTNC4

Not a netCDF-4 file.

NF_ENOTVAR

Can’t find this variable.

NF_ELATEDEF

This variable has already been the subject of a NF ENDDEF call. In netCDF-4 files NF ENDDEF will be called automatically for any data read or write. Once enddef has been called, it is impossible to set the deflate for a variable.

NF_ENOTINDEFINE

Not in define mode. This is returned for netCDF classic or 64-bit offset files, or for netCDF-4 files, when they were been created with NF STRICT NC3 flag. (see Section 2.5 [NF CREATE], page 9).

NF_EPERM Attempt to create object in read-only file.

NF_EINVAL

Invalid deflate level. The deflate level must be between 0 and 9, inclusive.

Example

In this example from nf test/ftst vars.F, a file is created with two dimensions and one variable. Chunking, deflate, and the fletcher32 filter are turned on. The deflate level is set to 4 below.

     C     Create the netCDF file.
           retval = nf_create(FILE_NAME, NF_NETCDF4, ncid)
           if (retval .ne. nf_noerr) call handle_err(retval)
     C     Define the dimensions.
           retval = nf_def_dim(ncid, "x", NX, x_dimid)
           if (retval .ne. nf_noerr) call handle_err(retval)
           retval = nf_def_dim(ncid, "y", NY, y_dimid)
           if (retval .ne. nf_noerr) call handle_err(retval)
     C     Define the variable.
           dimids(1) = y_dimid
           dimids(2) = x_dimid
           retval = NF_DEF_VAR(ncid, "data", NF_INT, NDIMS, dimids, varid)
           if (retval .ne. nf_noerr) call handle_err(retval)

     C     Turn on chunking.
           chunks(1) = NY
           chunks(2) = NX
           retval = NF_DEF_VAR_CHUNKING(ncid, varid, NF_CHUNKED, chunks)
           if (retval .ne. nf_noerr) call handle_err(retval)
     C     Turn on deflate compression, fletcher32 checksum.
           retval = NF_DEF_VAR_deflate(ncid, varid, 0, 1, 4)
           if (retval .ne. nf_noerr) call handle_err(retval)
           retval = NF_DEF_VAR_FLETCHER32(ncid, varid, NF_FLETCHER32)
           if (retval .ne. nf_noerr) call handle_err(retval)

6.9 Learn About Deflate Parameters for a Variable: NF_INQ_

    VAR_DEFLATE

The function NF INQ VAR DEFLATE returns the deflate settings for a variable in a netCDF-4 file.

It is not necessary to know the deflate settings to read the variable. (Deflate is completely transparent to readers of the data).

Usage

     NF_INQ_VAR_DEFLATE(INTEGER NCID, INTEGER VARID, INTEGER SHUFFLE,
                        INTEGER DEFLATE, INTEGER DEFLATE_LEVEL);

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

SHUFFLE NF INQ VAR DEFLATE will set this to a 1 if the shuffle filter is turned on for this variable, and a 0 otherwise.

DEFLATE NF INQ VAR DEFLATE will set this to a 1 if the deflate filter is turned on for this variable, and a 0 otherwise.

DEFLATE_LEVEL

NF INQ VAR DEFLATE function will write the deflate level here, if deflate is in use.

Errors

NF INQ VAR DEFLATE returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error.

Possible return codes include:

NF_NOERR No error.
NF_BADID Bad ncid.
NF_ENOTNC4

Not a netCDF-4 file.

NF_ENOTVAR

Can’t find this variable.

Example

In this example code from nf test/ftst vars.F, a file with a variable using deflate is opened, and the deflate level checked.

     C     Is everything set that is supposed to be?
           retval = nf_inq_var_deflate(ncid, varid, shuffle, deflate,
          +     deflate_level)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (shuffle .ne. 0 .or. deflate .ne. 1 .or.
          +     deflate_level .ne. 4) stop 2

6.10 Define Checksum Parameters for a Variable: NF_DEF_

     VAR_FLETCHER32

The function NF DEF VAR FLETCHER32 sets the checksum property for a variable in a netCDF-4 file.

This function may only be called after the variable is defined, but before NF ENDDEF is called.

Usage

     NF_DEF_VAR_FLETCHER32(INTEGER NCID, INTEGER VARID, INTEGER CHECKSUM);

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

CHECKSUM If this is NF FLETCHER32, fletcher32 checksums will be turned on for this variable.

Errors

NF DEF VAR FLETCHER32 returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error.

Possible return codes include:

NF_NOERR No error.
NF_BADID Bad ncid.
NF_ENOTNC4

Not a netCDF-4 file.

NF_ENOTVAR

Can’t find this variable.

NF_ELATEDEF

This variable has already been the subject of a NF ENDDEF call. In netCDF-4 files NF ENDDEF will be called automatically for any data read or write.

Once enddef has been called, it is impossible to set the checksum property for a variable.

NF_ENOTINDEFINE

Not in define mode. This is returned for netCDF classic or 64-bit offset files, or for netCDF-4 files, when they were been created with NF STRICT NC3 flag. (see Section 2.5 [NF CREATE], page 9).

NF_EPERM Attempt to create object in read-only file.

Example

In this example from nf test/ftst vars.F, the variable in a file has the Fletcher32 checksum filter turned on.

C     Create the netCDF file.
retval = nf_create(FILE_NAME, NF_NETCDF4, ncid)
if (retval .ne. nf_noerr) call handle_err(retval)
C     Define the dimensions.
retval = nf_def_dim(ncid, "x", NX, x_dimid)
if (retval .ne. nf_noerr) call handle_err(retval)
retval = nf_def_dim(ncid, "y", NY, y_dimid)
if (retval .ne. nf_noerr) call handle_err(retval)
C     Define the variable.
dimids(1) = y_dimid
dimids(2) = x_dimid
retval = NF_DEF_VAR(ncid, "data", NF_INT, NDIMS, dimids, varid)
if (retval .ne. nf_noerr) call handle_err(retval)
C     Turn on chunking.
chunks(1) = NY
chunks(2) = NX
retval = NF_DEF_VAR_CHUNKING(ncid, varid, NF_CHUNKED, chunks)
if (retval .ne. nf_noerr) call handle_err(retval)
C     Turn on deflate compression, fletcher32 checksums.
retval = NF_DEF_VAR_DEFLATE(ncid, varid, 0, 1, 4)
if (retval .ne. nf_noerr) call handle_err(retval)
retval = NF_DEF_VAR_FLETCHER32(ncid, varid, NF_FLETCHER32)
if (retval .ne. nf_noerr) call handle_err(retval)

6.11 Learn About Checksum Parameters for a Variable: NF_

INQ_VAR_FLETCHER32

The function NF INQ VAR FLETCHER32 returns the checksum settings for a variable in a netCDF-4 file.

Usage

     NF_INQ_VAR_FLETCHER32(INTEGER NCID, INTEGER VARID, INTEGER CHECKSUM);

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE. VARID Variable ID. CHECKSUM NF INQ VAR FLETCHER32 will set this to NF FLETCHER32 if the

fletcher32 filter is turned on for this variable, and NF NOCHECKSUM if it is not.

Errors

NF INQ VAR FLETCHER32 returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error.

Possible return codes include:

NF_NOERR No error.
NF_BADID Bad ncid.
NF_ENOTNC4

Not a netCDF-4 file.

NF_ENOTVAR

Can’t find this variable.

Example

In this example from nf test/ftst vars.F the checksum filter is checked for a file. Since it was turned on for this variable, the checksum variable is set to NF FLETCHER32.

           retval = nf_inq_var_fletcher32(ncid, varid, checksum)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (checksum .ne. NF_FLETCHER32) stop 2

6.12 Define Endianness of a Variable: NF_DEF_VAR_ENDIAN

The function NF DEF VAR ENDIAN sets the endianness for a variable in a netCDF-4 file.

This function must be called after the variable is defined, but before NF ENDDEF is called.

By default, netCDF-4 variables are in native endianness. That is, they are big-endian on a big-endian machine, and little-endian on a little endian machine.

In some cases a user might wish to change from native endianness to either big or littleendianness. This function allows them to do that.

Usage

     NF_DEF_VAR_ENDIAN(INTEGER NCID, INTEGER VARID, INTEGER ENDIAN)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE. VARID Variable ID.

ENDIAN Set to NF ENDIAN NATIVE for native endianness. (This is the default). Set to NF ENDIAN LITTLE for little endian, or NF ENDIAN BIG for big endian.

Errors

NF DEF VAR ENDIAN returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error.

Possible return codes include:

NF_NOERR No error.
NF_BADID Bad ncid.
NF_ENOTNC4

Not a netCDF-4 file.

NF_ENOTVAR

Can’t find this variable.

NF_ELATEDEF

This variable has already been the subject of a NF ENDDEF call. In netCDF-4 files NF ENDDEF will be called automatically for any data read or write. Once enddef has been called, it is impossible to set the endianness of a variable.

NF_ENOTINDEFINE

Not in define mode. This is returned for netCDF classic or 64-bit offset files, or for netCDF-4 files, when they were been created with NF STRICT NC3 flag, and the file is not in define mode. (see Section 2.5 [NF CREATE], page 9).

NF_EPERM Attempt to create object in read-only file.

Example

In this example from nf test/ftst vars.c, a file is created with one variable, and its endianness is set to NF ENDIAN BIG.

     C     Create the netCDF file.
           retval = nf_create(FILE_NAME, NF_NETCDF4, ncid)
           if (retval .ne. nf_noerr) call handle_err(retval)
     C     Define the dimensions.
           retval = nf_def_dim(ncid, "x", NX, x_dimid)
           if (retval .ne. nf_noerr) call handle_err(retval)
           retval = nf_def_dim(ncid, "y", NY, y_dimid)
           if (retval .ne. nf_noerr) call handle_err(retval)
     C     Define the variable.
           dimids(1) = y_dimid
           dimids(2) = x_dimid
           retval = NF_DEF_VAR(ncid, "data", NF_INT, NDIMS, dimids, varid)
           if (retval .ne. nf_noerr) call handle_err(retval)

     C     Turn on chunking.
           chunks(1) = NY
           chunks(2) = NX
           retval = NF_DEF_VAR_chunking(ncid, varid, 0, chunks)
           if (retval .ne. nf_noerr) call handle_err(retval)
     C     Set variable to big-endian (default is whatever is native to
     C     writing machine).
           retval = NF_DEF_VAR_endian(ncid, varid, NF_ENDIAN_BIG)
           if (retval .ne. nf_noerr) call handle_err(retval)

6.13 Learn About Endian Parameters for a Variable: NF_

     INQ_VAR_ENDIAN

The function NF INQ VAR ENDIAN returns the endianness settings for a variable in a netCDF-4 file.

Usage

     NF_INQ_VAR_ENDIAN(INTEGER NCID, INTEGER VARID, INTEGER ENDIAN)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

ENDIAN NF INQ VAR ENDIAN will set this to NF ENDIAN LITTLE if this variable is stored in little-endian format, NF ENDIAN BIG if it is stored in big-endian format, and NF ENDIAN NATIVE if the endianness is not set, and the variable is not created yet.

Errors

NF INQ VAR ENDIAN returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error.

Possible return codes include:

NF_NOERR No error.
NF_BADID Bad ncid.
NF_ENOTNC4

Not a netCDF-4 file.

NF_ENOTVAR

Can’t find this variable.

Example

In this example from nf test/ftst vars.F, the endianness of a variable is checked to make sure it is NF ENDIAN BIG.

           retval = nf_inq_var_endian(ncid, varid, endianness)
           if (retval .ne. nf_noerr) call handle_err(retval)
           if (endianness .ne. NF_ENDIAN_BIG) stop 2

6.14 Get a Variable ID from Its Name: NF INQ VARID

The function NF INQ VARID returns the ID of a netCDF variable, given its name.

Usage

     INTEGER FUNCTION NF_INQ_VARID(INTEGER NCID, CHARACTER*(*) NAME,
                                   INTEGER varid)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

NAME Variable name for which ID is desired.

varid Returned variable ID.

Errors

NF INQ VARID returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The specified variable name is not a valid name for a variable in the specified netCDF dataset.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF INQ VARID to find out the ID of a variable named rh in an existing netCDF dataset named foo.nc:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER  STATUS, NCID, RHID
        ...
     STATUS = NF_OPEN ('foo.nc', NF_NOWRITE, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     STATUS = NF_INQ_VARID (NCID, 'rh', RHID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.15 Get Information about a Variable from Its ID:

NF INQ VAR family

A family of functions that returns information about a netCDF variable, given its ID. Information about a variable includes its name, type, number of dimensions, a list of dimension IDs describing the shape of the variable, and the number of variable attributes that have been assigned to the variable.

The function NF INQ VAR returns all the information about a netCDF variable, given its ID. The other functions each return just one item of information about a variable.

These other functions include NF INQ VARNAME, NF INQ VARTYPE, NF INQ VARNDIMS, NF INQ VARDIMID, and NF INQ VARNATTS.

Usage

     INTEGER FUNCTION NF_INQ_VAR      (INTEGER NCID, INTEGER VARID,
                                       CHARACTER*(*) name, INTEGER xtype,
                                       INTEGER ndims, INTEGER dimids(*),
                                       INTEGER natts)
     INTEGER FUNCTION NF_INQ_VARNAME  (INTEGER NCID, INTEGER VARID,
                                       CHARACTER*(*) name)
     INTEGER FUNCTION NF_INQ_VARTYPE  (INTEGER NCID, INTEGER VARID,
                                       INTEGER xtype)
     INTEGER FUNCTION NF_INQ_VARNDIMS (INTEGER NCID, INTEGER VARID,
                                       INTEGER ndims)
     INTEGER FUNCTION NF_INQ_VARDIMID (INTEGER NCID, INTEGER VARID,
                                       INTEGER dimids(*))
     INTEGER FUNCTION NF_INQ_VARNATTS (INTEGER NCID, INTEGER VARID,
                                       INTEGER natts)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

NAME Returned variable name. The caller must allocate space for the returned name. The maximum possible length, in characters, of a variable name is given by the predefined constant NF MAX NAME.

xtype Returned variable type, one of the set of predefined netCDF external data types. The type of this parameter, NF TYPE, is defined in the netCDF header file. The valid netCDF external data types are NF BYTE, NF CHAR, NF SHORT, NF INT, NF FLOAT, AND NF DOUBLE.

ndims Returned number of dimensions the variable was defined as using. For example, 2 indicates a matrix, 1 indicates a vector, and 0 means the variable is a scalar with no dimensions.

dimids Returned vector of *ndimsp dimension IDs corresponding to the variable dimensions. The caller must allocate enough space for a vector of at least *ndimsp integers to be returned. The maximum possible number of dimensions for a variable is given by the predefined constant NF MAX VAR DIMS.

natts Returned number of variable attributes assigned to this variable.

These functions return the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF INQ VAR to find out about a variable named rh in an existing netCDF dataset named foo.nc:

     INCLUDE 'netcdf.inc'
        ...

INTEGER  STATUS, NCID
INTEGER  RHID               ! variable ID
CHARACTER*31 RHNAME         ! variable name
INTEGER  RHTYPE             ! variable type
INTEGER  RHN                ! number of dimensions
INTEGER  RHDIMS(NF_MAX_VAR_DIMS)   ! variable shape
INTEGER  RHNATT                    ! number of attributes
   ...
STATUS = NF_OPEN ('foo.nc', NF_NOWRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID (NCID, 'rh', RHID)  ! get ID
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_INQ_VAR (NCID, RHID, RHNAME, RHTYPE, RHN, RHDIMS, RHNATT)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.16 Write a Single Data Value: NF PUT VAR1 type

The functions NF PUT VAR1 type put a single data value of the specified type into a variable of an open netCDF dataset that is in data mode. Inputs are the netCDF ID, the variable ID, an index that specifies which value to add or alter, and the data value. The value is converted to the external data type of the variable, if necessary.

Take care when using the simplest forms of this interface with record variables when you don’t specify how many records are to be read. If you try to read all the values of a record variable into an array but there are more records in the file than you assume, more data will be read than you expect, which may cause a segmentation violation.

Usage

INTEGER FUNCTION  NF_PUT_VAR1_TEXT(INTEGER NCID, INTEGER VARID,
                                   INTEGER INDEX(*), CHARACTER CHVAL)
INTEGER FUNCTION  NF_PUT_VAR1_INT1(INTEGER NCID, INTEGER VARID,
                                   INTEGER INDEX(*), INTEGER*1 I1VAL)
INTEGER FUNCTION  NF_PUT_VAR1_INT2(INTEGER NCID, INTEGER VARID,
                                   INTEGER INDEX(*), INTEGER*2 I2VAL)
INTEGER FUNCTION  NF_PUT_VAR1_INT (INTEGER NCID, INTEGER VARID,
                                   INTEGER INDEX(*), INTEGER   IVAL)
INTEGER FUNCTION  NF_PUT_VAR1_REAL(INTEGER NCID, INTEGER VARID,
                                   INTEGER INDEX(*), REAL      RVAL)
INTEGER FUNCTION  NF_PUT_VAR1_DOUBLE(INTEGER NCID, INTEGER VARID,
                                   INTEGER INDEX(*), DOUBLE    DVAL)
INTEGER FUNCTION  NF_PUT_VAR1(INTEGER NCID, INTEGER VARID,
                                   INTEGER INDEX(*), *)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

INDEX The index of the data value to be written. The indices are relative to 1, so for example, the first data value of a two-dimensional variable would have index (1,1).

The elements of index must correspond to the variable’s dimensions. Hence, if the variable uses the unlimited dimension, the last index would correspond to the record number.

CHVAL
I1VAL
I2VAL
IVAL
RVAL

DVAL Pointer to the data value to be written. If the type of data values differs from the netCDF variable type, type conversion will occur. See Section “Type Conversion” in The NetCDF Users Guide.

Errors

NF PUT VAR1 type returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• The specified indices were out of range for the rank of the specified variable. For example, a negative index or an index that is larger than the corresponding dimension length will cause an error.

• The specified value is out of the range of values representable by the external data type of the variable.

• The specified netCDF is in define mode rather than data mode.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF PUT VAR1 DOUBLE to set the (4,3,2) element of the variable named rh to 0.5 in an existing netCDF dataset named foo.nc. For simplicity in this example, we assume that we know that rh is dimensioned with lon, lat, and time, so we want to set the value of rh that corresponds to the fourth lon value, the third lat value, and the second time value:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER  STATUS             ! error status
     INTEGER  NCID
     INTEGER  RHID               ! variable ID
     INTEGER  RHINDX(3)          ! where to put value
     DATA RHINDX /4, 3, 2/
        ...
     STATUS = NF_OPEN ('foo.nc', NF_WRITE, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     STATUS = NF_INQ_VARID (NCID, 'rh', RHID)  ! get ID
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
     STATUS = NF_PUT_VAR1_DOUBLE (NCID, RHID, RHINDX, 0.5)

     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.17 Write an Entire Variable: NF PUT VAR type

The NF PUT VAR type family of functions write all the values of a variable into a netCDF variable of an open netCDF dataset. This is the simplest interface to use for writing a value in a scalar variable or whenever all the values of a multidimensional variable can all be written at once. The values to be written are associated with the netCDF variable by assuming that the last dimension of the netCDF variable varies fastest in the C interface. The values are converted to the external data type of the variable, if necessary.

Take care when using the simplest forms of this interface with record variables when you don’t specify how many records are to be written. If you try to write all the values of a record variable into a netCDF file that has no record data yet (hence has 0 records), nothing will be written. Similarly, if you try to write all of a record variable but there are more records in the file than you assume, more data may be written to the file than you supply, which may result in a segmentation violation.

Usage

     INTEGER FUNCTION NF_PUT_VAR_TEXT  (INTEGER NCID, INTEGER VARID,
                                        CHARACTER*(*) TEXT)
     INTEGER FUNCTION NF_PUT_VAR_INT1  (INTEGER NCID, INTEGER VARID,
                                        INTEGER*1 I1VALS(*))
     INTEGER FUNCTION NF_PUT_VAR_INT2  (INTEGER NCID, INTEGER VARID,
                                        INTEGER*2 I2VALS(*))
     INTEGER FUNCTION NF_PUT_VAR_INT   (INTEGER NCID, INTEGER VARID,
                                        INTEGER IVALS(*))
     INTEGER FUNCTION NF_PUT_VAR_REAL  (INTEGER NCID, INTEGER VARID,
                                        REAL RVALS(*))
     INTEGER FUNCTION NF_PUT_VAR_DOUBLE(INTEGER NCID, INTEGER VARID,
                                        DOUBLE DVALS(*))
     INTEGER FUNCTION NF_PUT_VAR       (INTEGER NCID, INTEGER VARID,
                                        VALS(*))

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

TEXT
I1VALS
I2VALS
IVALS
RVALS
DVALS

VALS The block of data values to be written. The data should be of the type appropriate for the function called. You cannot put CHARACTER data into a numeric variable or numeric data into a text variable. For numeric data, if the type of data differs from the netCDF variable type, type conversion will occur (see Section “Type Conversion” in The NetCDF Users Guide). The order in which the data will be written into the specified variable is with the first dimension varying fastest (like the ordinary FORTRAN convention).

Errors

Members of the NF PUT VAR type family return the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• One or more of the specified values are out of the range of values representable by the external data type of the variable.

• One or more of the specified values are out of the range of values representable by the external data type of the variable.

• The specified netCDF dataset is in define mode rather than data mode.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF PUT VAR DOUBLE to add or change all the values of the variable named rh to 0.5 in an existing netCDF dataset named foo.nc. For simplicity in this example, we assume that we know that rh is dimensioned with lon, lat, and time, and that there are ten lon values, five lat values, and three time values.

INCLUDE 'netcdf.inc'
   ...
PARAMETER (TIMES=3, LATS=5, LONS=10) ! dimension lengths
INTEGER  STATUS, NCID, TIMES
INTEGER  RHID                        ! variable ID
DOUBLE RHVALS(LONS, LATS, TIMES)
   ...
STATUS = NF_OPEN ('foo.nc', NF_WRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID (NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
DO 10 ILON = 1, LONS
   DO 10 ILAT = 1, LATS
      DO 10 ITIME = 1, TIMES
         RHVALS(ILON, ILAT, ITIME) = 0.5
10 CONTINUE
STATUS = NF_PUT_var_DOUBLE (NCID, RHID, RHVALS)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.18 Write an Array of Values: NF PUT VARA type

The function NF PUT VARA type writes values into a netCDF variable of an open netCDF dataset. The part of the netCDF variable to write is specified by giving a corner and a vector of edge lengths that refer to an array section of the netCDF variable. The values to be written are associated with the netCDF variable by assuming that the first dimension of the netCDF variable varies fastest in the FORTRAN interface. The netCDF dataset must be in data mode.

Usage

INTEGER FUNCTION NF_PUT_VARA_TEXT(INTEGER NCID, INTEGER VARID,
INTEGER START(*), INTEGER COUNT(*),
CHARACTER*(*) TEXT)
INTEGER FUNCTION NF_PUT_VARA_INT1(INTEGER NCID, INTEGER VARID,
INTEGER START(*), INTEGER COUNT(*),
INTEGER*1 I1VALS(*))
INTEGER FUNCTION NF_PUT_VARA_INT2(INTEGER NCID, INTEGER VARID,
INTEGER START(*), INTEGER COUNT(*),
INTEGER*2 I2VALS(*))
INTEGER FUNCTION NF_PUT_VARA_INT (INTEGER NCID, INTEGER VARID,
INTEGER START(*), INTEGER COUNT(*),
INTEGER IVALS(*))
INTEGER FUNCTION NF_PUT_VARA_REAL(INTEGER NCID, INTEGER VARID,
INTEGER START(*), INTEGER COUNT(*),
REAL RVALS(*))
INTEGER FUNCTION NF_PUT_VARA_DOUBLE(INTEGER NCID, INTEGER VARID,
INTEGER START(*), INTEGER COUNT(*),
DOUBLE DVALS(*))
INTEGER FUNCTION NF_PUT_VARA     (INTEGER NCID, INTEGER VARID,
INTEGER START(*), INTEGER COUNT(*),
VALS(*))

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

START A vector of integers specifying the index in the variable where the first of the data values will be written. The indices are relative to 1, so for example, the first data value of a variable would have index (1, 1, ..., 1). The length of START must be the same as the number of dimensions of the specified variable. The elements of START must correspond to the variable’s dimensions in order. Hence, if the variable is a record variable, the last index would correspond to the starting record number for writing the data values.

COUNT A vector of integers specifying the edge lengths along each dimension of the block of data values to written. To write a single value, for example, specify COUNT as (1, 1, ..., 1). The length of COUNT is the number of dimensions of the specified variable. The elements of COUNT correspond to the variable’s dimensions. Hence, if the variable is a record variable, the last element of COUNT corresponds to a count of the number of records to write.

Note: setting any element of the count array to zero causes the function to exit without error, and without doing anything.

TEXT
I1VALS
I2VALS
IVALS
RVALS
DVALS

VALS The block of data values to be written. The data should be of the type appropriate for the function called. You cannot put CHARACTER data into a numeric variable or numeric data into a text variable. For numeric data, if the type of data differs from the netCDF variable type, type conversion will occur (see Section “Type Conversion” in The NetCDF Users Guide).

Errors

NF PUT VARA type returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• The specified corner indices were out of range for the rank of the specified variable. For example, a negative index, or an index that is larger than the corresponding dimension length will cause an error.

• The specified edge lengths added to the specified corner would have referenced data out of range for the rank of the specified variable. For example, an edge length that is larger than the corresponding dimension length minus the corner index will cause an error.

• One or more of the specified values are out of the range of values representable by the external data type of the variable.

• The specified netCDF dataset is in define mode rather than data mode.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF PUT VARA DOUBLE to add or change all the values of the variable named rh to 0.5 in an existing netCDF dataset named foo.nc. For simplicity in this example, we assume that we know that rh is dimensioned with time, lat, and lon, and that there are three time values, five lat values, and ten lon values.

     INCLUDE 'netcdf.inc'
        ...
     PARAMETER (NDIMS=3)         ! number of dimensions
     PARAMETER (TIMES=3, LATS=5, LONS=10) ! dimension lengths
     INTEGER  STATUS, NCID, TIMES
     INTEGER  RHID               ! variable ID
     INTEGER  START(NDIMS), COUNT(NDIMS)
     DOUBLE RHVALS(LONS, LATS, TIMES)
     DATA START /1, 1, 1/        ! start at first value
     DATA COUNT /LONS, LATS, TIMES/
        ...

STATUS = NF_OPEN ('foo.nc', NF_WRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID (NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
DO 10 ILON = 1, LONS
   DO 10 ILAT = 1, LATS
      DO 10 ITIME = 1, TIMES
         RHVALS(ILON, ILAT, ITIME) = 0.5
10 CONTINUE
STATUS = NF_PUT_VARA_DOUBLE (NCID, RHID, START, COUNT, RHVALS)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.19 NF PUT VARS type

Each member of the family of functions NF PUT VARS type writes a subsampled (strided) array section of values into a netCDF variable of an open netCDF dataset. The subsampled array section is specified by giving a corner, a vector of counts, and a stride vector. The netCDF dataset must be in data mode.

Usage

INTEGER FUNCTION NF_PUT_VARS_TEXT  (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*),CHARACTER*(*) TEXT)
INTEGER FUNCTION NF_PUT_VARS_INT1  (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*),INTEGER*1 I1VALS(*))
INTEGER FUNCTION NF_PUT_VARS_INT2  (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*),INTEGER*2 I2VALS(*))
INTEGER FUNCTION NF_PUT_VARS_INT   (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*), INTEGER IVALS(*))
INTEGER FUNCTION NF_PUT_VARS_REAL  (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*),  REAL RVALS(*))
INTEGER FUNCTION NF_PUT_VARS_DOUBLE(INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*),  DOUBLE DVALS(*))
INTEGER FUNCTION NF_PUT_VARS       (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*), VALS(*))

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

START A vector of integers specifying the index in the variable where the first of the data values will be written. The indices are relative to 1, so for example, the first data value of a variable would have index (1, 1, ..., 1). The elements of START correspond, in order, to the variable’s dimensions. Hence, if the variable is a record variable, the last index would correspond to the starting record number for writing the data values.

COUNT A vector of integers specifying the number of indices selected along each dimension. To write a single value, for example, specify COUNT as (1, 1, ..., 1). The elements of COUNT correspond, in order, to the variable’s dimensions. Hence, if the variable is a record variable, the last element of COUNT corresponds to a count of the number of records to write. Note: setting any element of the count array to zero causes the function to exit without error, and without doing anything.

STRIDE A vector of integers that specifies the sampling interval along each dimension of the netCDF variable. The elements of the stride vector correspond, in order, to the netCDF variable’s dimensions (STRIDE(1) gives the sampling interval along the most rapidly varying dimension of the netCDF variable). Sampling intervals are specified in type-independent units of elements (a value of 1 selects consecutive elements of the netCDF variable along the corresponding dimension, a value of 2 selects every other element, etc.).

TEXT
I1VALS
I2VALS
IVALS
RVALS
DVALS

VALS The block of data values to be written. The data should be of the type appropriate for the function called. You cannot put CHARACTER data into a numeric variable or numeric data into a text variable. For numeric data, if the type of data differs from the netCDF variable type, type conversion will occur (see Section “Type Conversion” in The NetCDF Users Guide).

Errors

NF PUT VARS type returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• The specified start, count and stride generate an index which is out of range.

• One or more of the specified values are out of the range of values representable by the external data type of the variable.

• The specified netCDF is in define mode rather than data mode.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example of using NF PUT VARS REAL to write – from an internal array – every other point of a netCDF variable named rh which is described by the FORTRAN declaration REAL RH(6,4) (note the size of the dimensions):

INCLUDE 'netcdf.inc'
   ...
PARAMETER (NDIM=2)   ! rank of netCDF variable
INTEGER NCID         ! netCDF dataset ID
INTEGER STATUS       ! return code
INTEGER RHID         ! variable ID
INTEGER START(NDIM)  ! netCDF variable start point
INTEGER COUNT(NDIM)  ! size of internal array
INTEGER STRIDE(NDIM) ! netCDF variable subsampling intervals
REAL RH(3,2)         ! note subsampled sizes for netCDF variable
                     ! dimensions
DATA START   /1, 1/  ! start at first netCDF variable value
DATA COUNT   /3, 2/  ! size of internal array: entire (subsampled)
                     ! netCDF variable
DATA STRIDE /2, 2/   ! access every other netCDF element
   ...
STATUS = NF_OPEN('foo.nc', NF_WRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID(NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_PUT_VARS_REAL(NCID, RHID, START, COUNT, STRIDE, RH)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.20 NF PUT VARM type

The NF PUT VARM type family of functions writes a mapped array section of values into a netCDF variable of an open netCDF dataset. The mapped array section is specified by giving a corner, a vector of counts, a stride vector, and an index mapping vector. The index mapping vector is a vector of integers that specifies the mapping between the dimensions of a netCDF variable and the in-memory structure of the internal data array. No assumptions are made about the ordering or length of the dimensions of the data array. The netCDF dataset must be in data mode.

Usage

INTEGER FUNCTION NF_PUT_VARM_TEXT  (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*), INTEGER IMAP(*),
                            CHARACTER*(*) TEXT)
INTEGER FUNCTION NF_PUT_VARM_INT1  (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*), INTEGER IMAP(*),
                            INTEGER*1 I1VALS(*))
INTEGER FUNCTION NF_PUT_VARM_INT2  (INTEGER NCID, INTEGER VARID,

                       INTEGER START(*), INTEGER COUNT(*),
                       INTEGER STRIDE(*), INTEGER IMAP(*),
                       INTEGER*2 I2VALS(*))
INTEGER FUNCTION NF_PUT_VARM_INT   (INTEGER NCID, INTEGER VARID,
                       INTEGER START(*), INTEGER COUNT(*),
                       INTEGER STRIDE(*), INTEGER IMAP(*),
                       INTEGER IVALS(*))
INTEGER FUNCTION NF_PUT_VARM_REAL  (INTEGER NCID, INTEGER VARID,
                       INTEGER START(*), INTEGER COUNT(*),
                       INTEGER STRIDE(*), INTEGER IMAP(*),
                       REAL RVALS(*))
INTEGER FUNCTION NF_PUT_VARM_DOUBLE(INTEGER NCID, INTEGER VARID,
                       INTEGER START(*), INTEGER COUNT(*),
                       INTEGER STRIDE(*), INTEGER IMAP(*),
                       DOUBLE DVALS(*))

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

START A vector of integers specifying the index in the variable where the first of the data values will be written. The indices are relative to 1, so for example, the first data value of a variable would have index (1, 1, ..., 1). The elements of START correspond, in order, to the variable’s dimensions. Hence, if the variable is a record variable, the last index would correspond to the starting record number for writing the data values.

COUNT A vector of integers specifying the number of indices selected along each dimension. To write a single value, for example, specify COUNT as (1, 1, ..., 1). The elements of COUNT correspond, in order, to the variable’s dimensions. Hence, if the variable is a record variable, the last element of COUNT corresponds to a count of the number of records to write. Note: setting any element of the count array to zero causes the function to exit without error, and without doing anything.

STRIDE A vector of integers that specifies the sampling interval along each dimension of the netCDF variable. The elements of the stride vector correspond, in order, to the netCDF variable’s dimensions (STRIDE(1) gives the sampling interval along the most rapidly varying dimension of the netCDF variable). Sampling intervals are specified in type-independent units of elements (a value of 1 selects consecutive elements of the netCDF variable along the corresponding dimension, a value of 2 selects every other element, etc.).

IMAP A vector of integers that specifies the mapping between the dimensions of a netCDF variable and the in-memory structure of the internal data array. The elements of the index mapping vector correspond, in order, to the netCDF variable’s dimensions (IMAP(1) gives the distance between elements of the internal array corresponding to the most rapidly varying dimension of the netCDF variable). Distances between elements are specified in units of elements (the distance between internal elements that occupy adjacent memory locations is 1 and not the element’s byte-length as in netCDF 2).

TEXT
I1VALS
I2VALS
IVALS
RVALS

DVALS The data values to be written. The data should be of the type appropriate for the function called. You cannot put CHARACTER data into a numeric variable or numeric data into a text variable. For numeric data, if the type of data differs from the netCDF variable type, type conversion will occur (see Section “Type Conversion” in The NetCDF Users Guide).

Errors

NF PUT VARM type returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• The specified START, COUNT, and STRIDE generate an index which is out of range. Note that no error checking is possible on the imap vector.

• One or more of the specified values are out of the range of values representable by the external data type of the variable.

• The specified netCDF is in define mode rather than data mode.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

The following IMAP vector maps in the trivial way a 2x3x4 netCDF variable and an internal array of the same shape:

     REAL A(2,3,4)       ! same shape as netCDF variable
     INTEGER IMAP(3)
     DATA IMAP /1, 2, 6/ ! netCDF dimension       inter-element distance
                         ! ----------------       ----------------------
                         ! most rapidly varying       1
                         ! intermediate               2 (=IMAP(1)*2)
                         ! most slowly varying        6 (=IMAP(2)*3)

Using the IMAP vector above with NF PUT VARM REAL obtains the same result as simply using NF PUT VAR REAL.

Here is an example of using NF PUT VARM REAL to write – from a transposed, internal array – a netCDF variable named rh which is described by the FORTRAN declaration REAL RH(4,6) (note the size and order of the dimensions):

     INCLUDE 'netcdf.inc'
        ...
     PARAMETER (NDIM=2)   ! rank of netCDF variable
     INTEGER NCID         ! netCDF ID
     INTEGER STATUS       ! return code
     INTEGER RHID         ! variable ID
     INTEGER START(NDIM)  ! netCDF variable start point

INTEGER COUNT(NDIM)  ! size of internal array
INTEGER STRIDE(NDIM) ! netCDF variable subsampling intervals
INTEGER IMAP(NDIM)   ! internal array inter-element distances
REAL RH(6,4)         ! note transposition of netCDF variable dimensions
DATA START   /1, 1/  ! start at first netCDF variable element
DATA COUNT   /4, 6/  ! entire netCDF variable; order corresponds
                     ! to netCDF variable -- not internal array
DATA STRIDE /1, 1/   ! sample every netCDF element
DATA IMAP   /6, 1/   ! would be /1, 4/ if not transposing
STATUS = NF_OPEN('foo.nc', NF_WRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID(NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_PUT_VARM_REAL(NCID, RHID, START, COUNT, STRIDE, IMAP, RH)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

Here is another example of using NF PUT VARM REAL to write – from a transposed, internal array – a subsample of the same netCDF variable, by writing every other point of the netCDF variable:

INCLUDE 'netcdf.inc'
   ...
PARAMETER (NDIM=2)   ! rank of netCDF variable
INTEGER NCID         ! netCDF dataset ID
INTEGER STATUS       ! return code
INTEGER RHID         ! variable ID
INTEGER START(NDIM)  ! netCDF variable start point
INTEGER COUNT(NDIM)  ! size of internal array
INTEGER STRIDE(NDIM) ! netCDF variable subsampling intervals
INTEGER IMAP(NDIM)   ! internal array inter-element distances
REAL RH(3,2)         ! note transposition of (subsampled) dimensions
DATA START   /1, 1/  ! start at first netCDF variable value
DATA COUNT   /2, 3/  ! order of (subsampled) dimensions corresponds
                     ! to netCDF variable -- not internal array
DATA STRIDE /2, 2/   ! sample every other netCDF element
DATA IMAP   /3, 1/   ! would be `1, 2' if not transposing
   ...
STATUS = NF_OPEN('foo.nc', NF_WRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID(NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_PUT_VARM_REAL(NCID, RHID, START, COUNT, STRIDE, IMAP, RH)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.21 NF GET VAR1 type

The functions NF GET VAR1 type get a single data value from a variable of an open netCDF dataset that is in data mode. Inputs are the netCDF ID, the variable ID, a multidimensional index that specifies which value to get, and the address of a location into which the data value will be read. The value is converted from the external data type of the variable, if necessary.

Usage

     INTEGER FUNCTION  NF_GET_VAR1_TEXT(INTEGER NCID, INTEGER VARID,
                                        INTEGER INDEX(*), CHARACTER CHVAL)
     INTEGER FUNCTION  NF_GET_VAR1_INT1(INTEGER NCID, INTEGER VARID,
                                        INTEGER INDEX(*), INTEGER*1 I1VAL)
     INTEGER FUNCTION  NF_GET_VAR1_INT2(INTEGER NCID, INTEGER VARID,
                                        INTEGER INDEX(*), INTEGER*2 I2VAL)
     INTEGER FUNCTION  NF_GET_VAR1_INT (INTEGER NCID, INTEGER VARID,
                                        INTEGER INDEX(*), INTEGER   IVAL)
     INTEGER FUNCTION  NF_GET_VAR1_REAL(INTEGER NCID, INTEGER VARID,
                                        INTEGER INDEX(*), REAL      RVAL)
     INTEGER FUNCTION  NF_GET_VAR1_DOUBLE(INTEGER NCID, INTEGER VARID,
                                        INTEGER INDEX(*), DOUBLE    DVAL)
     INTEGER FUNCTION  NF_GET_VAR1(INTEGER NCID, INTEGER VARID,
                                        INTEGER INDEX(*), VAL)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

INDEX The index of the data value to be read. The indices are relative to 1, so for example, the first data value of a two-dimensional variable has index (1,1). The elements of index correspond to the variable’s dimensions. Hence, if the variable is a record variable, the last index is the record number.

CHVAL
I1VAL
I2VAL
IVAL
RVAL
DVAL

VAL The location into which the data value will be read. You cannot get CHARAC-TER data from a numeric variable or numeric data from a character variable. For numeric data, if the type of data differs from the netCDF variable type, type conversion will occur. (see Section “Type Conversion” in The NetCDF Users Guide).

Errors

NF GET VAR1 type returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• The specified indices were out of range for the rank of the specified variable. For example, a negative index or an index that is larger than the corresponding dimension length will cause an error.

• The value is out of the range of values representable by the desired data type.

• The specified netCDF is in define mode rather than data mode.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF GET VAR1 DOUBLE to get the (4,3,2) element of the variable named rh in an existing netCDF dataset named foo.nc. For simplicity in this example, we assume that we know that rh is dimensioned with lon, lat, and time, so we want to get the value of rh that corresponds to the fourth lon value, the third lat value, and the second time value:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER STATUS, NCID
     INTEGER RHID           ! variable ID
     INTEGER RHINDX(3)      ! where to get value
     DOUBLE PRECISION RHVAL ! put it here
     DATA RHINDX /4, 3, 2/
        ...
     STATUS = NF_OPEN ('foo.nc', NF_NOWRITE, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     STATUS = NF_INQ_VARID (NCID, 'rh', RHID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
     STATUS = NF_GET_VAR1_DOUBLE (NCID, RHID, RHINDX, RHVAL)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.22 NF GET VAR type

The members of the NF GET VAR type family of functions read all the values from a netCDF variable of an open netCDF dataset. This is the simplest interface to use for reading the value of a scalar variable or when all the values of a multidimensional variable can be read at once. The values are read into consecutive locations with the first dimension varying fastest. The netCDF dataset must be in data mode.

Take care when using the simplest forms of this interface with record variables when you don’t specify how many records are to be read. If you try to read all the values of a record variable into an array but there are more records in the file than you assume, more data will be read than you expect, which may cause a segmentation violation.

Usage

     INTEGER FUNCTION NF_GET_VAR_TEXT  (INTEGER NCID, INTEGER VARID,
                                        CHARACTER*(*) text)

     INTEGER FUNCTION NF_GET_VAR_INT1  (INTEGER NCID, INTEGER VARID,
                                        INTEGER*1 i1vals(*))
     INTEGER FUNCTION NF_GET_VAR_INT2  (INTEGER NCID, INTEGER VARID,
                                        INTEGER*2 i2vals(*))
     INTEGER FUNCTION NF_GET_VAR_INT   (INTEGER NCID, INTEGER VARID,
                                        INTEGER ivals(*))
     INTEGER FUNCTION NF_GET_VAR_REAL  (INTEGER NCID, INTEGER VARID,
                                        REAL rvals(*))
     INTEGER FUNCTION NF_GET_VAR_DOUBLE(INTEGER NCID, INTEGER VARID,
                                        DOUBLE dvals(*))
     INTEGER FUNCTION NF_GET_VAR       (INTEGER NCID, INTEGER VARID,
                                        vals(*))

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

TEXT
I1VALS
I2VALS
IVALS
RVALS
DVALS

VALS The block of data values to be read. The data should be of the type appropriate for the function called. You cannot read CHARACTER data from a numeric variable or numeric data from a text variable. For numeric data, if the type of data differs from the netCDF variable type, type conversion will occur (see Section “Type Conversion” in The NetCDF Users Guide).

Errors

NF GET VAR type returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• One or more of the values are out of the range of values representable by the desired type.

• The specified netCDF is in define mode rather than data mode.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF GET VAR DOUBLE to read all the values of the variable named rh from an existing netCDF dataset named foo.nc. For simplicity in this example, we assume that we know that rh is dimensioned with lon, lat, and time, and that there are ten lon values, five lat values, and three time values.

     INCLUDE 'netcdf.inc'
        ...
     PARAMETER (TIMES=3, LATS=5, LONS=10) ! dimension lengths

INTEGER STATUS, NCID
INTEGER RHID                         ! variable ID
DOUBLE RHVALS(LONS, LATS, TIMES)
   ...
STATUS = NF_OPEN ('foo.nc', NF_NOWRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID (NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_GET_VAR_DOUBLE (NCID, RHID, RHVALS)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.23 NF GET VARA type

The members of the NF GET VARA type family of functions read an array of values from a netCDF variable of an open netCDF dataset. The array is specified by giving a corner and a vector of edge lengths. The values are read into consecutive locations with the first dimension varying fastest. The netCDF dataset must be in data mode.

Usage

INTEGER FUNCTION NF_GET_VARA_TEXT(INTEGER NCID, INTEGER VARID,
                                  INTEGER START(*), INTEGER COUNT(*),
                                  CHARACTER*(*) text)
INTEGER FUNCTION NF_GET_VARA_INT1(INTEGER NCID, INTEGER VARID,
                                  INTEGER START(*), INTEGER COUNT(*),
                                  INTEGER*1 i1vals(*))
INTEGER FUNCTION NF_GET_VARA_INT2(INTEGER NCID, INTEGER VARID,
                                  INTEGER START(*), INTEGER COUNT(*),
                                  INTEGER*2 i2vals(*))
INTEGER FUNCTION NF_GET_VARA_INT (INTEGER NCID, INTEGER VARID,
                                  INTEGER START(*), INTEGER COUNT(*),
                                  INTEGER ivals(*))
INTEGER FUNCTION NF_GET_VARA_REAL(INTEGER NCID, INTEGER VARID,
                                  INTEGER START(*), INTEGER COUNT(*),
                                  REAL rvals(*))
INTEGER FUNCTION NF_GET_VARA_DOUBLE(INTEGER NCID, INTEGER VARID,
                                  INTEGER START(*), INTEGER COUNT(*),
                                  DOUBLE dvals(*))

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

START A vector of integers specifying the index in the variable where the first of the data values will be read. The indices are relative to 1, so for example, the first data value of a variable would have index (1, 1, ..., 1). The length of START must be the same as the number of dimensions of the specified variable. The elements of START correspond, in order, to the variable’s dimensions. Hence, if the variable is a record variable, the last index would correspond to the starting record number for reading the data values.

COUNT A vector of integers specifying the edge lengths along each dimension of the block of data values to be read. To read a single value, for example, specify COUNT as (1, 1, ..., 1). The length of COUNT is the number of dimensions of the specified variable. The elements of COUNT correspond, in order, to the variable’s dimensions. Hence, if the variable is a record variable, the last element of COUNT corresponds to a count of the number of records to read. Note: setting any element of the count array to zero causes the function to exit without error, and without doing anything.

text
i1vals
i2vals
ivals
rvals

dvals The block of data values to be read. The data should be of the type appropriate for the function called. You cannot read CHARACTER data from a numeric variable or numeric data from a text variable. For numeric data, if the type of data differs from the netCDF variable type, type conversion will occur (see Section “Type Conversion” in The NetCDF Users Guide).

Errors

NF GET VARA type returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• The specified corner indices were out of range for the rank of the specified variable. For example, a negative index or an index that is larger than the corresponding dimension length will cause an error.

• The specified edge lengths added to the specified corner would have referenced data out of range for the rank of the specified variable. For example, an edge length that is larger than the corresponding dimension length minus the corner index will cause an error.

• One or more of the values are out of the range of values representable by the desired type.

• The specified netCDF is in define mode rather than data mode.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF GET VARA DOUBLE to read all the values of the variable named rh from an existing netCDF dataset named foo.nc. For simplicity in this example, we assume that we know that rh is dimensioned with lon, lat, and time, and that there are ten lon values, five lat values, and three time values.

     INCLUDE 'netcdf.inc'

   ...
PARAMETER (NDIMS=3)                  ! number of dimensions
PARAMETER (TIMES=3, LATS=5, LONS=10) ! dimension lengths
INTEGER STATUS, NCID
INTEGER RHID                         ! variable ID
INTEGER START(NDIMS), COUNT(NDIMS)
DOUBLE RHVALS(LONS, LATS, TIMES)
DATA START /1, 1, 1/                 ! start at first value
DATA COUNT /LONS, LATS, TIMES/       ! get all the values
   ...
STATUS = NF_OPEN ('foo.nc', NF_NOWRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID (NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_GET_VARA_DOUBLE (NCID, RHID, START, COUNT, RHVALS)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.24 NF GET VARS type

The NF GET VARS type family of functions read a subsampled (strided) array section of values from a netCDF variable of an open netCDF dataset. The subsampled array section is specified by giving a corner, a vector of edge lengths, and a stride vector. The values are read with the first dimension of the netCDF variable varying fastest. The netCDF dataset must be in data mode.

Usage

INTEGER FUNCTION NF_GET_VARS_TEXT  (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*),CHARACTER*(*) text)
INTEGER FUNCTION NF_GET_VARS_INT1  (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*),INTEGER*1 i1vals(*))
INTEGER FUNCTION NF_GET_VARS_INT2  (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*),INTEGER*2 i2vals(*))
INTEGER FUNCTION NF_GET_VARS_INT   (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*), INTEGER ivals(*))
INTEGER FUNCTION NF_GET_VARS_REAL  (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*), REAL rvals(*))
INTEGER FUNCTION NF_GET_VARS_DOUBLE(INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*), DOUBLE dvals(*))

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

START A vector of integers specifying the index in the variable from which the first of the data values will be read. The indices are relative to 1, so for example, the first data value of a variable would have index (1, 1, ..., 1). The elements of START correspond, in order, to the variable’s dimensions. Hence, if the variable is a record variable, the last index would correspond to the starting record number for reading the data values.

COUNT A vector of integers specifying the number of indices selected along each dimension. To read a single value, for example, specify COUNT as (1, 1, ..., 1). The elements of COUNT correspond, in order, to the variable’s dimensions. Hence, if the variable is a record variable, the last element of COUNT corresponds to a count of the number of records to read. Note: setting any element of the count array to zero causes the function to exit without error, and without doing anything.

STRIDE A vector of integers specifying, for each dimension, the interval between selected indices or the value 0. The elements of the vector correspond, in order, to the variable’s dimensions. A value of 1 accesses adjacent values of the netCDF variable in the corresponding dimension; a value of 2 accesses every other value of the netCDF variable in the corresponding dimension; and so on. A 0 argument is treated as (1, 1, ..., 1).

text
i1vals
i2vals
ivals
rvals

dvals The block of data values to be read. The data should be of the type appropriate for the function called. You cannot read CHARACTER data from a numeric variable or numeric data from a text variable. For numeric data, if the type of data differs from the netCDF variable type, type conversion will occur (see Section “Type Conversion” in The NetCDF Users Guide).

Errors

NF GET VARS type returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• The specified start, count and stride generate an index which is out of range.

• One or more of the values are out of the range of values representable by the desired type.

• The specified netCDF is in define mode rather than data mode.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF GET VARS DOUBLE to read every other value in each dimension of the variable named rh from an existing netCDF dataset named foo.nc. Values are assigned, using the same dimensional strides, to a 2-parameter array. For simplicity in this example, we assume that we know that rh is dimensioned with lon, lat, and time, and that there are ten lon values, five lat values, and three time values.

INCLUDE 'netcdf.inc'
   ...
PARAMETER (NDIMS=3)                  ! number of dimensions
PARAMETER (TIMES=3, LATS=5, LONS=10) ! dimension lengths
INTEGER STATUS, NCID
INTEGER RHID ! variable ID
INTEGER START(NDIMS), COUNT(NDIMS), STRIDE(NDIMS)
DOUBLE DATA(LONS, LATS, TIMES)
DATA START /1, 1, 1/                 ! start at first value
DATA COUNT /LONS, LATS, TIMES/
DATA STRIDE /2, 2, 2/
   ...
STATUS = NF_OPEN ('foo.nc', NF_NOWRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID (NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_GET_VARS_DOUBLE(NCID,RHID,START,COUNT,STRIDE,DATA(1,1,1))
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.25 NF GET VARM type

The NF GET VARM type family of functions reads a mapped array section of values from a netCDF variable of an open netCDF dataset. The mapped array section is specified by giving a corner, a vector of edge lengths, a stride vector, and an index mapping vector. The index mapping vector is a vector of integers that specifies the mapping between the dimensions of a netCDF variable and the in-memory structure of the internal data array. No assumptions are made about the ordering or length of the dimensions of the data array. The netCDF dataset must be in data mode.

Usage

INTEGER FUNCTION NF_GET_VARM_TEXT  (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*), INTEGER IMAP(*),
                            CHARACTER*(*) text)
INTEGER FUNCTION NF_GET_VARM_INT1  (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*), INTEGER IMAP(*),
                            INTEGER*1 i1vals(*))
INTEGER FUNCTION NF_GET_VARM_INT2  (INTEGER NCID, INTEGER VARID,
                            INTEGER START(*), INTEGER COUNT(*),
                            INTEGER STRIDE(*), INTEGER IMAP(*),
                            INTEGER*2 i2vals(*))
INTEGER FUNCTION NF_GET_VARM_INT   (INTEGER NCID, INTEGER VARID,

                       INTEGER START(*), INTEGER COUNT(*),
                       INTEGER STRIDE(*), INTEGER IMAP(*),
                       INTEGER ivals(*))
INTEGER FUNCTION NF_GET_VARM_REAL  (INTEGER NCID, INTEGER VARID,
                       INTEGER START(*), INTEGER COUNT(*),
                       INTEGER STRIDE(*), INTEGER IMAP(*),
                       REAL rvals(*))
INTEGER FUNCTION NF_GET_VARM_DOUBLE(INTEGER NCID, INTEGER VARID,
                       INTEGER START(*), INTEGER COUNT(*),
                       INTEGER STRIDE(*), INTEGER IMAP(*),
                       DOUBLE dvals(*))

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

START A vector of integers specifying the index in the variable from which the first of the data values will be read. The indices are relative to 1, so for example, the first data value of a variable would have index (1, 1, ..., 1). The elements of START correspond, in order, to the variable’s dimensions. Hence, if the variable is a record variable, the last index would correspond to the starting record number for reading the data values.

COUNT A vector of integers specifying the number of indices selected along each dimension. To read a single value, for example, specify COUNT as (1, 1, ..., 1). The elements of COUNT correspond, in order, to the variable’s dimensions. Hence, if the variable is a record variable, the last element of COUNT corresponds to a count of the number of records to read. Note: setting any element of the count array to zero causes the function to exit without error, and without doing anything.

STRIDE A vector of integers specifying, for each dimension, the interval between selected indices or the value 0. The elements of the vector correspond, in order, to the variable’s dimensions. A value of 1 accesses adjacent values of the netCDF variable in the corresponding dimension; a value of 2 accesses every other value of the netCDF variable in the corresponding dimension; and so on. A 0 argument is treated as (1, 1, ..., 1).

IMAP A vector of integers that specifies the mapping between the dimensions of a netCDF variable and the in-memory structure of the internal data array. IMAP(1) gives the distance between elements of the internal array corresponding to the most rapidly varying dimension of the netCDF variable. IMAP(N) (where N is the rank of the netCDF variable) gives the distance between elements of the internal array corresponding to the most slowly varying dimension of the netCDF variable. Intervening IMAP elements correspond to other dimensions of the netCDF variable in the obvious way. Distances between elements are specified in units of elements (the distance between internal elements that occupy adjacent memory locations is 1 and not the element’s byte-length as in netCDF 2).

text
i1vals
i2vals
ivals
rvals

dvals The block of data values to be read. The data should be of the type appropriate for the function called. You cannot read CHARACTER data from a numeric variable or numeric data from a text variable. For numeric data, if the type of data differs from the netCDF variable type, type conversion will occur (see Section “Type Conversion” in The NetCDF Users Guide).

Errors

NF GET VARM type returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• The specified START, COUNT, and STRIDE generate an index which is out of range. Note that no error checking is possible on the imap vector.

• One or more of the values are out of the range of values representable by the desired type.

• The specified netCDF is in define mode rather than data mode.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

The following IMAP vector maps in the trivial way a 2x3x4 netCDF variable and an internal array of the same shape:

     REAL A(2,3,4)       ! same shape as netCDF variable
     INTEGER IMAP(3)
     DATA IMAP /1, 2, 6/ ! netCDF dimension       inter-element distance
                         ! ----------------       ----------------------
                         ! most rapidly varying       1
                         ! intermediate               2 (=IMAP(1)*2)
                         ! most slowly varying        6 (=IMAP(2)*3)

Using the IMAP vector above with NF GET VARM REAL obtains the same result as simply using NF GET VAR REAL.

Here is an example of using NF GET VARM REAL to transpose a netCDF variable named rh which is described by the FORTRAN declaration REAL RH(4,6) (note the size and order of the dimensions):

     INCLUDE 'netcdf.inc'
        ...
     PARAMETER (NDIM=2)   ! rank of netCDF variable
     INTEGER NCID         ! netCDF dataset ID
     INTEGER STATUS       ! return code
     INTEGER RHID         ! variable ID
     INTEGER START(NDIM)  ! netCDF variable start point

INTEGER COUNT(NDIM)  ! size of internal array
INTEGER STRIDE(NDIM) ! netCDF variable subsampling intervals
INTEGER IMAP(NDIM)   ! internal array inter-element distances
REAL    RH(6,4)      ! note transposition of netCDF variable dimensions
DATA START   /1, 1/  ! start at first netCDF variable element
DATA COUNT   /4, 6/  ! entire netCDF variable; order corresponds
                     ! to netCDF variable -- not internal array
DATA STRIDE /1, 1/   ! sample every netCDF element
DATA IMAP   /6, 1/   ! would be /1, 4/ if not transposing
   ...
STATUS = NF_OPEN('foo.nc', NF_NOWRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID(NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_GET_VARM_REAL(NCID, RHID, START, COUNT, STRIDE, IMAP, RH)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

Here is another example of using NF GET VARM REAL to simultaneously transpose and subsample the same netCDF variable, by accessing every other point of the netCDF variable:

INCLUDE 'netcdf.inc'
   ...
PARAMETER (NDIM=2)   ! rank of netCDF variable
INTEGER NCID         ! netCDF dataset ID
INTEGER STATUS       ! return code
INTEGER RHID         ! variable ID
INTEGER START(NDIM)  ! netCDF variable start point
INTEGER COUNT(NDIM)  ! size of internal array
INTEGER STRIDE(NDIM) ! netCDF variable subsampling intervals
INTEGER IMAP(NDIM)   ! internal array inter-element distances
REAL    RH(3,2)      ! note transposition of (subsampled) dimensions
DATA START   /1, 1/  ! start at first netCDF variable value
DATA COUNT   /2, 3/  ! order of (subsampled) dimensions corresponds
                     ! to netCDF variable -- not internal array
DATA STRIDE /2, 2/   ! sample every other netCDF element
DATA IMAP   /3, 1/   ! would be `1, 2' if not transposing
   ...
STATUS = NF_OPEN('foo.nc', NF_NOWRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID(NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_GET_VARM_REAL(NCID, RHID, START, COUNT, STRIDE, IMAP, RH)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.26 Reading and Writing Character String Values

Character strings are not a primitive netCDF external data type, in part because FOR-TRAN does not support the abstraction of variable-length character strings (the FORTRAN LEN function returns the static length of a character string, not its dynamic length). As a result, a character string cannot be written or read as a single object in the netCDF interface. Instead, a character string must be treated as an array of characters, and array access must be used to read and write character strings as variable data in netCDF datasets. Furthermore, variable-length strings are not supported by the netCDF interface except by convention; for example, you may treat a zero byte as terminating a character string, but you must explicitly specify the length of strings to be read from and written to netCDF variables.

Character strings as attribute values are easier to use, since the strings are treated as a single unit for access. However, the value of a character-string attribute is still an array of characters with an explicit length that must be specified when the attribute is defined.

When you define a variable that will have character-string values, use a character-position dimension as the most quickly varying dimension for the variable (the first dimension for the variable in FORTRAN). The length of the character-position dimension will be the maximum string length of any value to be stored in the character-string variable. Space for maximum-length strings will be allocated in the disk representation of character-string variables whether you use the space or not. If two or more variables have the same maximum length, the same character-position dimension may be used in defining the variable shapes.

To write a character-string value into a character-string variable, use either entire variable access or array access. The latter requires that you specify both a corner and a vector of edge lengths. The character-position dimension at the corner should be one for FOR-TRAN. If the length of the string to be written is n, then the vector of edge lengths will specify n in the character-position dimension, and one for all the other dimensions:(n, 1, 1, ..., 1).

In FORTRAN, fixed-length strings may be written to a netCDF dataset without a terminating character, to save space. Variable-length strings should follow the C convention of writing strings with a terminating zero byte so that the intended length of the string can be determined when it is later read by either C or FORTRAN programs.

The FORTRAN interface for reading and writing strings requires the use of different functions for accessing string values and numeric values, because standard FORTRAN does not permit the same formal parameter to be used for both character values and numeric values. An additional argument, specifying the declared length of the character string passed as a value, is required for NF PUT VARA TEXT and NF GET VARA TEXT. The actual length of the string is specified as the value of the edge-length vector corresponding to the character-position dimension.

Here is an example that defines a record variable, tx, for character strings and stores a character-string value into the third record using NF PUT VARA TEXT. In this example, we assume the string variable and data are to be added to an existing netCDF dataset named foo.nc that already has an unlimited record dimension time.

     INCLUDE 'netcdf.inc'
        ...
     INTEGER   TDIMS, TXLEN

PARAMETER (TDIMS=2)    ! number of TX dimensions
PARAMETER (TXLEN = 15) ! length of example string
INTEGER  NCID
INTEGER  CHID          ! char position dimension id
INTEGER  TIMEID        ! record dimension id
INTEGER  TXID          ! variable ID
INTEGER  TXDIMS(TDIMS) ! variable shape
INTEGER  TSTART(TDIMS), TCOUNT(TDIMS)
CHARACTER*40 TXVAL     ! max length 40
DATA TXVAL /'example string'/
   ...
TXVAL(TXLEN:TXLEN) = CHAR(0)   ! null terminate
   ...
STATUS = NF_OPEN('foo.nc', NF_WRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_REDEF(NCID)        ! enter define mode
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
! define character-position dimension for strings of max length 40
STATUS = NF_DEF_DIM(NCID, "chid", 40, CHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
! define a character-string variable
TXDIMS(1) = CHID   ! character-position dimension first
TXDIMS(2) = TIMEID
STATUS = NF_DEF_VAR(NCID, "tx", NF_CHAR, TDIMS, TXDIMS, TXID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_ENDDEF(NCID) ! leave define mode
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
! write txval into tx netCDF variable in record 3
TSTART(1) = 1      ! start at beginning of variable
TSTART(2) = 3      ! record number to write
TCOUNT(1) = TXLEN  ! number of chars to write
TCOUNT(2) = 1      ! only write one record
STATUS = NF_PUT_VARA_TEXT (NCID, TXID, TSTART, TCOUNT, TXVAL)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.27 Fill Values

What happens when you try to read a value that was never written in an open netCDF dataset? You might expect that this should always be an error, and that you should get an error message or an error status returned. You do get an error if you try to read data from a netCDF dataset that is not open for reading, if the variable ID is invalid for the specified netCDF dataset, or if the specified indices are not properly within the range defined by the dimension lengths of the specified variable. Otherwise, reading a value that was not written returns a special fill value used to fill in any undefined values when a netCDF variable is first written.

You may ignore fill values and use the entire range of a netCDF external data type, but in this case you should make sure you write all data values before reading them. If you know you will be writing all the data before reading it, you can specify that no prefilling of variables with fill values will occur by calling NF SET FILL before writing. This may provide a significant performance gain for netCDF writes.

The variable attribute FillValue may be used to specify the fill value for a variable. Their are default fill values for each type, defined in the include file netcdf.inc: NF FILL CHAR, NF FILL INT1 (same as NF FILL BYTE), NF FILL INT2 (same as NF FILL SHORT), NF FILL INT, NF FILL REAL (same as NF FILL FLOAT), and NF FILL DOUBLE.

The netCDF byte and character types have different default fill values. The default fill value for characters is the zero byte, a useful value for detecting the end of variable-length C character strings. If you need a fill value for a byte variable, it is recommended that you explicitly define an appropriate FillValue attribute, as generic utilities such as ncdump will not assume a default fill value for byte variables.

Type conversion for fill values is identical to type conversion for other values: attempting to convert a value from one type to another type that can’t represent the value results in a range error. Such errors may occur on writing or reading values from a larger type (such as double) to a smaller type (such as float), if the fill value for the larger type cannot be represented in the smaller type.

6.28 NF RENAME VAR

The function NF RENAME VAR changes the name of a netCDF variable in an open netCDF dataset. If the new name is longer than the old name, the netCDF dataset must be in define mode. You cannot rename a variable to have the name of any existing variable.

Usage

     INTEGER FUNCTION NF_RENAME_VAR (INTEGER NCID, INTEGER VARID,
                                     CHARACTER*(*) NEWNAM)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID.

NAME New name for the specified variable.

Errors

NF RENAME VAR returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The new name is in use as the name of another variable.

• The variable ID is invalid for the specified netCDF dataset.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF RENAME VAR to rename the variable rh to rel hum in an existing netCDF dataset named foo.nc:

     INCLUDE 'netcdf.inc'
        ...
     INTEGER  STATUS, NCID
     INTEGER  RHID             ! variable ID
        ...
     STATUS = NF_OPEN ('foo.nc', NF_WRITE, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     STATUS = NF_REDEF (NCID)  ! enter definition mode
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
     STATUS = NF_INQ_VARID (NCID, 'rh', RHID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
     STATUS = NF_RENAME_VAR (NCID, RHID, 'rel_hum')
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
     STATUS = NF_ENDDEF (NCID) ! leave definition mode
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

6.29 Change between Collective and Independent Parallel Access: NF VAR PAR ACCESS

The function NF VAR PAR ACCESS changes whether read/write operations on a parallel file system are performed collectively (the default) or independently on the variable. This function can only be called if the file was created with NF CREATE PAR (see Section 2.7 [NF CREATE PAR], page 12) or opened with NF OPEN PAR (see Section 2.10 [NF OPEN PAR], page 16).

This function is only available if the netCDF library was built with a HDF5 library for which –enable-parallel was used, and which was linked (like HDF5) to MPI libraries.

Calling this function affects only the open file - information about whether a variable is to be accessed collectively or independently is not written to the data file. Every time you open a file on a parallel file system, all variables default to collective operations. The change a variable to independent lasts only as long as that file is open.

The variable can be changed from collective to independent, and back, as often as desired.

Usage

     INTEGER NF_VAR_PAR_ACCESS(INTEGER NCID, INTEGER VARID, INTEGER ACCESS);

NCID NetCDF ID, from a previous call to NF OPEN PAR (see Section 2.10 [NF OPEN PAR], page 16) or NF CREATE PAR (see Section 2.7 [NF CREATE PAR], page 12).

varid Variable ID.

access NF INDEPENDENT to set this variable to independent operations. NF COLLECTIVE to set it to collective operations.

Return Values

NF_NOERR No error.
NF_ENOTVAR

No variable found.

NF_ENOTNC4

Not a netCDF-4 file.

NF_NOPAR File not opened for parallel access.

Example

This example comes from test program nf test/ftst parallel.F. For this test to be run, netCDF must have been built with a parallel-enabled HDF5, and –enable-parallel-tests must have been used when configuring netcdf.

           retval = nf_var_par_access(ncid, varid, nf_collective)
           if (retval .ne. nf_noerr) stop 2

7 Attributes

7.1 Attributes Introduction

Attributes may be associated with each netCDF variable to specify such properties as units, special values, maximum and minimum valid values, scaling factors, and offsets. Attributes for a netCDF dataset are defined when the dataset is first created, while the netCDF dataset is in define mode. Additional attributes may be added later by reentering define mode. A netCDF attribute has a netCDF variable to which it is assigned, a name, a type, a length, and a sequence of one or more values. An attribute is designated by its variable ID and name. When an attribute name is not known, it may be designated by its variable ID and number in order to determine its name, using the function NF INQ ATTNAME.

THE attributes associated with a variable are typically defined immediately after the variable is created, while still in define mode. The data type, length, and value of an attribute may be changed even when in data mode, as long as the changed attribute requires no more space than the attribute as originally defined.

It is also possible to have attributes that are not associated with any variable. These are called global attributes and are identified by using NF GLOBAL as a variable pseudo-ID. Global attributes are usually related to the netCDF dataset as a whole and may be used for purposes such as providing a title or processing history for a netCDF dataset.

Operations supported on attributes are:

• Create an attribute, given its variable ID, name, data type, length, and value.

• Get attribute’s data type and length from its variable ID and name.

• Get attribute’s value from its variable ID and name.

• Copy attribute from one netCDF variable to another.

• Get name of attribute from its number.

• Rename an attribute.

• Delete an attribute.

7.2 Attribute Conventions

Names commencing with underscore (’ ’) are reserved for use by the netCDF library. Most generic applications that process netCDF datasets assume standard attribute conventions and it is strongly recommended that these be followed unless there are good reasons for not doing so. Below we list the names and meanings of recommended standard attributes that have proven useful. Note that some of these (e.g. units, valid range, scale factor) assume numeric data and should not be used with character data. units

A character string that specifies the units used for the variable’s data. Unidata has developed a freely-available library of routines to convert between character string and binary forms of unit specifications and to perform various useful operations on the binary forms. This library is used in some netCDF applications. Using the recommended units syntax permits data represented in conformable units to be automatically converted to common units for arithmetic operations. See Section “Units” in The NetCDF Users Guide.

long_name

A long descriptive name. This could be used for labeling plots, for example. If a variable has no long name attribute assigned, the variable name should be used as a default.

valid_min

A scalar specifying the minimum valid value for this variable.

valid_max

A scalar specifying the maximum valid value for this variable.

valid_range

A vector of two numbers specifying the minimum and maximum valid values for this variable, equivalent to specifying values for both valid min and valid max attributes. Any of these attributes define the valid range. The attribute valid range must not be defined if either valid min or valid max is defined. Generic applications should treat values outside the valid range as missing. The type of each valid range, valid min and valid max attribute should match the type of its variable (except that for byte data, these can be of a signed integral type to specify the intended range). If neither valid min, valid max nor valid range is defined then generic applications should define a valid range as follows. If the data type is byte and FillValue is not explicitly defined, then the valid range should include all possible values. Otherwise, the valid range should exclude the FillValue (whether defined explicitly or by default) as follows. If the FillValue is positive then it defines a valid maximum, otherwise it defines a valid minimum. For integer types, there should be a difference of 1 between the FillValue and this valid minimum or maximum. For floating point types, the difference should be twice the minimum possible (1 in the least significant bit) to allow for rounding error.

scale_factor

If present for a variable, the data are to be multiplied by this factor after the data are read by the application that accesses the data.

add_offset

If present for a variable, this number is to be added to the data after it is read by the application that accesses the data. If both scale factor and add offset attributes are present, the data are first scaled before the offset is added. The attributes scale factor and add offset can be used together to provide simple data compression to store low-resolution floating-point data as small integers in a netCDF dataset. When scaled data are written, the application should first subtract the offset and then divide by the scale factor, rounding the result to the nearest integer to avoid a bias caused by truncation towards zero. When scale factor and add offset are used for packing, the associated variable (containing the packed data) is typically of type byte or short, whereas the unpacked values are intended to be of type float or double. The attributes scale factor and add offset should both be of the type intended for the unpacked data, e.g. float or double.

_FillValue

The FillValue attribute specifies the fill value used to pre-fill disk space allocated to the variable. Such pre-fill occurs unless nofill mode is set using NF SET FILL. See Section 2.18 [NF SET FILL], page 24. The fill value is returned when reading values that were never written. If FillValue is defined then it should be scalar and of the same type as the variable. It is not necessary to define your own FillValue attribute for a variable if the default fill value for the type of the variable is adequate. However, use of the default fill value for data type byte is not recommended. Note that if you change the value of this attribute, the changed value applies only to subsequent writes; previously written data are not changed.

Generic applications often need to write a value to represent undefined or missing values. The fill value provides an appropriate value for this purpose because it is normally outside the valid range and therefore treated as missing when read by generic applications. It is legal (but not recommended) for the fill value to be within the valid range.

See Section 6.27 [Fill Values], page 106.

missing_value

This attribute is not treated in any special way by the library or conforming generic applications, but is often useful documentation and may be used by specific applications. The missing value attribute can be a scalar or vector containing values indicating missing data. These values should all be outside the valid range so that generic applications will treat them as missing.

signedness

Deprecated attribute, originally designed to indicate whether byte values should be treated as signed or unsigned. The attributes valid min and valid max may be used for this purpose. For example, if you intend that a byte variable store only nonnegative values, you can use valid min = 0 and valid max = 255. This attribute is ignored by the netCDF library.

C_format A character array providing the format that should be used by C applications to print values for this variable. For example, if you know a variable is only accurate to three significant digits, it would be appropriate to define the C format attribute as "%.3g". The ncdump utility program uses this attribute for variables for which it is defined. The format applies to the scaled (internal) type and value, regardless of the presence of the scaling attributes scale factor and add offset.

FORTRAN_format

A character array providing the format that should be used by FORTRAN applications to print values for this variable. For example, if you know a variable is only accurate to three significant digits, it would be appropriate to define the FORTRAN format attribute as "(G10.3)".

title A global attribute that is a character array providing a succinct description of what is in the dataset. history A global attribute for an audit trail. This is a character array with a line

for each invocation of a program that has modified the dataset. Well-behaved generic netCDF applications should append a line containing: date, time of day, user name, program name and command arguments.

Conventions

If present, ’Conventions’ is a global attribute that is a character array for the name of the conventions followed by the dataset, in the form of a string that is interpreted as a directory name relative to a directory that is a repository of documents describing sets of discipline-specific conventions. This permits a hierarchical structure for conventions and provides a place where descriptions and examples of the conventions may be maintained by the defining institutions and groups. The conventions directory name is currently interpreted relative to the directory pub/netcdf/Conventions/ on the host machine ftp.unidata.ucar.edu. Alternatively, a full URL specification may be used to name a WWW site where documents that describe the conventions are maintained. For example, if a group named NUWG agrees upon a set of conventions for dimension names, variable names, required attributes, and netCDF representations for certain discipline-specific data structures, they may store a document describing the agreed-upon conventions in a dataset in the NUWG/ subdirectory of the Conventions directory. Datasets that followed these conventions would contain a global Conventions attribute with value "NUWG". Later, if the group agrees upon some additional conventions for a specific subset of NUWG data, for example time series data, the description of the additional conventions might be stored in the NUWG/Time series/ subdirectory, and datasets that adhered to these additional conventions would use the global Conventions attribute with value "NUWG/Time series", implying that this dataset adheres to the NUWG conventions and also to the additional NUWG time-series conventions.

7.3 NF PUT ATT type

The function NF PUT ATT type adds or changes a variable attribute or global attribute of an open netCDF dataset. If this attribute is new, or if the space required to store the attribute is greater than before, the netCDF dataset must be in define mode.

Usage

Although it’s possible to create attributes of all types, text and double attributes are adequate for most purposes.

INTEGER FUNCTION  NF_PUT_ATT_TEXT  (INTEGER NCID, INTEGER VARID,
                               CHARACTER*(*) NAME, INTEGER LEN,
                               CHARACTER*(*) TEXT)
INTEGER FUNCTION  NF_PUT_ATT_INT1  (INTEGER NCID, INTEGER VARID,
                               CHARACTER*(*) NAME, INTEGER XTYPE,
                               LEN, INTEGER*1 I1VALS(*))
INTEGER FUNCTION  NF_PUT_ATT_INT2  (INTEGER NCID, INTEGER VARID,
                                CHARACTER*(*) NAME, INTEGER XTYPE,

                                         LEN, INTEGER*2 I2VALS(*))
     INTEGER FUNCTION  NF_PUT_ATT_INT   (INTEGER NCID, INTEGER VARID,
                                         CHARACTER*(*) NAME, INTEGER XTYPE,
                                         LEN, INTEGER IVALS(*))
     INTEGER FUNCTION  NF_PUT_ATT_REAL  (INTEGER NCID, INTEGER VARID,
                                         CHARACTER*(*) NAME, INTEGER XTYPE,
                                         LEN, REAL RVALS(*))
     INTEGER FUNCTION  NF_PUT_ATT_DOUBLE(INTEGER NCID, INTEGER VARID,
                                         CHARACTER*(*) NAME, INTEGER XTYPE,
                                         LEN, DOUBLE DVALS(*))
     INTEGER FUNCTION  NF_PUT_ATT       (INTEGER NCID, INTEGER VARID,
                                         CHARACTER*(*) NAME, INTEGER XTYPE,
                                         LEN, * VALS(*))

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID of the variable to which the attribute will be assigned or NF GLOBAL for a global attribute.

NAME Attribute name. Attribute name conventions are assumed by some netCDF generic applications, e.g., ‘units’ as the name for a string attribute that gives the units for a netCDF variable. For examples of attribute conventions see Section 7.2 [Attribute Conventions], page 111.

XTYPE One of the set of predefined netCDF external data types. The type of this parameter, NF TYPE, is defined in the netCDF header file. The valid netCDF external data types are NF BYTE, NF CHAR, NF SHORT, NF INT, NF FLOAT, and NF DOUBLE. Although it’s possible to create attributes of all types, NF CHAR and NF DOUBLE attributes are adequate for most purposes.

LEN Number of values provided for the attribute.

TEXT
I1VALS
I2VALS
IVALS
RVALS
DVALS

VALS An array of LEN attribute values. The data should be of a type appropriate for the function called. You cannot write CHARACTER data into a numeric attribute or numeric data into a text attribute. For numeric data, if the type of data differs from the attribute type, type conversion will occur See Section “Type Conversion” in The NetCDF Users Guide.

Errors

NF PUT ATT type returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• The specified netCDF type is invalid.

• The specified length is negative.

• The specified open netCDF dataset is in data mode and the specified attribute would expand.

• The specified open netCDF dataset is in data mode and the specified attribute does not already exist.

• The specified netCDF ID does not refer to an open netCDF dataset.

• The number of attributes for this variable exceeds NF MAX ATTRS.

Example

Here is an example using NF PUT ATT DOUBLE to add a variable attribute named valid range for a netCDF variable named rh and a global attribute named title to an existing netCDF dataset named foo.nc:

INCLUDE 'netcdf.inc'
   ...
INTEGER STATUS, NCID
INTEGER RHID                 ! variable ID
DOUBLE RHRNGE(2)
DATA RHRNGE /0.0D0, 100.0D0/
   ...
STATUS = NF_OPEN ('foo.nc', NF_WRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_REDEF (NCID)     ! enter define mode
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_INQ_VARID (NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_PUT_ATT_DOUBLE (NCID, RHID, 'valid_range', NF_DOUBLE, &
                            2, RHRNGE)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_PUT_ATT_TEXT (NCID, NF_GLOBAL, 'title', 19,
                          'example netCDF dataset')
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_ENDDEF (NCID)    ! leave define mode
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

7.4 NF INQ ATT Family

This family of functions returns information about a netCDF attribute. All but one of these functions require the variable ID and attribute name; the exception is NF INQ ATTNAME. Information about an attribute includes its type, length, name, and number. See the NF GET ATT family for getting attribute values.

The function NF INQ ATTNAME gets the name of an attribute, given its variable ID and number. This function is useful in generic applications that need to get the names of all the attributes associated with a variable, since attributes are accessed by name rather than number in all other attribute functions. The number of an attribute is more volatile than the name, since it can change when other attributes of the same variable are deleted. This is why an attribute number is not called an attribute ID.

The function NF INQ ATT returns the attribute’s type and length. The other functions each return just one item of information about an attribute.

Usage

     INTEGER FUNCTION NF_INQ_ATT    (INTEGER NCID, INTEGER VARID,
                                     CHARACTER*(*) NAME, INTEGER xtype,
                                     INTEGER len)
     INTEGER FUNCTION NF_INQ_ATTTYPE(INTEGER NCID, INTEGER VARID,
                                     CHARACTER*(*) NAME, INTEGER xtype)
     INTEGER FUNCTION NF_INQ_ATTLEN (INTEGER NCID, INTEGER VARID,
                                     CHARACTER*(*) NAME, INTEGER len)
     INTEGER FUNCTION NF_INQ_ATTNAME(INTEGER NCID, INTEGER VARID,
                                     INTEGER ATTNUM, CHARACTER*(*) name)
     INTEGER FUNCTION NF_INQ_ATTID  (INTEGER NCID, INTEGER VARID,
                                     CHARACTER*(*) NAME, INTEGER attnum)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID of the attribute’s variable, or NF GLOBAL for a global attribute.

NAME Attribute name. For NF INQ ATTNAME, this is a pointer to the location for the returned attribute name.

xtype Returned attribute type, one of the set of predefined netCDF external data types. The valid netCDF external data types are NF BYTE, NF CHAR, NF SHORT, NF INT, NF FLOAT, and NF DOUBLE.

len Returned number of values currently stored in the attribute. For a string-valued attribute, this is the number of characters in the string.

attnum For NF INQ ATTNAME, the input attribute number; for NF INQ ATTID, the returned attribute number. The attributes for each variable are numbered from 1 (the first attribute) to NATTS, where NATTS is the number of attributes for the variable, as returned from a call to NF INQ VARNATTS. (If you already know an attribute name, knowing its number is not very useful, because accessing information about an attribute requires its name.)

Errors

Each function returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• The specified attribute does not exist.

• The specified netCDF ID does not refer to an open netCDF dataset.

• For NF INQ ATTNAME, the specified attribute number is negative or more than the number of attributes defined for the specified variable.

Example

Here is an example using NF INQ ATT to find out the type and length of a variable attribute named valid range for a netCDF variable named rh and a global attribute named title in an existing netCDF dataset named foo.nc:

INCLUDE 'netcdf.inc'
   ...
INTEGER STATUS, NCID
INTEGER RHID               ! variable ID
INTEGER VRLEN, TLEN        ! attribute lengths
   ...
STATUS = NF_OPEN ('foo.nc', NF_NOWRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID (NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_ATTLEN (NCID, RHID, 'valid_range', VRLEN)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_INQ_ATTLEN (NCID, NF_GLOBAL, 'title', TLEN)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

7.5 NF GET ATT type

Members of the NF GET ATT type family of functions get the value(s) of a netCDF attribute, given its variable ID and name.

Usage

INTEGER FUNCTION NF_GET_ATT_TEXT   (INTEGER NCID, INTEGER VARID,
                                    CHARACTER*(*) NAME,
                                    CHARACTER*(*) text)
INTEGER FUNCTION NF_GET_ATT_INT1   (INTEGER NCID, INTEGER VARID,
                                    CHARACTER*(*) NAME,
                                    INTEGER*1 i1vals(*))
INTEGER FUNCTION NF_GET_ATT_INT2   (INTEGER NCID, INTEGER VARID,
                                    CHARACTER*(*) NAME,
                                    INTEGER*2 i2vals(*))
INTEGER FUNCTION NF_GET_ATT_INT    (INTEGER NCID, INTEGER VARID,
                                    CHARACTER*(*) NAME,
                                    INTEGER ivals(*))
INTEGER FUNCTION NF_GET_ATT_REAL   (INTEGER NCID, INTEGER VARID,
                                    CHARACTER*(*) NAME,
                                    REAL rvals(*))
INTEGER FUNCTION NF_GET_ATT_DOUBLE (INTEGER NCID, INTEGER VARID,
                                    CHARACTER*(*) NAME,
                                    DOUBLE dvals(*))
INTEGER FUNCTION NF_GET_ATT        (INTEGER NCID, INTEGER VARID,

                                         CHARACTER*(*) NAME, * vals(*))

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID Variable ID of the attribute’s variable, or NF GLOBAL for a global attribute.

NAME Attribute name.

TEXT
I1VALS
I2VALS
IVALS
RVALS
DVALS

VALS Returned attribute values. All elements of the vector of attribute values are returned, so you must provide enough space to hold them. If you don’t know how much space to reserve, call NF INQ ATTLEN first to find out the length of the attribute. You cannot read character data from a numeric variable or numeric data from a text variable. For numeric data, if the type of data differs from the netCDF variable type, type conversion will occur. See Section “Type Conversion” in The The NetCDF Users Guide.

Errors

NF GET ATT type returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The variable ID is invalid for the specified netCDF dataset.

• The specified attribute does not exist.

• The specified netCDF ID does not refer to an open netCDF dataset.

• One or more of the attribute values are out of the range of values representable by the desired type.

Example

Here is an example using NF GET ATT DOUBLE to determine the values of a variable attribute named valid range for a netCDF variable named rh and a global attribute named title in an existing netCDF dataset named foo.nc. In this example, it is assumed that we don’t know how many values will be returned, but that we do know the types of the attributes. Hence, to allocate enough space to store them, we must first inquire about the length of the attributes.

     INCLUDE 'netcdf.inc'
        ...
     PARAMETER (MVRLEN=3)           ! max number of "valid_range" values
     PARAMETER (MTLEN=80)           ! max length of "title" attribute
     INTEGER STATUS, NCID
     INTEGER RHID                   ! variable ID
     INTEGER VRLEN, TLEN            ! attribute lengths
     DOUBLE PRECISION VRVAL(MVRLEN) ! vr attribute values
     CHARACTER*80 TITLE             ! title attribute values

   ...
STATUS = NF_OPEN ('foo.nc', NF_WRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID (NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
! find out attribute lengths, to make sure we have enough space
STATUS = NF_INQ_ATTLEN (NCID, RHID, 'valid_range', VRLEN)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_INQ_ATTLEN (NCID, NF_GLOBAL, 'title', TLEN)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
! get attribute values, if not too big
IF (VRLEN .GT. MVRLEN) THEN
    WRITE (*,*) 'valid_range attribute too big!'
    CALL EXIT
ELSE
    STATUS = NF_GET_ATT_DOUBLE (NCID, RHID, 'valid_range', VRVAL)
    IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
ENDIF
IF (TLEN .GT. MTLEN) THEN
    WRITE (*,*) 'title attribute too big!'
    CALL EXIT
ELSE
    STATUS = NF_GET_ATT_TEXT (NCID, NF_GLOBAL, 'title', TITLE)
    IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
ENDIF

7.6 NF COPY ATT

The function NF COPY ATT copies an attribute from one open netCDF dataset to another. It can also be used to copy an attribute from one variable to another within the same netCDF.

Usage

INTEGER FUNCTION NF_COPY_ATT (INTEGER NCID_IN, INTEGER VARID_IN,
                              CHARACTER*(*) NAME, INTEGER NCID_OUT,
                              INTEGER VARID_OUT)

NCID_IN The netCDF ID of an input netCDF dataset from which the attribute will be copied, from a previous call to NF OPEN or NF CREATE.

VARID_IN ID of the variable in the input netCDF dataset from which the attribute will be copied, or NF GLOBAL for a global attribute.

NAME Name of the attribute in the input netCDF dataset to be copied.

NCID_OUT The netCDF ID of the output netCDF dataset to which the attribute will be copied, from a previous call to NF OPEN or NF CREATE. It is permissible for the input and output netCDF IDs to be the same. The output netCDF dataset should be in define mode if the attribute to be copied does not already exist for the target variable, or if it would cause an existing target attribute to grow.

VARID_OUT

ID of the variable in the output netCDF dataset to which the attribute will be copied, or NF GLOBAL to copy to a global attribute.

Errors

NF COPY ATT returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The input or output variable ID is invalid for the specified netCDF dataset.

• The specified attribute does not exist.

• The output netCDF is not in define mode and the attribute is new for the output dataset is larger than the existing attribute.

• The input or output netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF COPY ATT to copy the variable attribute units from the variable rh in an existing netCDF dataset named foo.nc to the variable avgrh in another existing netCDF dataset named bar.nc, assuming that the variable avgrh already exists, but does not yet have a units attribute:

INCLUDE 'netcdf.inc'
   ...
INTEGER STATUS             ! error status
INTEGER NCID1, NCID2       ! netCDF IDs
INTEGER RHID, AVRHID       ! variable IDs
   ...
STATUS = NF_OPEN ('foo.nc', NF_NOWRITE, NCID1)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_OPEN ('bar.nc', NF_WRITE, NCID2)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID (NCID1, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_INQ_VARID (NCID2, 'avgrh', AVRHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_REDEF (NCID2)  ! enter define mode
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
! copy variable attribute from "rh" to "avgrh"
STATUS = NF_COPY_ATT (NCID1, RHID, 'units', NCID2, AVRHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...

     STATUS = NF_ENDDEF (NCID2) ! leave define mode
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

7.7 NF RENAME ATT

The function NF RENAME ATT changes the name of an attribute. If the new name is longer than the original name, the netCDF dataset must be in define mode. You cannot rename an attribute to have the same name as another attribute of the same variable.

Usage

     INTEGER FUNCTION NF_RENAME_ATT (INTEGER NCID, INTEGER VARID,
                                     CHARACTER*(*) NAME,
                                     CHARACTER*(*) NEWNAME)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE

VARID ID of the attribute’s variable, or NF GLOBAL for a global attribute

NAME The current attribute name.

NEWNAME The new name to be assigned to the specified attribute. If the new name is longer than the current name, the netCDF dataset must be in define mode.

Errors

NF RENAME ATT returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The specified variable ID is not valid.

• The new attribute name is already in use for another attribute of the specified variable.

• The specified netCDF dataset is in data mode and the new name is longer than the old name.

• The specified attribute does not exist.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF RENAME ATT to rename the variable attribute units to Units for a variable rh in an existing netCDF dataset named foo.nc:

     INCLUDE "netcdf.inc"
        ...
     INTEGER STATUS   ! error status
     INTEGER NCID     ! netCDF ID
     INTEGER RHID     ! variable ID
        ...
     STATUS = NF_OPEN ("foo.nc", NF_NOWRITE, NCID)
     IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
        ...
     STATUS = NF_INQ_VARID (NCID, "rh", RHID)

IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
! rename attribute
STATUS = NF_RENAME_ATT (NCID, RHID, "units", "Units")
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

7.8 NF DEL ATT

The function NF DEL ATT deletes a netCDF attribute from an open netCDF dataset. The netCDF dataset must be in define mode.

Usage

INTEGER FUNCTION NF DEL ATT (INTEGER NCID, INTEGER VARID, CHARAC-TER*(*) NAME)

NCID NetCDF ID, from a previous call to NF OPEN or NF CREATE.

VARID ID of the attribute’s variable, or NF GLOBAL for a global attribute.

NAME The name of the attribute to be deleted.

Errors

NF DEL ATT returns the value NF NOERR if no errors occurred. Otherwise, the returned status indicates an error. Possible causes of errors include:

• The specified variable ID is not valid.

• The specified netCDF dataset is in data mode.

• The specified attribute does not exist.

• The specified netCDF ID does not refer to an open netCDF dataset.

Example

Here is an example using NF DEL ATT to delete the variable attribute Units for a variable rh in an existing netCDF dataset named foo.nc:

INCLUDE 'netcdf.inc'
   ...
INTEGER STATUS            ! error status
INTEGER NCID              ! netCDF ID
INTEGER RHID              ! variable ID
   ...
STATUS = NF_OPEN ('foo.nc', NF_WRITE, NCID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
STATUS = NF_INQ_VARID (NCID, 'rh', RHID)
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
   ...
! delete attribute
STATUS = NF_REDEF (NCID)  ! enter define mode

IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_DEL_ATT (NCID, RHID, 'Units')
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)
STATUS = NF_ENDDEF (NCID) ! leave define mode
IF (STATUS .NE. NF_NOERR) CALL HANDLE_ERR(STATUS)

Appendix A NetCDF 2 to NetCDF 3 Fortran 77 Transition Guide

A.1 Overview of FORTRAN interface changes

NetCDF version 3 includes a complete rewrite of the netCDF library. It is about twice as fast as the previous version. The netCDF file format is unchanged, so files written with version 3 can be read with version 2 code and vice versa.

The core library is now written in ANSI C. You must have an ANSI C compiler to compile this version. The FORTRAN interface is layered on top of the C interface using a different technique than was used in netCDF-2.

Rewriting the library offered an opportunity to implement improved C and FORTRAN interfaces that provide some significant benefits:

• type safety, by eliminating the need to use type punning in arguments;

• automatic type conversions, by eliminating the undesirable coupling between the language-independent external netCDF types (NF BYTE, ..., NF DOUBLE) and language-dependent internal data types (INT*1, ..., DOUBLE PRECISION);

• support for future enhancements, by eliminating obstacles to the clean addition of support for packed data and multithreading;

• more standard error behavior, by uniformly communicating an error status back to the calling program in the return value of each function.

It is not necessary to rewrite programs that use the version 2 FORTRAN interface, because the netCDF-3 library includes a backward compatibility interface that supports all the old functions, globals, and behavior. We are hoping that the benefits of the new interface will be an incentive to use it in new netCDF applications. It is possible to convert old applications to the new interface incrementally, replacing netCDF-2 calls with the corresponding netCDF-3 calls one at a time.

Other changes in the implementation of netCDF result in improved portability, maintainability, and performance on most platforms. A clean separation between I/O and type layers facilitates platform-specific optimizations. The new library no longer uses a vendorprovided XDR library, which simplifies linking programs that use netCDF and speeds up data access significantly in most cases.

A.2 The New FORTRAN Interface

First, here’s an example of FORTRAN code that uses the netCDF-2 interface:

     ! Use a buffer big enough for values of any type
     DOUBLE PRECISION DBUF(NDATA)
     REAL RBUF(NDATA)
     ...
     EQUIVALENCE (RBUF, DBUF), ...
     INT XTYPE     ! to hold the actual type of the data
     INT STATUS    ! for error status
     ! Get the actual data type
     CALL NCVINQ(NCID, VARID, ...,XTYPE, ...)
     ...

! Get the data
CALL NCVGT(NCID, VARID, START, COUNT, DBUF, STATUS)
IF(STATUS .NE. NCNOERR) THEN
   PRINT *, 'Cannot get data, error code =', STATUS
   ! Deal with error
   ...
ENDIF
IF (XTYPE .EQ. NCDOUBLE) THEN
   CALL DANALYZE(DBUF)
ELSEIF (XTYPE .EQ. NCFLOAT) THEN
   CALL RANALYZE(RBUF)
...
ENDIF

Here’s how you might handle this with the new netCDF-3 FORTRAN interface:

! I want to use doubles for my analysis
DOUBLE PRECISION DBUF(NDATA)
INT STATUS
! So I use a function that gets the data as doubles.
STATUS = NF_GET_VARA_DOUBLE(NCID, VARID, START, COUNT, DBUF)
IF(STATUS .NE. NF_NOERR) THEN
   PRINT *, 'Cannot get data, ', NF_STRERROR(STATUS)
   ! Deal with error
   ...
ENDIF
CALL DANALYZE(DBUF)

The example above illustrates changes in function names, data type conversion, and error handling, discussed in detail in the sections below.

A.3 Function Naming Conventions

The netCDF-3 C library employs a new naming convention, intended to make netCDF programs more readable. For example, the name of the function to rename a variable is now NF RENAME VAR instead of the previous NCVREN.

All netCDF-3 FORTRAN function names begin with the NF prefix. The second part of the name is a verb, like GET, PUT, INQ (for inquire), or OPEN. The third part of the name is typically the object of the verb: for example DIM, VAR, or ATT for functions dealing with dimensions, variables, or attributes. To distinguish the various I/O operations for variables, a single character modifier is appended to VAR:

• VAR entire variable access

• VAR1 single value access

• VARA array or array section access

• VARS strided access to a subsample of values

• VARM mapped access to values not contiguous in memory

At the end of the name for variable and attribute functions, there is a component indicating the type of the final argument: TEXT, INT1, INT2, INT, REAL, or DOUBLE.

This part of the function name indicates the type of the data container you are using in your program: character string, 1-byte integer, and so on.

Also, all PARAMETER names in the public FORTRAN interface begin with the prefix NF . For example, the PARAMETER which was formerly MAXNCNAM is now NF MAX NAME, and the former FILFLOAT is now NF FILL FLOAT.

As previously mentioned, all the old names are still supported for backward compatibility.

A.4 Type Conversion

With the new interface, users need not be aware of the external type of numeric variables, since automatic conversion to or from any desired numeric type is now available. You can use this feature to simplify code, by making it independent of external types. The elimination of type punning prevents some kinds of type errors that could occur with the previous interface. Programs may be made more robust with the new interface, because they need not be changed to accommodate a change to the external type of a variable.

If conversion to or from an external numeric type is necessary, it is handled by the library. This automatic conversion and separation of external data representation from internal data types will become even more important in netCDF version 4, when new external types will be added for packed data for which there is no natural corresponding internal type, for example, arrays of 11-bit values.

Converting from one numeric type to another may result in an error if the target type is not capable of representing the converted value. (In netCDF-2, such overflows can only happen in the XDR layer.) For example, a REAL may not be able to hold data stored externally as an NF DOUBLE (an IEEE floating-point number). When accessing an array of values, an NF ERANGE error is returned if one or more values are out of the range of representable values, but other values are converted properly.

Note that mere loss of precision in type conversion does not return an error. Thus, if you read double precision values into an INTEGER, for example, no error results unless the magnitude of the double precision value exceeds the representable range of INTEGERs on your platform. Similarly, if you read a large integer into a REAL incapable of representing all the bits of the integer in its mantissa, this loss There are two new functions in netCDF-3 that don’t correspond to any netCDF-2 functions: NF INQ LIBVERS and NF STRERROR. The version ation The previous implementation returned an error when the same dimension was used more than once in specifying the shape of a variable in ncvardef. This restriction is relaxed in the netCDF-3 implementation, because an autocorrelation matrix is a good example where using the same dimension twice makes sense.

In the new interface, units for the IMAP argument to the NF PUT VARM and NF GET VARM families of functions are now in terms of the number of data elements of the desired internal type, not in terms of bytes as in the netCDF version-2 mapped access interfaces.

Following is a table of netCDF-2 function names and names of the corresponding netCDF-3 functions. For parameter lists of netCDF-2 functions, see the netCDF-2 User’s Guide.

NCABOR NF ABORT NCACPY NF COPY ATT NCADEL NF DEL ATT NCAGT NF GET ATT DOUBLE, NF GET ATT REAL, NF GET ATT INT,

NF GET ATT INT1, NF GET ATT INT2

NCAGTC NF GET ATT TEXT NCAINQ NF INQ ATT, NF INQ ATTID, NF INQ ATTLEN, NF INQ ATTTYPE NCANAM NF INQ ATTNAME NCAPT NF PUT ATT DOUBLE, NF PUT ATT REAL, NF PUT ATT INT,

NF PUT ATT INT1NF PUT

Appendix B Summary of FORTRAN 77 Interface

Input parameters are in upper case, output parameters are in lower case. The FORTRAN types of all the parameters are listed alphabetically by parameter name below the function declarations.

CHARACTER*80 FUNCTION  NF_INQ_LIBVERS()
CHARACTER*80 FUNCTION  NF_STRERROR  (NCERR)
INTEGER FUNCTION  NF_CREATE         (PATH, CMODE, ncid)
INTEGER FUNCTION  NF_OPEN           (PATH, MODE, ncid)
INTEGER FUNCTION  NF_SET_FILL       (NCID, FILLMODE, old_mode)
INTEGER FUNCTION  NF_REDEF          (NCID)
INTEGER FUNCTION  NF_ENDDEF         (NCID)
INTEGER FUNCTION  NF_SYNC           (NCID)
INTEGER FUNCTION  NF_ABORT          (NCID)
INTEGER FUNCTION  NF_CLOSE          (NCID)
INTEGER FUNCTION  NF_INQ            (NCID, ndims, nvars, ngatts,
                                     unlimdimid)
INTEGER FUNCTION  NF_INQ_NDIMS      (NCID, ndims)
INTEGER FUNCTION  NF_INQ_NVARS      (NCID, nvars)
INTEGER FUNCTION  NF_INQ_NATTS      (NCID, ngatts)
INTEGER FUNCTION  NF_INQ_UNLIMDIM   (NCID, unlimdimid)
INTEGER FUNCTION  NF_DEF_DIM        (NCID, NAME, LEN, dimid)
INTEGER FUNCTION  NF_INQ_DIMID      (NCID, NAME, dimid)
INTEGER FUNCTION  NF_INQ_DIM        (NCID, DIMID, name, len)
INTEGER FUNCTION  NF_INQ_DIMNAME    (NCID, DIMID, name)
INTEGER FUNCTION  NF_INQ_DIMLEN     (NCID, DIMID, len)
INTEGER FUNCTION  NF_RENAME_DIM     (NCID, DIMID, NAME)
INTEGER FUNCTION  NF_DEF_VAR        (NCID, NAME, XTYPE, NDIMS, DIMIDS,
                                     varid)
INTEGER FUNCTION  NF_INQ_VAR        (NCID, VARID, name, xtype, ndims,
                                     dimids, natts)
INTEGER FUNCTION  NF_INQ_VARID      (NCID, NAME, varid)
INTEGER FUNCTION  NF_INQ_VARNAME    (NCID, VARID, name)
INTEGER FUNCTION  NF_INQ_VARTYPE    (NCID, VARID, xtype)
INTEGER FUNCTION  NF_INQ_VARNDIMS   (NCID, VARID, ndims)
INTEGER FUNCTION  NF_INQ_VARDIMID   (NCID, VARID, DIMIDS)
INTEGER FUNCTION  NF_INQ_VARNATTS   (NCID, VARID, natts)
INTEGER FUNCTION  NF_RENAME_VAR     (NCID, VARID, NAME)
INTEGER FUNCTION  NF_PUT_VAR_TEXT   (NCID, VARID, TEXT)
INTEGER FUNCTION  NF_GET_VAR_TEXT   (NCID, VARID, text)
INTEGER FUNCTION  NF_PUT_VAR_INT1   (NCID, VARID, I1VAL)
INTEGER FUNCTION  NF_GET_VAR_INT1   (NCID, VARID, i1val)
INTEGER FUNCTION  NF_PUT_VAR_INT2   (NCID, VARID, I2VAL)
INTEGER FUNCTION  NF_GET_VAR_INT2   (NCID, VARID, i2val)
INTEGER FUNCTION  NF_PUT_VAR_INT    (NCID, VARID, IVAL)
INTEGER FUNCTION  NF_GET_VAR_INT    (NCID, VARID, ival)

INTEGER FUNCTION  NF_PUT_VAR_REAL   (NCID, VARID, RVAL)
INTEGER FUNCTION  NF_GET_VAR_REAL   (NCID, VARID, rval)
INTEGER FUNCTION  NF_PUT_VAR_DOUBLE (NCID, VARID, DVAL)
INTEGER FUNCTION  NF_GET_VAR_DOUBLE (NCID, VARID, dval)
INTEGER FUNCTION  NF_PUT_VAR1_TEXT  (NCID, VARID, INDEX, TEXT)
INTEGER FUNCTION  NF_GET_VAR1_TEXT  (NCID, VARID, INDEX, text)
INTEGER FUNCTION  NF_PUT_VAR1_INT1  (NCID, VARID, INDEX, I1VAL)
INTEGER FUNCTION  NF_GET_VAR1_INT1  (NCID, VARID, INDEX, i1val)
INTEGER FUNCTION  NF_PUT_VAR1_INT2  (NCID, VARID, INDEX, I2VAL)
INTEGER FUNCTION  NF_GET_VAR1_INT2  (NCID, VARID, INDEX, i2val)
INTEGER FUNCTION  NF_PUT_VAR1_INT   (NCID, VARID, INDEX, IVAL)
INTEGER FUNCTION  NF_GET_VAR1_INT   (NCID, VARID, INDEX, ival)
INTEGER FUNCTION  NF_PUT_VAR1_REAL  (NCID, VARID, INDEX, RVAL)
INTEGER FUNCTION  NF_GET_VAR1_REAL  (NCID, VARID, INDEX, rval)
INTEGER FUNCTION  NF_PUT_VAR1_DOUBLE(NCID, VARID, INDEX, DVAL)
INTEGER FUNCTION  NF_GET_VAR1_DOUBLE(NCID, VARID, INDEX, dval)
INTEGER FUNCTION  NF_PUT_VARA_TEXT  (NCID, VARID, START, COUNT, TEXT)
INTEGER FUNCTION  NF_GET_VARA_TEXT  (NCID, VARID, START, COUNT, text)
INTEGER FUNCTION  NF_PUT_VARA_INT1  (NCID, VARID, START, COUNT, I1VALS)
INTEGER FUNCTION  NF_GET_VARA_INT1  (NCID, VARID, START, COUNT, i1vals)
INTEGER FUNCTION  NF_PUT_VARA_INT2  (NCID, VARID, START, COUNT, I2VALS)
INTEGER FUNCTION  NF_GET_VARA_INT2  (NCID, VARID, START, COUNT, i2vals)
INTEGER FUNCTION  NF_PUT_VARA_INT   (NCID, VARID, START, COUNT, IVALS)
INTEGER FUNCTION  NF_GET_VARA_INT   (NCID, VARID, START, COUNT, ivals)
INTEGER FUNCTION  NF_PUT_VARA_REAL  (NCID, VARID, START, COUNT, RVALS)
INTEGER FUNCTION  NF_GET_VARA_REAL  (NCID, VARID, START, COUNT, rvals)
INTEGER FUNCTION  NF_PUT_VARA_DOUBLE(NCID, VARID, START, COUNT, DVALS)
INTEGER FUNCTION  NF_GET_VARA_DOUBLE(NCID, VARID, START, COUNT, dvals)
INTEGER FUNCTION  NF_PUT_VARS_TEXT  (NCID, VARID, START, COUNT, STRIDE,
                                     TEXT)
INTEGER FUNCTION  NF_GET_VARS_TEXT  (NCID, VARID, START, COUNT, STRIDE,
                                     text)
INTEGER FUNCTION  NF_PUT_VARS_INT1  (NCID, VARID, START, COUNT, STRIDE,
                                     I1VALS)
INTEGER FUNCTION  NF_GET_VARS_INT1  (NCID, VARID, START, COUNT, STRIDE,
                                     i1vals)
INTEGER FUNCTION  NF_PUT_VARS_INT2  (NCID, VARID, START, COUNT, STRIDE,
                                     I2VALS)
INTEGER FUNCTION  NF_GET_VARS_INT2  (NCID, VARID, START, COUNT, STRIDE,
                                     i2vals)
INTEGER FUNCTION  NF_PUT_VARS_INT   (NCID, VARID, START, COUNT, STRIDE,
                                     IVALS)
INTEGER FUNCTION  NF_GET_VARS_INT   (NCID, VARID, START, COUNT, STRIDE,
                                     ivals)
INTEGER FUNCTION  NF_PUT_VARS_REAL  (NCID, VARID, START, COUNT, STRIDE,
                                     RVALS)
INTEGER FUNCTION  NF_GET_VARS_REAL  (NCID, VARID, START, COUNT, STRIDE,

                                     rvals)
INTEGER FUNCTION  NF_PUT_VARS_DOUBLE(NCID, VARID, START, COUNT, STRIDE,
                                     DVALS)
INTEGER FUNCTION  NF_GET_VARS_DOUBLE(NCID, VARID, START, COUNT, STRIDE,
                                     dvals)
INTEGER FUNCTION  NF_PUT_VARM_TEXT  (NCID, VARID, START, COUNT, STRIDE,
                                     IMAP, TEXT)
INTEGER FUNCTION  NF_GET_VARM_TEXT  (NCID, VARID, START, COUNT, STRIDE,
                                     IMAP, text)
INTEGER FUNCTION  NF_PUT_VARM_INT1  (NCID, VARID, START, COUNT, STRIDE,
                                     IMAP, I1VALS)
INTEGER FUNCTION  NF_GET_VARM_INT1  (NCID, VARID, START, COUNT, STRIDE,
                                     IMAP, i1vals)
INTEGER FUNCTION  NF_PUT_VARM_INT2  (NCID, VARID, START, COUNT, STRIDE,
                                     IMAP, I2VALS)
INTEGER FUNCTION  NF_GET_VARM_INT2  (NCID, VARID, START, COUNT, STRIDE,
                                     IMAP, i2vals)
INTEGER FUNCTION  NF_PUT_VARM_INT   (NCID, VARID, START, COUNT, STRIDE,
                                     IMAP, IVALS)
INTEGER FUNCTION  NF_GET_VARM_INT   (NCID, VARID, START, COUNT, STRIDE,
                                     IMAP, ivals)
INTEGER FUNCTION  NF_PUT_VARM_REAL  (NCID, VARID, START, COUNT, STRIDE,
                                     IMAP, RVALS)
INTEGER FUNCTION  NF_GET_VARM_REAL  (NCID, VARID, START, COUNT, STRIDE,
                                     IMAP, rvals)
INTEGER FUNCTION  NF_PUT_VARM_DOUBLE(NCID, VARID, START, COUNT, STRIDE,
                                     IMAP, DVALS)
INTEGER FUNCTION  NF_GET_VARM_DOUBLE(NCID, VARID, START, COUNT, STRIDE,
                                     IMAP, dvals)
INTEGER FUNCTION  NF_INQ_ATT        (NCID, VARID, NAME, xtype, len)
INTEGER FUNCTION  NF_INQ_ATTID      (NCID, VARID, NAME, attnum)
INTEGER FUNCTION  NF_INQ_ATTTYPE    (NCID, VARID, NAME, xtype)
INTEGER FUNCTION  NF_INQ_ATTLEN     (NCID, VARID, NAME, len)
INTEGER FUNCTION  NF_INQ_ATTNAME    (NCID, VARID, ATTNUM, name)
INTEGER FUNCTION  NF_COPY_ATT       (NCID_IN, VARID_IN, NAME,
                                     NCID_OUT, VARID_OUT)
INTEGER FUNCTION  NF_RENAME_ATT     (NCID, VARID, CURNAME, NEWNAME)
INTEGER FUNCTION  NF_DEL_ATT        (NCID, VARID, NAME)
INTEGER FUNCTION  NF_PUT_ATT_TEXT   (NCID, VARID, NAME, LEN, TEXT)
INTEGER FUNCTION  NF_GET_ATT_TEXT   (NCID, VARID, NAME, text)
INTEGER FUNCTION  NF_PUT_ATT_INT1   (NCID, VARID, NAME, XTYPE, LEN,
                                     I1VALS)
INTEGER FUNCTION  NF_GET_ATT_INT1   (NCID, VARID, NAME, i1vals)
INTEGER FUNCTION  NF_PUT_ATT_INT2   (NCID, VARID, NAME, XTYPE, LEN,
                                     I2VALS)
INTEGER FUNCTION  NF_GET_ATT_INT2   (NCID, VARID, NAME, i2vals)

INTEGER FUNCTION  NF_PUT_ATT_INT    (NCID, VARID, NAME, XTYPE, LEN,
                                     IVALS)
INTEGER FUNCTION  NF_GET_ATT_INT    (NCID, VARID, NAME, ivals)
INTEGER FUNCTION  NF_PUT_ATT_REAL   (NCID, VARID, NAME, XTYPE, LEN,
                                     RVALS)
INTEGER FUNCTION  NF_GET_ATT_REAL   (NCID, VARID, NAME, rvals)
INTEGER FUNCTION  NF_PUT_ATT_DOUBLE (NCID, VARID, NAME, XTYPE, LEN,
                                     DVALS)
INTEGER FUNCTION  NF_GET_ATT_DOUBLE (NCID, VARID, NAME, dvals)
INTEGER       ATTNUM       ! attribute number
INTEGER       attnum       ! returned attribute number
INTEGER       CMODE        ! NF_NOCLOBBER, NF_SHARE flags expression
INTEGER       COUNT        ! array of edge lengths of block of values
CHARACTER(*)  CURNAME      ! current name (before renaming)
INTEGER       DIMID        ! dimension ID
INTEGER       dimid        ! returned dimension ID
INTEGER       DIMIDS       ! list of dimension IDs
INTEGER       dimids       ! list of returned dimension IDs
DOUBLEPRECISION  DVAL      ! single data value
DOUBLEPRECISION  dval      ! returned single data value
DOUBLEPRECISION  DVALS     ! array of data values
DOUBLEPRECISION  dvals     ! array of returned data values
INTEGER       FILLMODE     ! NF_NOFILL or NF_FILL, for setting fill mode
INTEGER*1     I1VAL        ! single data value
INTEGER*1     I1val        ! returned single data value
INTEGER*1     I1VALS       ! array of data values
INTEGER*1     i1vals       ! array of returned data values
INTEGER*2     I2VAL        ! single data value
INTEGER*2     i2val        ! returned single data value
INTEGER*2     I2VALS       ! array of data values
INTEGER*2     i2vals       ! array of returned data values
INTEGER       IMAP         ! index mapping vector
INTEGER       INDEX        ! variable array index vector
INTEGER       IVAL         ! single data value
INTEGER       ival         ! returned single data value
INTEGER       IVALS        ! array of data values
INTEGER       ivals        ! array of returned data values
INTEGER       LEN          ! dimension or attribute length
INTEGER       len          ! returned dimension or attribute length
INTEGER       MODE         ! open mode, one of NF_WRITE or NF_NOWRITE
CHARACTER(*)  NAME         ! dimension, variable, or attribute name
CHARACTER(*)  name         ! returned dim, var, or att name
INTEGER       natts        ! returned number of attributes
INTEGER       NCERR        ! error returned from NF_xxx function call
INTEGER       NCID         ! netCDF ID of an open netCDF dataset
INTEGER       ncid         ! returned netCDF ID

INTEGER       NCID_IN      ! netCDF ID of open source netCDF dataset
INTEGER       NCID_OUT     ! netCDF ID of open destination netCDF dataset
INTEGER       NDIMS        ! number of dimensions
INTEGER       ndims        ! returned number of dimensions
CHARACTER(*)  NEWNAME      ! new name for dim, var, or att
INTEGER       ngatts       ! returned number of global attributes
INTEGER       nvars        ! returned number of variables
INTEGER       old_mode     ! previous fill mode, NF_NOFILL or NF_FILL,
CHARACTER(*)  PATH         ! name of netCDF dataset
REAL          RVAL         ! single data value
REAL          rval         ! returned single data value
REAL          RVALS        ! array of data values
REAL          rvals        ! array of returned data values
INTEGER       START        ! variable array indices of first value
INTEGER       STRIDE       ! variable array dimensional strides
CHARACTER(*)  TEXT         ! input text value
CHARACTER(*)  text         ! returned text value
INTEGER       unlimdimid   ! returned ID of unlimited dimension
INTEGER       VARID        ! variable ID
INTEGER       varid        ! returned variable ID
INTEGER       VARID_IN     ! variable ID
INTEGER       VARID_OUT    ! variable ID
INTEGER       XTYPE        ! external type: NF_BYTE, NF_CHAR, ... ,
INTEGER       xtype        ! returned external type

Index 135

Index

A

attributes, adding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

B

big-endian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

C

checksum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 chunking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 chunksizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 compiling with netCDF library. . . . . . . . . . . . . . . . . . 5 compound types, overview . . . . . . . . . . . . . . . . . . . . . 46 compression, setting parameters . . . . . . . . . . . . . . . 72 contiguous. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 creating dataset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

D

datasets, introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . 7 deflate. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 dimensions, adding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

E

endianness. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 enum type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

F

fill. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 fletcher32. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

G

groups, overview. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

I

interface descriptions . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

L

linking to netCDF library . . . . . . . . . . . . . . . . . . . . . . . 5 little-endian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

N

NF CREATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 NF ENDDEF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 NF OPEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 NF ABORT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 NF CLOSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 NF CLOSE, typical use. . . . . . . . . . . . . . . . . . . . . . . . . 1 NF COPY ATT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 NF CREATE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 NF CREATE, typical use. . . . . . . . . . . . . . . . . . . . . . . 1 NF CREATE PAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 NF DEF COMPOUND . . . . . . . . . . . . . . . . . . . . . . . 46 NF DEF DIM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 NF DEF DIM, typical use . . . . . . . . . . . . . . . . . . . . . . 1 NF DEF ENUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 NF DEF GRP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 NF DEF OPAQUE. . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 NF DEF VAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 NF DEF VAR, typical use. . . . . . . . . . . . . . . . . . . . . . 1 NF DEF VAR CHUNKING . . . . . . . . . . . . . . . . . . . 68 NF DEF VAR DEFLATE. . . . . . . . . . . . . . . . . . . . . 72 NF DEF VAR ENDIAN . . . . . . . . . . . . . . . . . . . . . . 77 NF DEF VAR FILL . . . . . . . . . . . . . . . . . . . . . . . . . . 70 NF DEF VAR FLETCHER32. . . . . . . . . . . . . . . . . 75 NF DEF VLEN . . . . . . . . . . . . . . . . . . . . . . . . . . . 54, 55 NF DEL ATT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 NF ENDDEF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 NF ENDDEF, typical use . . . . . . . . . . . . . . . . . . . . . . 1 NF FREE VLEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 NF GET ATT, typical use . . . . . . . . . . . . . . . . . . . . . 2 NF GET ATT type . . . . . . . . . . . . . . . . . . . . . . . . . 118 NF GET VAR, typical use . . . . . . . . . . . . . . . . . . . . . 2 NF GET VAR type . . . . . . . . . . . . . . . . . . . . . . . . . . 95 NF GET VAR1 type . . . . . . . . . . . . . . . . . . . . . . . . . 94 NF GET VARA type . . . . . . . . . . . . . . . . . . . . . . . . 97 NF GET VARM type . . . . . . . . . . . . . . . . . . . . . . . 101 NF GET VARS type . . . . . . . . . . . . . . . . . . . . . . . . . 99 NF GET VLEN ELEMENT . . . . . . . . . . . . . . . . . . 57 NF INQ Family . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 NF INQ, typical use . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 NF INQ ATT Family . . . . . . . . . . . . . . . . . . . . . . . . 116 NF INQ ATTNAME, typical use . . . . . . . . . . . . . . . 3 NF INQ COMPOUND . . . . . . . . . . . . . . . . . . . . . . . . 50 NF INQ COMPOUND FIELD . . . . . . . . . . . . . . . . 52 NF INQ COMPOUND FIELDDIM SIZES …. 52 NF INQ COMPOUND FIELDINDEX . . . . . . . . 52 NF INQ COMPOUND FIELDNAME . . . . . . . . . 52 NF INQ COMPOUND FIELDNDIMS . . . . . . . . 52 NF INQ COMPOUND FIELDOFFSET . . . . . . . 52 NF INQ COMPOUND FIELDTYPE. . . . . . . . . . 52 NF INQ COMPOUND NAME . . . . . . . . . . . . . . . . 50 NF INQ COMPOUND NFIELDS . . . . . . . . . . . . . 50 NF INQ COMPOUND SIZE . . . . . . . . . . . . . . . . . . 50 NF INQ DIM Family. . . . . . . . . . . . . . . . . . . . . . . . . . 39 NF INQ DIMID. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 NF INQ DIMID, typical use . . . . . . . . . . . . . . . . . . . . 2 NF INQ DIMIDS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 NF INQ ENUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 NF INQ ENUM IDENT . . . . . . . . . . . . . . . . . . . . . . 63 nf inq enum member . . . . . . . . . . . . . . . . . . . . . . . . . . 63

136 NetCDF Fortran 77 Interface Guide

NF INQ FORMAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 NF INQ GRP PARENT . . . . . . . . . . . . . . . . . . . . . . 35 NF INQ GRPNAME . . . . . . . . . . . . . . . . . . . . . . . . . . 33 NF INQ GRPNAME FULL . . . . . . . . . . . . . . . . . . . 34 NF INQ GRPNAME LEN . . . . . . . . . . . . . . . . . . . . 32 NF INQ GRPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 NF INQ LIBVERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 NF INQ NATTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 NF INQ NCID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 NF INQ NDIMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 NF INQ NVARS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 NF INQ OPAQUE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 NF INQ TYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 NF INQ TYPEIDS. . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 NF INQ UNLIMDIM . . . . . . . . . . . . . . . . . . . . . . . . . 21 NF INQ USER TYPE . . . . . . . . . . . . . . . . . . . . . . . . 45 NF INQ VAR family . . . . . . . . . . . . . . . . . . . . . . . . . . 80 NF INQ VAR CHUNKING . . . . . . . . . . . . . . . . . . . 69 NF INQ VAR DEFLATE . . . . . . . . . . . . . . . . . . . . . 74 NF INQ VAR ENDIAN . . . . . . . . . . . . . . . . . . . . . . . 79 NF INQ VAR FILL . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 NF INQ VAR FLETCHER32 . . . . . . . . . . . . . . . . . 76 NF INQ VARID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 NF INQ VARID, typical use. . . . . . . . . . . . . . . . . . . . 2 NF INQ VARIDS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 NF INSERT ARRAY COMPOUND. . . . . . . . . . . 49 NF INSERT COMPOUND . . . . . . . . . . . . . . . . . . . . 47 NF INSERT ENUM. . . . . . . . . . . . . . . . . . . . . . . . . . . 61 NF OPEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 NF OPEN PAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 NF PUT ATT, typical use. . . . . . . . . . . . . . . . . . . . . . 1 NF PUT ATT type . . . . . . . . . . . . . . . . . . . . . . . . . 114 NF PUT VAR, typical use. . . . . . . . . . . . . . . . . . . . . . 1 NF PUT VAR type . . . . . . . . . . . . . . . . . . . . . . . . . . 84 NF PUT VAR1 type . . . . . . . . . . . . . . . . . . . . . . . . . 82 NF PUT VARA type. . . . . . . . . . . . . . . . . . . . . . . . . 85 NF PUT VARM type . . . . . . . . . . . . . . . . . . . . . . . . 90 NF PUT VARS type . . . . . . . . . . . . . . . . . . . . . . . . . 88 NF PUT VLEN ELEMENT . . . . . . . . . . . . . . . . . . 56 NF REDEF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 NF RENAME ATT . . . . . . . . . . . . . . . . . . . . . . . . . . 122 NF RENAME DIM . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 NF RENAME VAR . . . . . . . . . . . . . . . . . . . . . . . . . . 107 NF SET DEFAULT FORMAT . . . . . . . . . . . . . . . . 26 NF SET FILL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 NF STRERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 NF SYNC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 NF VAR PAR ACCESS . . . . . . . . . . . . . . . . . . . . . 108 NF VAR PAR ACCESS, example . . . . . . . . . . . . 108

O

opaque type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

R

reading dataset with unknown names . . . . . . . . . . . 3 reading datasets with known names . . . . . . . . . . . . . 2

U

user defined types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 user defined types, overview . . . . . . . . . . . . . . . . . . . 43

V

variable length array type, overview. . . . . . . . . . . . 43 variable length arrays . . . . . . . . . . . . . . . . . . . . . . . . . . 53 variables, adding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 variables, checksum . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 variables, chunking . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 variables, contiguous. . . . . . . . . . . . . . . . . . . . . . . . . . . 68 variables, endian. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 variables, fill. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 variables, fletcher32. . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 variables, setting deflate . . . . . . . . . . . . . . . . . . . . . . . 72 VLEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 VLEN, defining . . . . . . . . . . . . . . . . . . . . . . . . 54, 55, 56